### v1

# Adaptive Visual Scene Understanding: Incremental Scene Graph Generation

Abstract
Scene graph generation (SGG) involves analyzing images to extract meaningful information about objects and their relationships. Given the dynamic nature of the visual world, it becomes crucial for AI systems to detect new objects and establish their new relationships with existing objects. To address the lack of continual learning methodologies in SGG, we introduce the comprehensive Continual ScenE Graph Generation (CSEGG) dataset along with 3 learning scenarios and 8 evaluation metrics. Our research investigates the continual learning performances of existing SGG methods on the retention of previous object entities and relationships as they learn new ones. Moreover, we also explore how continual object detection enhances generalization in classifying known relationships on unknown objects. We conduct extensive experiments benchmarking and analyzing the classical two-stage SGG methods and the most recent transformer-based SGG methods in continual learning settings, and gain valuable insights into the CSEGG problem. We invite the research community to explore this emerging field of study. All data and source code are publicly available at here.
1 Introduction
Scene graph generation (SGG) aims to extract object entities and their relationships in a scene. The resulting scene graph, carrying semantic scene structures, can be used for a variety of downstream tasks such as object detection(Szegedy et al., 2013), image captioning (Hassan et al., 2023; Aditya et al., 2015) , and visual question answering (Ghosh et al., 2019). Despite the notable advancements in SGG, current works have largely overlooked the critical aspect of continual learning. In the dynamic visual world, new objects and relationships are introduced incrementally, posing challenges for SGG models to adapt and generalize without forgetting previously acquired knowledge. This problem of Continual ScenE Graph Generation (CSEGG) holds great potential for various applications, such as real-time robotic navigation in dynamic environments and adaptive augmented reality experiences.
(a) Scene Graph Generation (SGG)
(b) Continual ScenE Graph Generation (CSEGG)
Figure 1: (a) A scene graph is a graph structure, where objects are represented as nodes (red boxes), and the relationships between objects are represented as edges connecting the corresponding nodes (green boxes). Each node in the graph contains information such as the object’s class label, and spatial location. The edges in the graph indicate the relationships between objects, often described by predicates. A scene graph can be parsed into a set of triplets, consisting of three components: a subject, a relationship predicate, and an object that serves as the target or object of the relationship. The graph allows for a compact and structured representation of the objects and their relationships within a visual scene. (b) An example CSEGG application is presented, where a robot continuously encounters new objects (blue) and new relationships (yellow) over time across new scenes.
Due to a scarcity of research specifically addressing the challenges of CSEGG, there is a pressing need for specialized investigations and methodologies to enable CSEGG. While the field of continual learning has witnessed significant growth in recent years, with a major focus on tasks such as image classification (Mai et al., 2021) , object detection (Wang et al., 2021a) , and visual question answering (Lei et al., 2022), these endeavors have largely neglected the distinctive complexities associated with CSEGG. Factors such as the long-tailed distribution of objects and relationships from each task in CSEGG and the intricate interplay in forgetting and generalization between incremental object detection and relationship classification remain unaddressed by previous works.
In this study, we establish a methodology for investigating CSEGG. Building upon existing SGG datasets (Krishna et al., 2017; Kuznetsova et al., 2020), we contribute a comprehensive CSEGG dataset containing 11 tasks, 108,249 images, 150 object classes, and 50 relationships, along with 3 learning protocols, 8 metrics, and a benchmark of continual learning baselines. Our research focuses on analyzing the extent to which existing scene graph generation methods, when combined with common continual learning techniques, experience forgetting as they learn new object entities and relationships. Additionally, we assess the generalization ability of CSEGG models in classifying known relationships between unknown objects. We expand our discussions on the problem motivation in Sec.A.1.4 and highlight the existence of a significant performance gap in this field. We believe that our work provides a valuable testbed for the research community to explore and advance this emerging area of study.
Main Contributions
1. Building upon existing SGG datasets, we introduced a large CSEGG dataset containing 108,249 images, 150 object classes, and 50 relationships.
2. We designed a systematic methodology for conducting CSEGG experiments and evaluating CSEGG models over a sequence of SGG tasks across 3 learning protocols. This helps the community expand the studies of CSEGG and benchmark future AI models.
3. In the experiments, we investigated the unique long-tailed challenge in CSEGG and observed that existing SGG models with continual learning methods do not perform well.
4. We assessed the ability of CSEGG models to generalize and classify known relationships on unfamiliar objects. Our findings revealed that continual learning models exhibiting less forgetting demonstrated better generalization capabilities in this regard.
2 Related Works
Scene Graph Generation Datasets.
Visual Phrase (Sadeghi & Farhadi, 2011) stands as one of the earliest datasets in the field of visual phrase recognition and detection. Over time, various large-scale datasets have emerged to tackle the challenges of Scene Graph Generation (SGG) (Johnson et al., 2015; Lu et al., 2016; Krishna et al., 2017; Kuznetsova et al., 2020; Liang et al., 2019; Zareian et al., 2020; Yang et al., 2019; Xu et al., 2017; Zhang et al., 2017; Dai et al., 2017; Li et al., 2017b; Zhang et al., 2019a). Among these, the Visual Genome dataset (Krishna et al., 2017) has played a pioneering role by providing comprehensive annotations of objects, attributes, and relationships in images. Another notable dataset is the Visual Relationship Detection dataset (Lu et al., 2016), which specifically focuses on detecting and classifying object relationships. More recently, the Open Image V6 dataset (Kuznetsova et al., 2020) has contributed to scene graph generation tasks by offering a vast collection of images accompanied by localized narratives. Despite the significant contributions of these datasets to SGG, none have been explicitly designed for the task of Continual ScenE Graph Generation (CSEGG). To address this gap, we leverage the existing datasets (Krishna et al., 2017) and curate a new one tailored specifically for CSEGG.
Scene Graph Generation Models. SGG models are categorized into two main approaches: top-down and bottom-up. Top-down approaches(Liao et al., 2019; Yu et al., 2017) typically rely on object detection as a precursor to relationship prediction. They involve detecting objects and then explicitly modeling their relationships using techniques such as rule-based reasoning(Lu et al., 2016) or graph convolutional networks (Yang et al., 2018). On the other hand, bottom-up approaches focus on jointly predicting objects and their relationships in an end-to-end manner (Li et al., 2017a; b; Xu et al., 2017). These methods often employ graph neural networks (Li et al., 2021; Zhang et al., 2019b) or message-passing algorithms (Xu et al., 2017) to capture the contextual information and dependencies between objects. Furthermore, recent works have explored the integration of language priors (Plummer et al., 2017; Lu et al., 2016; Wang et al., 2019) and attention mechanisms in transformers(Andrews et al., 2019) to enhance the accuracy and interpretability of scene graph generation. However, none of these works evaluate SGG models in incremental relationship learning and continual adaptation to new object entities. We benchmark SGG models in CSEGG, identify research gaps, and provide valuable insights about the development of future CSEGG models.
Continual Learning Methods. Existing continual learning works can be categorized into several approaches. (1) Regularization-based methods (Chaudhry et al., 2018; Zenke et al., 2017; Aljundi et al., 2018; Benzing, 2022) aim to mitigate catastrophic forgetting by employing regularization techniques in the parameter space, such as Elastic Weight Consolidation (EWC)(Kirkpatrick et al., 2017) or Synaptic Intelligence (SI)(Zenke et al., 2017). (2) Replay-based methods (Rolnick et al., 2019; Chaudhry et al., 2019; Riemer et al., 2018; Vitter, 1985; Rebuffi et al., 2017; Castro et al., 2018) utilize a memory buffer to store and replay past data during training, enabling the model to revisit and learn from previously seen examples, thereby reducing forgetting. The variants of these methods include generative replays(Shin et al., 2017; Wu et al., 2018; Ye & Bors, 2020; Rao et al., 2019), where synthetic data is generated and replayed. (3) Dynamic architecture-based approaches(Wang et al., 2022a) adapt the model’s architecture dynamically to accommodate new tasks. Techniques like network expansion(Yoon et al., 2017; Hung et al., 2019; Ostapenko et al., 2019) grow network capacity to accommodate additional knowledge without interfering with the existing ones. Despite the extensive investigation of continual learning in domains like image classification (Mai et al., 2021; Wang et al., 2022b; Cha et al., 2021) and object detection(Wang et al., 2021b; Shieh et al., 2020; Menezes et al., 2023), there is a notable dearth of research focusing on CSEGG. This work aims to bridge this gap by addressing the distinctive challenges of SGG, including the issue of long-tailed distribution (Desai et al., 2021; Nan et al., 2021; Chiou et al., 2021), within continual learning.
Figure 2: Three learning scenarios are introduced. From left to right, they are S1. relationship (Rel.) incremental learning (Incre.); S2. relationship and object (Rel. $+$ Obj.) Incre.; and S3. relationship generalization (Rel. Gen.) in Object Incre.. In S1 and S2, example triplets in the training (solid line) and test sets (dash line) from each task are presented. The training and test sets from the same task are color-coded. The new objects or relationships in each task are bold and underlined. In S3, one single test set (dashed gray box) is used for benchmarking the relationship generalization ability of object incre. learning models across all the tasks.
3 Continual ScenE Graph Generation (CSEGG) Benchmark
In CSEGG, we consider a sequence of tasks consisting of images and corresponding scene graphs with new objects or new relationships, or both. Let $D_{t}=\{(I_{i},G_{i})\}_{i=1}^{N_{t}}$ represent the dataset at task $t$, where $I_{i}$ denotes the $i$-th image and $G_{i}$ represents the associated scene graph. The scene graph $G_{i}$ comprises a set of object nodes $O_{i}$ and their corresponding relationships $R_{i}$. Each object node $o_{j}$ is defined by its class label $c_{j}$ and its spatial coordinates. Each relationship $r_{k}$ is represented by a triplet $(o_{s},p_{k},o_{o})$, where $o_{s}$ and $o_{o}$ denote the subject and object nodes, respectively, and $p_{k}$ represents the relationship predicate. The goal of CSEGG is to develop models that can incrementally learn new objects, new relationships, or both, without forgetting previously learned knowledge. Next, we introduce three continual learning scenarios.
(a) S1: Rel. Incre.
(b) S2: Rel. $+$ Obj. Incre.
(c) S3 Rel. Gen. In Obj. Incre.
Figure 3: Label distribution in each task in each learning scenario is presented. In scenario S1 (a) and scenario S3 (c), we use different colors to denote different tasks. The color gradient indicates the frequency of data within a task, with the lighter color denoting the smaller frequency of data in that category. Only the most frequent labels (relationship labels in (a) and object labels in (c)) are provided. See the legend for the total data size per task. In (b) scenario S2 on both objects and relationships, data distributions are presented in the form of small-world networks, where nodes denote object categories and the edges linking object pairs indicate relationships. Thickness in edges implies the diversity of relationships between object pairs. Same color conventions as (a) and (c) are applied. See the legend for triplet sizes. See Fig. S1,S2,S3 in Sec. A.3 for the full statistics of S1-3.
3.1 Learning Scenarios
We re-organize and divide the data from Visual Genome (Krishna et al., 2017) to cater to three continual learning scenarios below in our CSEGG dataset. The dataset contains 108,249 images, 150 objects, and 50 relationship categories. We follow the standard image splits for training, validation, and test sets (Xu et al., 2017).
Scenario 1 (S1): Relationship Incremental Learning. While existing continual object detection literature focuses on incrementally learning object attributes (Mai et al., 2021; Wang et al., 2022b; Cha et al., 2021; Wang et al., 2021b; Shieh et al., 2020; Menezes et al., 2023), incremental relationship classifications are equally important as it provides a deeper and more holistic understanding of the interactions and connections between objects within a scene. See Sec. A.1.1 for a concrete example application in medical imaging. To uncover contextual information and go beyond studies of object attributes, we introduce this scenario where new relationship predicates $p_{k}$ are incrementally added in each task (Fig. 2S1). There are 5 tasks in S1. To simulate the naturalistic settings where the frequency of relationship distribution is often long-tailed, we randomly and uniformly sample relationship classes from head, body and tail categories in Visual Genome (Krishna et al., 2017), and form a set of 10 relationship classes for each task. Thus, the relationships within a task are long-tailed; and the number of relationships from the head categories of each task is of the same scale. Different from continual object recognition and detection tasks, it is difficult to find unique images, in which ground truth scene graphs only contain relationships belonging to one specific task. To tackle this issue, we allow CSEGG models to see the same images over tasks, but the relationship labels are only provided in their given task (see Sec. A.1.5 for the design motivation). The same reasoning applies in S2 and S3. Example relationship classes from each task and their distributions are provided in Fig. 3(a).
Scenario 2 (S2): Scene Incremental Learning. To simulate the real-world scenario when there are demands for detecting new objects and new relationships over time in old and new scenes, we introduce this learning scenario where new objects $O_{i}$ and new relationship predicates $p_{k}$ are incrementally introduced over tasks (Fig. 2S2). See Sec. A.1.2 for two real-world example applications in robot collaborations on construction sites and video surveillance systems. To select the object and relationship classes from the original Visual Genome (Krishna et al., 2017) for S2, we have two design motivations in mind. First, in real-world applications, such as robotic navigation, robots might have already learned common relationships and objects in one environment. Incremental learning only happens on less frequent relationships and objects. (2) Transformer-based AI models typically require large amounts of training data to yield good performances. Training only on a small amount of data from tail classes often leads to close-to-chance performances. Thus, we take the common objects and relationships from the head classes in Visual Genome as one task, while the remaining less frequent objects and relationships from tail classes as the other task. This results in 2 tasks in total with the first task containing 100 object classes and 40 relationship classes. In the subsequent task, the CSEGG models are trained to continuously learn to detect 25 more object classes and 5 more relationship classes. Same as S1, both the object class and relationship class distributions are still long-tailed within a task (Fig. 3(b)).
Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning. We, as humans, have no problem at all recognizing the relationships of unknown objects with other nearby objects, even though we do not know the class labels of the unknown objects. This scenario is designed to investigate whether the CSEGG models can generalize as well as humans. See Sec. A.1.3 for two real-world applications in the deep sea and space explorations for autonomous navigation systems. Specifically, there are 4 tasks in total with each task containing 30 object classes and 35 relationship classes. In each subsequent task, the CSEGG models are trained to continuously learn to detect 30 more object classes and learn to classify the same set of 35 relationships among these objects. The class selection criteria for each task follow the same as S1, where the selections occur uniformly over head, body, and tail classes. Example object classes and their label distributions for each task are provided in Fig. 3(c). Different from S1 and S2, a standalone generalization test set is curated, where the objects are unknown and their classes do not overlap with any object classes in the training set but the relationships among these unknown objects are common to the training set of every task. The CSEGG models trained after every task are tested on the same generalization test sets.
3.2 Scene Graph Generation Backbone
(a) Schematic of SGTR (Li et al., 2022b)
(b) Overview of three continual learning baselines
Figure 4: Introduction to backbone SGG models and continual learning baselines. We use Scene graph Generation TRansformer (SGTR) (Li et al., 2022b) as the backbone SGG model (Sec. 3.2). SGTR consists of four modules indicated by each blue box. Arrows indicate the signal flows among modules. (b) Three continual learning baselines are listed: EWC (Kirkpatrick et al., 2017), Replay (Rolnick et al., 2019), Naive (Sec. 3.3) and PackNet (Mallya & Lazebnik, 2018)(Sec. 3.3). $\theta^{*}_{A}$ denotes the optimal network parameters after learning on task A. The arrows in colors indicate the shifts of network weights in the parameter space when learning Task B for different baselines.
We use the state-of-the-art one-stage Scene graph Generation TRansformer (SGTR) (Li et al., 2022b) and the traditional two-stage SGG model (Xu et al., 2017), named as CNN-SGG. CNN-SGG detects objects with Faster-RCNN(Girshick, 2015) backbone and infers their relationships separately via Iterative message passing (IMP)(Xu et al., 2017). For simplicity and consistency, we focus discussions solely on SGTR in the main text. For CNN-SGG, see Sec. A.2.2 for the introduction, training and implementation details over all three learning scenarios. We observed consistent relative CSEGG performance among all continual learning methods across both backbones (Fig. S15 and Sec. A.5).
Different from CNN-SGG, SGTR formulates the task as a bipartite graph construction problem. Briefly, we introduce how SGTR works (Fig. 4a). Given a scene image $I_{i}$, SGTR utilizes a 2D-CNN followed by a transformer-based encoder to extract image features. These features are further incorporated into a transformer-based decoder to predict object and subject nodes $O_{i}$. After that, the object-aware predicate nodes $R_{i}$ are formed based on both image features and object node features. Finally, a bipartite graph $G_{i}$ is constructed to collectively represent the scene with object $o_{i}$ and predicate nodes $r_{k}$, where the correspondence between these nodes is established based on the Hungarian matching algorithm (Kuhn, 1955). All experimental results are based on the average over 3 runs. We adapt the public source codes from (Li et al., 2022b) and (Wang et al., 2021b) for implementations of the continual learning algorithms on SGTR. We use the hyper-parameters provided in the code as the default values. See Sec. A.2.1 for training and implementation details of SGTR. All source code and data are publicly available here.
Techniques for Learning in Long-tailed Distribution. As shown in Fig. 3 and Sec. 3.1, the data in the training set of every task in all learning scenarios is long-tailed. We adopt the three existing techniques to alleviate the problem of imbalanced data distribution during training. LVIS(Gupta et al., 2019) is an image-level over-sampling strategy. The number of image repeats depends on the object classes with minimal frequency over the entire dataset. Bi-level sampling (BLS) (Li et al., 2021) combines image-level oversampling and instance-level undersampling. LVIS is used for image-level oversampling. At instance levels, a drop-out rate is applied. The more frequent an instance belonging to the common classes, the higher the dropout rate. The two-level data re-sampling achieves an effective trade-off between the head and tail classes. Equalized Focal Loss (EFL) (Li et al., 2022a) is an effective loss function, re-balancing the loss contribution of head and tail classes independently according to their imbalance degrees. During the training of SGTR, EFL is enabled all the time, regardless of the learning scenarios or continual learning algorithms.
3.3 Continual Learning Algorithms
We benchmark the continual learning baselines below (Fig.4b). Naive (lower bound) with the SGTR model as the backbone, is trained on each task in sequence without any measures to prevent catastrophic forgetting. EWC(Kirkpatrick et al., 2017) is a weight-regularization method, where the weights of the network are regularized in the parameter space, based on their “importance" to the previous tasks. PackNet(Mallya & Lazebnik, 2018) is a parameter-isolation method, where it iteratively prunes and pre-trains the network parameters so that it can sequentially pack multiple tasks within one network. Replay(Rolnick et al., 2019) includes a memory buffer with the capacity of storing $M$ percentages of images in the entire dataset as well as their corresponding ground truth object and predicate notations depending on the task at each learning scenario. We vary $M=$ 10%, 20%, and 100%. We randomly select images present in the current task and add them to the buffer. As the image may contain multiple object/relationship ground truths, the number of replays on ground truth notations in each task might vary. See Fig. S1(c)(d), Fig. S2(c)(d) and Fig. S3(c)(d) in Sec. A.3.4 where we report the number of ground truth notations stored in the memory buffer for each task in each learning scenario. However, these variations do not interfere with fair comparisons across replay methods with different memory buffer capacities. We also introduce Joint Training — an upper bound where the SGG model is trained on the entire CSEGG dataset.
To alleviate the problem of long-tailed data distribution, we also introduce BLS and LVIS in replay methods (Replay + BLS and Replay + LVIS). In the data re-balancing replay, the data in the memory buffer of Replay gets first re-sampled and then mixed with the data from the current task for training. As LVIS is an image-level up-sampling technique, the images containing the tail classes in the memory buffer are over-sampled. In contrast, BLS adds an extra instance-level under-sampling technique, which under-samples instances of head classes on the replay images. With these two approaches, the exact number of ground truth notations per image in the memory buffer varies over tasks in each learning scenario. We report these numbers in the Fig. S4 in Sec. A.3. The total number of instance replays for Replay+LVIS is slightly larger than the ones for Replay+BLS. We show the results of this effect in Sec. 4.2.
3.4 Evaluation Metrics
Same as existing SGG works (Xu et al., 2017; Li et al., 2022b), we adopt the evaluation metric recall@K (R@K) on the predicted scene graphs $G$. As CSEGG is long-tailed, we further report the results in mean recall (mR@K) over head, body, and tail relationship classes in Sec. A.4.2. For consistency, we provide the R@K results in the main text, while we provide and analyze results in mR@K in Fig. S8 and Fig. S9 in Sec. A.4.2.
Figure 5: Results of average recall and forgetting over tasks in Learning Scenarios 1 (a and b) and 2 (c and d) based on the SGTR backbone as the SGG model. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. X-axis indicates the task numbers. The higher Avg.R@20 and F@20, the better. See Fig. S5,S6 in Sec. A.4.1 for results of Avg.R@K, F@K, and mR@K, when K$=50,100$. See Sec. A.4.3 for results on FWT@K and BWT@K. See Sec. A.5 for the results based on the CNN-SGG backbone (Sec.3.2).
To assess the catastrophic forgetting of CSEGG models, we define Forgetfullness (F@K) as the difference in R@K on $D_{t=1}$ between the CSEGG models trained at task $t$ and task 1. An ideal CSEGG model could maintain the same $R@K$ on $D_{t=1}$ over tasks; thus, $F=0$ for all tasks. The more negative F is, the more severe in forgetting an model gets. To assess the overall recall of CSEGG models over tasks, we also report the continual average recall (Avg. R@K). Avg. R@K is computed as the average recall on all the data at the previous and current tasks $D_{i}$, where $i\in\{1,2,...,t\}$.
To assess whether the knowledge at previous tasks facilitates learning the new task and whether the knowledge at new tasks enhances the performances at older tasks, we introduce Forward Transfer (FWT@K) (Lin et al., 2022) and Backward Transfer (BWT@K)(Lopez-Paz & Ranzato, 2017). See Sec. A.2.3 and Sec. A.4.3 for its definitions, results, and analysis.
[…]
4 Results
4.1 Continual Scene Graph Generation Remains a Great Challenge.
We present Avg. R@20, F@20, FWT@20, and BWT@20 results for learning scenario 1 (S1) in Fig. 5 (a)(b) and Fig.S11(a)(b). Notably, all continual learning baselines start from a similar Avg.R@20 in Task 1 and their performance drops over subsequent tasks. This implies that catastrophic forgetting about learned relationships occurs when the CSEGG models learn new relationships. Among all the methods, the naive method takes no measures of preventing catastrophic forgetting, resulting in the largest drop in Avg.R@20 and F@20. In contrast, a replay method with all the old data to rehearse in the current task (Replay(100%)) yields the least forgetting and maintains a high Avg.R@20. Surprisingly, even though Replay(100%), as an upper bound, replays all the data in the current and previous tasks, there is still a drop in performance. This could possibly be due to the long-tailed data distribution in the memory buffer, which makes the rehearsal of tail classes even less frequent in new tasks, and thus, deteriorate the recall performances of tail classes. We also compared EWC versus the replay methods. Though EWC outperforms the naive baseline in earlier tasks, it fails in longer task sequences. Different from EWC, Replay with 10% still achieves a higher Avg.R@20 score of 8.55% and a higher F@20 of -22.21%. This aligns with the existing continual learning literature that replay methods are more effective than weight-regularization methods in eliminating catastrophic forgetting (Lesort et al., 2019). PackNet is a parameter isolation method. While PackNet outperforms EWC, its performance is inferior to that of Replay(10%). As expected, we also compare the replay methods with different memory buffer sizes. Replaying more old data helps CSEGG performances. Joint training demonstrates superior performance over all the tasks in Learning Scenario 1 as seen in Fig 5. This aligns with the existing continual learning literature that joint training is a superior upper bound than Replay(100%). As the knowledge carried forward is important for the subsequent tasks, we also permuted the task sequences and explored their role in CSEGG performances. Aligning with the existing literature (Singh et al., 2022), we found a prominent effect of task sequences in CSEGG (Fig. S13 and Sec. A.4.4).
Learning scenario 2 (S2) approximates the real-world CSEGG setting where there are constantly new demands in detecting new objects and new relationships simultaneously. The results of S2 in Avg.R@20 and F@20 are provided in Fig. 5(c)(d). Compared with S1, the overall Avg.R@20 and F@20 drop more significantly over tasks. For example, even with 20% memory buffer size, the replay method only achieves Avg. R@20 score of 6.57% and F@20 of -17.17% in Task 2. This suggests that the real-world CSEGG remains a challenging task and there still exists a large performance gap for state-of-the-art CSEGG methods. Moreover, we also made an interesting observation that Replay (100%) outperforms the upper bound of the joint training in the first task of Scenario 2. This performance difference could be attributed to the presence of long-tailed data distribution across tasks, with the first task containing more tailed classes than head classes. This is in contrast to the task splits in Scenario 1 where both head and tail classes are uniformly sampled for every task. Consequently, joint training struggles in the first task due to sub-optimal performance in tailed classes. To gain a qualitative understanding of CSEGG performances, we provide visualization results of the predicted scene graphs on example images over tasks for all the CSEGG baselines in Scenario 1 (Fig. S18 and Sec. A.6.1) and Scenario 2 (Fig. S19 and Sec. A.6.2).
Model Avg. R@20 $\uparrow$ F@20 $\uparrow$ T1 T5 T1 T5 Replay(10%) 28.7 8.55 0 -22.21 Replay(10%)+LVIS 28.7 14.38 0 -15.39 Replay(10%)+BLS 28.7 9.56 0 -22.4 Replay(100%) 28.7 16.17 0 -12.24 Table 1: Results at Task 1 and 5 in Learning Scenario 1 when sampling techniques on long-tailed distribution are applied. See Sec. 3.2 for the introduction to techniques used for long-tailed distributions. The best results are in bold.
[Uncaptioned image] Figure 6: Generalization results in Learning Scenario 3. See Sec. 3.4 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning baselines. Line types denote the IoU thresholds for determining correctly predicted bounding box locations.
4.2 Accounting for Long-tailed Distributions Enhances CSEGG
Due to the imbalanced data distribution in the real world, long-tailed distribution remains a unique challenge for CSEGG. Here, we introduce two data sampling techniques (LVIS and BLS) to counter-balance the long-tailed data distributions in the memory buffers as well as the feed-forward training tasks (Sec. 3.2 and Sec. 3.3). We report the results of Replay(10%)+BLS and Replay(10%)+LVIS in learning scenario S1. In Tab. 1, both long-tailed methods with replay(10%) outperform naive replays by an average margin of 3.42% in Avg.R@20 and 3.35% in F@20. This implies that data re-sampling techniques enhance general continual learning performances in long-tailed incremental learning settings. Indeed, we made the same observations after splitting the classes from each task into tail, body, and head classes and reporting their mR@K in the Fig. S10 in A.4.2. Interestingly, we see that Replay(10%)+BLS underperforms Replay(10%)+LVIS by 4.32% in Avg.R@20 and 6.42% in F@20. This contradicts the findings that BLS is more effective than LVIS in the classical SGG problem (Li et al., 2021). The performance discrepancy could be due to the difference in the number of replay instances in both approaches after these two data re-sampling methods are applied to the memory buffer (see Sec. 3.3). This emphasizes that the long-tailed learning methods explored in the SGG problem may not be effective in CSEGG. We need to explore new long-tailed learning methods specifically for CSEGG.
4.3 CSEGG Improves Generalization in Unknown Scene Understanding
[…]
Notably, we also found that the SGTR model is very good at generalizing to classify relationships Fig. 6 (b). Even in Task 1, both the naive method and the Replay(10%) achieve 45% recall of known relationships among unknown objects in the generalization test set. As the CSEGG models continuously learn to detect more new objects and classify their relationships in subsequent tasks, their relationship generalization ability among unknown objects saturates around Task 3. See Fig. S20 in Sec. A.6.3 for visualization examples.
5 Discussion
In the dynamic world, the incremental introduction of new objects and new relationships in scenes presents significant challenges for scene graph generation (SGG) models to effectively adapt and generalize without forgetting previously acquired semantic knowledge. However, despite the progress made in SGG and continual learning research, there remains a lack of comprehensive investigations specifically targeting the unique challenges of Continual Scene Graph Generation (CSEGG). To close this research gap, we take the initial steps of operationalizing CSEGG and introducing benchmarks, datasets, and evaluation protocols. Our study delves into three distinct learning scenarios, thoroughly examining the interplay between continual object detection and continual relationship classification for existing CSEGG methods under long-tailed class-incremental settings.
Our experimental results reveal intriguing insights. First, applying standard continual learning approaches combined with long-tailed techniques to SGG models yields moderate improvements. However, a notable performance gap persists between current CSEGG methods and the joint training upper bound. Second, we investigated the model’s generalization ability and found that the models are capable of generalizing to classify known relationships involving unfamiliar objects. Third, we compared the CSEGG performance of the traditional CNN-based and the transformer-based SGG models as backbones. We observed consistent relative CSEGG performance across all continual learning methods using both backbones, with CNN-SGG models underperforming SGTR-based ones.
Moving forward, there are several key avenues for future research. Our current endeavors focus on learning CSEGG problems from static images in Independent and Identically Distributed (i.d.d) manner, diverging from human learning from video streams. Future research can look into CSEGG problems on video streams. Our plans also involve expanding continual learning baselines and integrating more long-tailed distribution sampling techniques. Furthermore, we aim to construct a synthetic SGG dataset to systematically quantify the aspects of SGG that influence continual learning performance under controlled conditions. Although the CSEGG method holds promise for many downstream applications like monitoring systems, medical imaging, and autonomous navigation, we should also be aware of its misuse in privacy, data biases, fairness, security concerns, and misinterpretation. We invite the research community to join us in maintaining and updating the safe use of CSEGG benchmarks, thereby fostering its advancements in this field.
Ethics Statement
The development and deployment of Scene Graph Generation (SGG) technology present potential negative societal impacts that warrant careful consideration (Li et al., 2022b). Firstly, privacy concerns arise as SGG may inadvertently capture sensitive information from images, potentially violating privacy rights and raising surveillance issues. Secondly, bias and fairness challenges persist, as SGG algorithms can perpetuate biases present in training data, leading to discriminatory outcomes that reinforce societal inequalities. Misinterpretation and misclassification by SGG algorithms could result in misinformation and incorrect actions, impacting decision-making. The risk of manipulation and misuse of SGG-generated scene representations for malicious purposes is also a concern. For example, attackers might manipulate scene graphs to deceive systems or disrupt applications that rely on scene understanding.
Reproducibility Statement
We are committed to ensuring the reproducibility and transparency of our research. In accordance with the guidelines set forth by ICLR 2024, we provide detailed information to facilitate the replication of our experiments and results.
  1. 1.
    Code Availability: All code used for our experiments is available at here.
  2. 2.
    Data Availability: Any publicly accessible datasets used in our research are specified in the paper, along with their sources and access information.
  3. 3.
    Experimental Details: We have documented the specific details of our experiments, including hyper-parameters, model architectures, and pre-processing steps, to enable others to replicate our results.
We are dedicated to supporting the scientific community in replicating and building upon our work. We welcome feedback and collaboration to ensure the robustness and reliability of our research findings.
Acknowledgement
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-025), its NRFF award NRF-NRFF15-2023-0001, Mengmi Zhang’s Startup Grant from Agency for Science, Technology, and Research (A*STAR), and Early Career Investigatorship from Center for Frontier AI Research (CFAR), A*STAR. The authors declare that they have no competing interests. The authors would like to thank Stan Weixian Lei, Difei Gao, and Mike Shou for their feedback and suggestions.
Appendix A Appendix
A.1 Motivation for Learning Scenarios
Within this section, we present practical applications that demonstrate the real-world relevance of investigating CSEGG across all three learning scenarios.
A.1.1 Scenario 1 (S1): Relationship Incremental Learning
Within medical imaging, an agent must acquire the ability to detect cancerous cells within primary tumors, like colon adenocarcinoma. Subsequently, it must extend this proficiency to identifying the same cell types within metastatic growths that manifest in different bodily regions, such as lymph nodes or the liver. In this instance, the identical cancer cell disseminates to fresh organs or tissues, progressively establishing new relationships with other cells over the course of time.
A.1.2 Scenario 2 (S2): Scene Incremental Learning
The CSEGG model’s capacity to incorporate new objects and new relationships while retaining existing knowledge finds pivotal application in security contexts. Consider a company developing video-based security systems for indoor environments, capturing prevalent indoor objects and relationships. Expanding to outdoor settings like parking lots or restricted compounds demands retraining the model with new outdoor data alongside previous indoor data, ensuring operational effectiveness in both realms. The outdoor context introduces new objects like "cars" and relationships like "driving", distinct from indoor scenarios featuring "chair" and "sitting." Employing CSEGG allows the company to focus on new objects and relationships while retaining indoor insights.
Another real-world example would be a construction site where a team of robots is tasked with assembling various components to build a complex structure. Initially, during the foundation-laying phase, the robots are introduced to objects like "concrete blocks" and relationships like "stacking". As the construction advances to the wiring and installation phase, they encounter new objects like "wires" and relationships like "connecting," which were absent from earlier stages. The SGG model deployed in these robots needs to adapt incrementally to learn these new relationships without forgetting the existing ones. This ensures that the robots can effectively communicate and collaborate while comprehending the evolving scene and tasks, optimizing their construction efficiency and accuracy.
A.1.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
A prime example is the ongoing research on deep sea exploration for autonomous navigation systems, where undiscovered flora and fauna reside beneath the ocean’s surface. Encountering new and unidentified species becomes manageable through SGG’s ability to understand spatial relations. The robot discerns the object’s proximity or orientation even without precise identification of the spieces, enhancing its autonomous navigation ability. Likewise, in deep space exploration, SGG aids in recognizing spatial relationships with previously unseen space debris, aiding in path-planning. In essence, SGG’s relationship generalization empowers robots to navigate and plan routes in unfamiliar terrains, such as deep sea and deep space, where novel encounters demand adaptable responses.
A.1.4 Profound Impacts Beyond the Problem of CSEGG
We wish to emphasize that the challenge of continuous Scene Graph Generation (SGG) applied to real-world images exhibits two distinctive attributes. Techniques developed to address these attributes possess the potential to be widely applicable across various domains and subsequent tasks. Firstly, SGG in a continuous learning context frequently deals with data stemming from distributions that exhibit long-tailed characteristics. The methodologies formulated within the realm of Continual SGG can be adapted to address broader long-tailed continuous learning issues, such as classifying bird sounds in ecological studies. Secondly, the process of continual SGG demands the ability to engage in reasoning and integrate knowledge over time. As an illustration, in the context of retail inventory management, merely learning to identify durians (a tropical fruit uncommon in US supermarkets) is insufficient. The model must also cultivate the capability to integrate this new information into its existing knowledge database, ensuring that this new product is positioned alongside other fruits within the grocery store.
A.1.5 Design Motivations for Overlapping Images and Non-overlapping Labels across Tasks
The design of such images and label splits over tasks aligns with human learning scenarios where a parent teaches the baby to recognize different toys and objects in the bedroom. Though the baby is exposed to the same bedroom scenes multiple times, the parent only teaches the baby to detect and recognize one object at a time in a continual learning setting. In the future, we will expand our studies to cases where the SGG models learn from non-overlapping sets of training images for each task.
A.2 Implementation Details
A.2.1 SGTR
The SGTR is trained in two stages in a supervised manner. In stage 1, only object detection losses in DETR is (Carion et al., 2020) applied on $O_{i}$. In stage 2, only predicate entity prediction loss is applied on $R_{i}$, which can be further decomposed into L1 and GIOU losses for object/subject/predicate localization (Rezatofighi et al., 2019) and cross-entropy loss for object/subject/predicate classification. In learning scenario S1, we skip Stage 1, and directly load pre-trained weights of DETR for object detection on the entire training set of Visual Genome (Li et al., 2022b). In stage 2 of S1, we freeze the feature extractor, and fine-tune the rest parts of SGTR for predicate entity predictions. As only relationship classes are incrementally introduced in S1, we freeze the entire weights of DETR for detecting all the objects in the scene over tasks. However, empirical results suggest that fine-tuning transformer-based encoders in DETR helps downstream predicate predictions (Li et al., 2022b). Even with fine-tuning DETR in S1, we verify that there is minimal forgetting of detecting all objects in the scene over tasks (see Sec. S14 and Sec. A.4.5). Thus, the forgetting observed in S1 could only be attributed to incremental relationship learning. In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the feature extractor, pre-trained on ImageNet (Deng et al., 2009), and fine-tune the entire DETR (Carion et al., 2020) over all the tasks. Stage 2 of S2 and S3 is the same as S1.
Training the SGTR model involves two stages:
Object Detection Training: In this stage, a batch size of 32 is used. All methods are optimized using the Adam optimizer with a base learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. Object detection training is conducted only in the S2 and S3 scenarios. Each task in S2 is trained for 100 epochs, while each task in S3 is trained for 50 epochs. To expedite convergence, pre-trained weights on ImageNet are utilized before training on Task 1 for both S2 and S3.
SGG (Scene Graph Generation) Training: In this stage, the entire SGTR model is fine-tuned while keeping the 2D-CNN feature extractor frozen. A batch size of 24 is employed, and the Adam optimizer is used with a base learning rate of $8\times 10^{-5}$. In S1 and S3, each model is trained for 50 epochs per task, while in S2, 80 epochs per task are used. All models are trained on 4 A5000 GPUs.
A.2.2 CNN-SGG Backbone
Given a scene image $I_{i}$, CNN-SGG utilizes Faster-RCNN(Girshick, 2015) to generate a set of object proposals. The model subsequently extracts visual features of nodes and edges from the set of object proposals. Finally, both edge and node GRUs output a structured scene graph via message passing. The CNN-SGG is trained in two stages in a supervised manner. In stage 1, only object detection losses in Faster-RCNN(Ren et al., 2015) are applied on $O_{i}$. We use the cross entropy loss for the object class and $L1$ loss for the bounding box offsets. In stage 2, the visual feature extractor (VGG-16(Simonyan & Zisserman, 2014) pre-trained on ImageNet (Deng et al., 2009)) and GRUs layers are trained to predict the final object classes, bounding boxes, and relationship predicates using cross-entropy loss and $L1$ loss. In Learning Scenario 1 (S1), similar to the implementation details of SGTR in Sec.3.2, we skip Stage 1, and directly load pre-trained weights of Faster-RCNN for object detection on the entire training set of Visual Genome (Li et al., 2022b). In stage 2 of S1, we load the pre-trained weights of the visual feature extractor (pre-trained on ImageNet) and fine-tune the rest parts of the model. In stage 2 of S1, we load the pre-trained weights of visual feature extractor (pre-trained on ImageNet) and fine-tune the rest parts of the model. In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the Faster-RCNN, pre-trained on ImageNet Deng et al. (2009), and fine-tune it over all the tasks. Stage 2 of S2 and S3 follows the same training regimes as Stage 2 of S1.
Object Detection Training: All methods are optimized using the SGD optimizer with a base learning rate of $1\times 10^{-2}$ and a weight decay of $1\times 10^{-4}$. For training on the entire VG dataset, we train the model for 60 epochs with a batch size of 8 for both S2 and S3. To expedite convergence, pre-trained weights on ImageNet are utilized before training on Task 1 for both S2 and S3.
SGG (Scene Graph Generation) Training: A batch size of 12 is employed, and the SGD optimizer is used with a base learning rate of $1\times 10^{-2}$ and a weight decay of $1\times 10^{-4}$ In S1, each model is trained for 30 epochs. In S2, each model is trained for 15 epochs. In S3, each model is trained for 25 epochs. All models are trained on 4 A5000 GPUs.
A.2.3 Definitions of BWT and FWT Evaluation Metrics
To assess the influence that learning a task $t$ has on the performance of any previous tasks in CSEGG models, we also report Backward Transfer (BWT@K)(Lopez-Paz & Ranzato, 2017). BWT@K is defined as $BWT@K=\frac{1}{T-1}\sum_{i=1}^{T-1}R@K_{T,i}\>-\>R@K_{i,i}$, where $T$ denotes the total number of tasks in a learning scenario and $R@K_{i,j}$ denotes the continual learning model trained after task $i$ and tested in task $j$.
To assess the influence of previous tasks on the current task $t$ in CSEGG models, we also report Forward Transfer (FWT@K) (Lin et al., 2022). FWT is defined as $FWT@K=\frac{1}{T-1}\sum_{i=2}^{T}R@K_{i,i}\>-\>\overline{b@K}_{i,i}$, where $\overline{b@K}_{i,i}$ is the R@K for an independent model with random initialization trained in task $i$ and tested in task $i$.
A.3 Data Statistics
In this section, we provide various types of data statistics for all three learning scenarios. Specifically, we present statistics regarding the number of images, objects, and relationships involved in each task of each learning scenario. Additionally, we include data statistics pertaining to replay buffers of sizes 10%, 20%, and 100%.
A.3.1 Scenario 1 (S1): Relationship Incremental Learning
In this scenario, new relationship predicates $p_{k}$ are incrementally added in each task. The Learning Scenario 1 (S1) comprises five tasks, with each task consisting of 10 mutually exclusive relationships. Across all tasks, there is a common set of 150 objects present. Fig. S1 shows comprehensive data statistics for stages 1 and 2 over all the tasks.
As mentioned in Sec. 3.2, we skip the Stage 1 training in S1 and directly load pre-trained weights of DETR (Zhu et al. (2020)) for object detection on the entire training set of Visual Genome. Fig. S1(a) shows the distribution of objects in the entire training set of Visual Genome.
Fig. S1(b) displays the distribution of relationships during Training Stage 2 for each task in the Learning Scenario 1 (S1). Each task consists of 10 relationships that are mutually exclusive. Notably, a long tail pattern is observed in the distribution of relationships for each task. The legend of Fig. S1(b) indicates that there is a relatively uniform number of images present for each task in Stage 2 of the Learning Scenario 1 (S1).
The test sets for each task in S1 exhibit the distributions of the number of images and relationships that closely align with the distributions of the training sets depicted in Fig. S1(b).
A.3.2 Scenario 2 (S2): Scene Incremental Learning
In this learning scenario, new objects $O_{i}$ and new relationship predicates $p_{k}$ are incrementally introduced over tasks. The Learning Scenario 2 (S2) comprises of 2 tasks with the first task containing 100 objects and 40 relationships and the second task containing 25 objects and 5 relationships. Fig. S2 shows comprehensive data statistics pertaining to Learning Scenario 2 (S2).
As mentioned in Sec. 3.2, both the Stage 1 and Stage 2 training are present in S2. Fig. S2(a) presents the distribution of objects during Stage 1 for each task in the S2. The first task consists of 100 objects and the second task consists of 25 objects which are mutually exclusive. Notably, a long tail pattern is observed in the distribution of objects for each task. The legend of Fig. S2(a) indicates that there is a uniform number of images present for each task in Stage 1 of the S2.
Fig. S2(b) displays the distribution of relationships during Stage 2 for each task in the S2. The first task consists of 40 relationships and the second task consists of 5 relationships that are mutually exclusive. Notably, a long tail pattern is observed in the distribution of relationships for each task. The legend of Fig. S2(b) highlights that in Stage 2 of the S2, Task 1 has a noticeably higher number of images compared to Task 2. However, the number of relationship notations per class from Task 2 is higher than Task 1, since relationships from head classes are assigned to Task 2 (see Sec. 3.1).
The test sets for each task in S2 exhibit similar distributions in the number of images, objects, and relationships to the distributions of the training sets depicted in Fig. S2(a)(b).
A.3.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
In this learning scenario, there are a total of four tasks, with each task encompassing 30 distinct objects and 35 common relationships. Fig. S3 shows comprehensive data statistics pertaining to Learning Scenario 3 (S3).
As mentioned in Sec. 3.2, both Stage 1 and Stage 2 training are present in S3. Fig. S3(a) presents the distribution of objects during Stage 1 for each task in the S3. Each task consists of 30 objects which are mutually exclusive. Notably, a long tail pattern is observed in the distribution of objects for each task. The legend of Fig. S3(a) indicates that there is a relatively uniform number of images present for each task in Stage 1 of the S3. Fig. S3(b) displays the distribution of relationships during Stage 2 for each task in the S3. Each task consists of 35 relationships which are common over all the tasks. Notably, a long tail pattern is observed in the distribution of relationships for each task.
As mentioned in Sec. 3.1 of Scenario 3 (S3), a standalone generalization test set is curated for S3. This test set consists of 1942 images and features a completely different set of objects compared to those present in the training data. However, the test set maintains the same set of relationships as observed in the training data.
A.3.4 Replay Buffers
In our benchmarking process, we evaluate multiple replay baselines using a memory buffer with a fixed capacity to store a certain percentage (M) of images from the entire training dataset, along with their corresponding ground truth object and relationship notations specific to each learning scenario. We vary the value of M to be 10%, 20%, and 100%. As there could be multiple ground truth objects and relationships per image, we provide statistics on the number of ground truth notations stored in the memory buffer for each task in all the learning scenarios in this section. Specifically, we report the histograms for ground truth notations in the replay buffer for scenario 1 (Fig. S1), scenario 2 (Fig. S2), and scenario 3 (Fig. S3).
In general, over all three learning scenarios, we observe a long-tail distribution of notations per object or relationship class within each task. At stage 1, as we fix the memory capacity over tasks, the number of images stored in the memory buffer remains constant over tasks in most cases, except for memory capacity $M=100\%$. This is to maximize the diversity of information for replays. Note that although the number of images remain constant in the memory buffer, the number of images allocated to each task decreases given a fixed memory buffer capacity; and hence, we saw a decrease in the number of notations from previous tasks.
At stage 2, the images stored in stage 1 are carried over. However, as some images might not contain task-relevant relationship classes, the number of images used for training at stage 2 might vary over tasks. The notations stored in the memory buffer also follow the long-tail distributions.
In learning scenario 1, we reported the histograms of ground truth notations after two sampling techniques, LVIS and BLS, are applied in the memory buffers (Fig. S4). Both LVIS and BLS resampling methods aim to address the long-tail distribution issues by either reducing the prominence of head classes or over-sampling the object instances from the tail classes. Indeed, by comparing the histograms in Fig. S4(a)(b), we noticed such changes in histograms for tail and head classes. A similar trend in the class distribution per task emerges when applying long-tail distribution techniques to the exemplar dataset, as indicated in Fig.S4(c)(d).
Figure S1: Data Statistics for Learning Scenario 1 (S1): (a) Distribution of objects in the entire training set of Visual Genome during Stage 1. (b) Distribution of relationships during Stage 2 for each task in S1. (c) and (d) show distributions of relationships in the memory buffer for Replay 10% and Replay 100 % during Stage 2 for each task in S1. See the legend for the color codes of each task. The numbers in brackets in the legend in (b-d) denote the number of training images in the particular task. Zoom in to the figure to get the exact labels and the frequency associated with them.
Figure S2: Data Statistics for Learning Scenario 2 (S2): (a) Distribution of objects during Stage 1 for each task in S2. (b) Distribution of relationships during Stage 2 for each task in S2.(c) and (d) show distributions of objects in the memory buffer for Replay 10% and Replay 20% during Stage 1 for each task in S2. (e) and (f) show distributions of relationships in the memory buffer for Replay 10% and Replay 20% during Stage 2 for each task in S2. See the legend for the color codes of each task. The numbers in brackets in the legend in (a-f) denote the number of training images in the particular task. Zoom in to the figure to get the exact labels and the frequency associated with them.
Figure S3: Data Statistics for Learning Scenario 3 (S3): (a) Distribution of objects during Stage 1 for each task in S3. (b) Distribution of relationships during Stage 2 for each task in S3. (c) Distribution of objects in the memory buffer for Replay 10% during Stage 1 for each task in S3. (d) Distribution of relationships in the memory buffer for Replay 10% during Stage 2 for each task in S3. See the legend for the color codes of each task. The numbers in brackets in the legend in (a-d) denote the number of training images in the particular task. Zoom in to the figure to get the exact labels and the frequency associated with them.
Figure S4: Data Statistics for Learning Scenario 1 (S1) when sampling techniques on long-tailed distribution are applied: (a) Distribution of relationships during Stage 2 for each task in S1 when LVIS is applied. (b) Distribution of relationships in the memory buffer for Replay 10% during Stage 2 for each task in S1 when LVIS is applied.(c) Distribution of relationships during Stage 2 for each task in S1 when BLS is applied. (d) Distribution of relationships in the memory buffer for Replay 10% during Stage 2 for each task in S1 when BLS is applied. The numbers in brackets in the legend in (a-d) denote the number of training images in the particular task. Zoom in to the figure to get the exact labels and the frequency associated with them.
A.4 More Quantitative Result Analysis on Continual SGTR Methods
A.4.1 Results for K = 50, 100
Fig. S5 and S6 present the results for Avg.R@K and F@K with K values of 50 and 100, focusing on the Learning Scenarios 1 and 2. These results align with the findings discussed in Sec. 4.1, which considered Avg.R@20 and F@20. Notably, when K is increased to 50 and 100, both Avg.R and F exhibit higher absolute values compared to K=20 but the trends remain consistent. Fig. S7 presents the results of Gen.R@K for K=50,100 for Learning Scenario 3. Similar findings in Sec. 4.3 can be made here as well.
Figure S5: Results in AvgR@K and F@K over tasks in Learning Scenario 1 (S1) when K=50,100. See Fig. 5 for design conventions.
Figure S6: Results in AvgR@K and F@K over tasks in Learning Scenario 2 (S2) when K=50,100. See Fig. 5 for design conventions.
Figure S7: Results in Gen R@K over tasks in Learning Scenario 3 (S3) when K=50,100. See Fig. 6 for design conventions.
A.4.2 Mean Recall Results
Mean Recall@K (mR@K) calculates Recall@K for each relationship category independently and then reports their mean. This is a more fair evaluation metric when dealing with long tail distributions as it averages out the recall over all the relationship classes. In this section, we define mF@K and Avg.mR@K based on mR as described in Sec. 3.4. We report the results of mF@K and Avg.mR@K for learning scenarios 1 and 2 in Fig. S8 and Fig. S9. Note that the evaluations with mF@K and Avg.mR@K are not applicable in learning scenario 3. Though the absolute values of mF@K and Avg.mR@K are generally smaller than the values of F@K and Avg.R@K in Fig. 5, the same observations made in Sec. 4.1 can be applied here as well.
Figure S8: Results in Avg. mR@K and mF@K over tasks in Learning Scenario 1 (S1) when K=20,50,100. See Fig. 5 for design conventions.
Figure S9: Results in Avg. mR@K and mF@K over tasks in Learning Scenario 2 (S2) when K=20,50,100. See Fig. 5 for design conventions.
Figure S10: Results in Avg. mR@K and mF@K over tasks in Learning Scenario 1 (S1) when sampling techniques on long-tailed distribution are applied.. See Sec. 3.2 for the introduction to techniques used for long-tailed distributions. See Tab. 1 for design conventions.
A.4.3 Result Analysis on BWT@K and FWT@K
We provide the results of BWT@K and FWT@K in Fig. S11 for Learning Scenario 1. Consistent with the results of Avg.R@20 and F@20, we observe that among all methods, the naive method has the lowest FWT@20 and BWT@20 implying catastrophic forgetting. Once again, EWC outperforms the naive baseline by having higher FWT@20 and BWT@20 though the difference is small as it fails in longer task sequences. Although all the baselines have negative BWT@20 indicating forgetting, Replay(10%) yields a higher BWT@20 than PackNet, EWC and naive baseline pointing to a reduced level of forgetting. Replay(100%) exhibits the greatest BWT@20, suggesting the significance of revisiting previous samples to prevent forgetting and facilitate backward knowledge transfer. Similar to BWT@20, most of the baselines have negative FWT@20 implying that knowledge acquired from prior tasks slightly interferes with the learning process of the current task. In contrast, Replay(10%) shows positive FWT@20 indicating that knowledge acquired from prior tasks enhances the learning process of the current task. Surprisingly, while Replay(100%) excels in Avg.R@20, F@20, and BWT@20, Replay(10%) achieves a superior FWT@20. This suggests that Replay(100%) might not grasp the current task as effectively as Replay(10%), as it preserves the knowledge from previous tasks by replaying more old samples and forgetting less at the sacrifice of slow learning about the current task. Similar trends are observed and the same analysis can be applied in Learning Scenario 2 (Fig. S12).
Figure S11: Results of FWT (a) and BWT (b) over baselines in Learning Scenario 1 based on the SGTR backbone. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. The x-axis indicates the various baselines. The higher FWT@20 and BWT@20, the better.
Figure S12: Results of FWT (a) and BWT (b) over baselines in Learning Scenario 2 based on the SGTR backbone. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. The x-axis indicates the various baselines. The higher FWT@20 and BWT@20, the better.
A.4.4 Results for Different Task Sequences
Recent work (Singh et al., 2022) has highlighted the consistent and substantial curriculum effects in class-incremental learning and continual visual question-answering tasks. Inspired by these findings, we conducted experiments to assess the impact of the curricula in the context of CSEGG. To delve into its potential influence, we trained baselines with three distinct task sequences for Learning Scenario 1. Our results demonstrated that curriculum learning indeed shapes CSEGG performance within class-incremental settings. Notably, in Fig. S13(d), a large difference in Avg.R@20 between Order 1 and Order 3 emerges for the Replay (100%) baseline. Similarly, Fig. S13(a) reveals a substantial Avg.R@20 disparity between Order 2 and Order 3 for the Naive baseline. This trend extends to F@20, as depicted in Fig. S13(e)(f)(g)(h). These insights collectively affirm the significance of the curricula within CSEGG.
Figure S13: Results of F@K=20, Avg. R@K=20 over tasks on CSEGG models with the SGTR backbone in Learning Scenario 1 with different permutations of task sequences. (a),(e) denotes Avg.R@K and F@K for naive baseline. (b),(f) denotes Avg.R@K and F@K for EWC baseline. (c),(g) denotes Avg.R@K and F@K for Replay(10%) baseline. (d),(h) denotes Avg.R@K and F@K for Replay(100%) baseline. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. X-axis indicates the task numbers. The higher F, Avg.R@20, the better.
A.4.5 Minimal Forgetting in DETR
To validate the impact of fine-tuning the DETR model in training Stage 2 of learning scenario S1 on relationship predicate predictions and to ensure minimal forgetting occurs in object detection (Sec 3.2), we compare the mean Average Precision (mAP) for object detection on the entire test set of Visual Genome between the pre-trained DETR checkpoint from the paper (Li et al., 2022b) and the DETR models after fine-tuning on each task of S1.
As shown in Fig. S14, the results indicate a slight decrease of 0.4 in mAP from the pre-trained checkpoint to the DETR models over 5 tasks. This study provides evidence that fine-tuning DETR in S1 has negligible effects on forgetting. The forgetfulness observed in S1 can only be attributed to relationship incremental learning.
Figure S14: mAP performance on the entire Visual Genome Test Set. We used a pre-trained DETR checkpoint (Li et al., 2022b) as well as a naive baseline of SGTR continually trained on each task from S1, evaluated them on the entire Visual Genome’s test set, and reported their mAP performances. We observed a minimal decrease in mAP across all tasks.
A.5 Comparision between CNN-SGG Backbone and the SGTR Model
[…]
Figure S15: Results of average recall and forgetting over tasks in Learning Scenario 1 (a and b) for CNN-SGG, when K-20. See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. X-axis indicates the task numbers. The higher Avg.R@20 and F@20, the better.
Figure S16: Results of average recall and forgetting over tasks in Learning Scenario 2 (a and b) for CNN-SGG (K=20). See Sec. 3.3 for introduction to continual learning baselines. See Sec. 3.4 for explanations about evaluation metrics. X-axis indicates the task numbers. The higher Avg.R@20 and F@50, the better.
Fig. S17 provides the generalization results in detecting unknown objects and classifying known relationships among these objects in Learning Scenario 3 (S3) for CNN-SGG backbone. As elaborated in Sec. 4.3, an intriguing trend emerges: the CSEGG models demonstrate an increasing ability to identify unknown objects as the number of tasks rises. Furthermore, similar to our observations in Sec. 4.3, the CNN-SGG model exhibits proficiency in identifying known relationships between unfamiliar objects. However, this capability reaches a plateau as the task count escalates.
Figure S17: Generalization results in Learning Scenario 3 for CNN-SGG (K=20). See Sec. 3.4 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning baselines. Line types denote the IoU thresholds for determining correctly predicted bounding box locations.
A.6 Visualization Examples for All Learning Scenarios for Continual SGTR-based Models
In this section, we present visualization examples from each learning scenario to showcase the performance of the three continual SGTR-based models, namely Replay10%, EWC, and Naive, in three learning scenarios.
A.6.1 Learning Scenario 1 (S1)
Figure S18: Visualization example for Learning Scenario 1 (S1). The leftmost column in the figure displays the ground truth bounding boxes and scene graphs for each task in Learning Scenario 1 (S1). The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay 10%, Naive, and EWC). In all the scene graphs, red boxes indicate objects, while green boxes represent relationships. The direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet. For example, in the scene graph predicted by the EWC model after Task 5, the triplet is "trunk behind women", as the arrow goes from "trunk" to "behind" to "women". The time arrow on the left side of the figure demonstrates that the model is exposed to new data over time, with new relationships incrementally added, as described in Sec. 3.1.
Figure S19: Visualization example for Learning Scenario 2 (S2). The leftmost column shows the ground truth bounding boxes and scene graphs in each task of Learning Scenario 2 (S2). The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay 10%, Naive, and EWC). In all the scene graphs, red boxes indicate objects, while green boxes represent relationships. As explained in Fig. S18 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet. The time arrow on the left side of the figure demonstrates that the model is exposed to new data over time, with new objects and relationships incrementally added, as described in Sec. 3.1.
From Fig.S18 we observe that, in Task 1, the ground truth scene graph contains triplets of "on" relationship: "plate on table" and "hair on women". After training on task 1, all three models (Replay 10%, EWC, Naive) can accurately predict these triplets of "on" relationship.
In Task 2, triplets of "has" relationship are introduced: "plate has food" and "women has hair". After training on task 2 data, the Replay 10% model successfully remembers the triplets of "on" relationship ("plate on table", "hair on women") from Task 1 and predicts "women has hair". The Naive model forgets the triplets of "on" relationship and only predicts "women has hair". The EWC model remembers the triplets of "on" relationships and predicts "women has hair". None of the models predict "plate has food".
In Task 3, triplet of "at" relationship is introduced: "women at table". After training on task 3 data, the Replay 10% model remembers the previous triplets ("plate on table", "women has hair", "hair on women") and predicts "women at table". The Naive model forgets the previous triplets and only predicts "women at table". In contrast to its previous performance, the EWC model forgets the previous triplets and only predicts "women at table".
In Task 4, triplets related to "of" relationship are introduced: "hand of women", "neck of women", "arm of women", and "head of women". After training on task 4 data, the Replay 10% model remembers the triplets related to "on" and "has" relationships ("plate on table", "women has hair", "hair on women") from previous tasks but forgets the "at" relationship triplet (“women at table”). It only predicts "head of women" from the triplets introduced in Task 4. The Naive and EWC models both forget the "at" relationship triplet from the previous task but predict "head of women" and "arm of women" from the triplets introduced in Task 4.
In Task 5, triplet belonging to the "behind" relationship is introduced: "trunk behind women". After training on task 5 data, the Replay 10% model forgets the triplets related to "on" relationship ("plate on table", "hair on women") and only remembers the triplets related to "has" and "of" relationships ("women has hair", "head of women") learned from the previous task. It is not able to predict "trunk behind women". The Naive model, similar to its performance after previous tasks, fails to remember any triplets previously learned and only predicts "trunk behind women". The EWC model also fails to remember any triplets from the previous task and only predicts "trunk behind women".
A.6.2 Learning Scenario 2 (S2)
From Fig.S19 we observe that, in Task 1, the ground truth scene graph contains triplets: “man riding skateboard”, “man above skateboard”, and “shoe of skateboard”. After training on task 1 data, all three models (Replay 10%, Naive, EWC) predict “man riding skateboard” and “man above skateboard”. None of the models predict “shoe of skateboard”.
In Task 2, new triplets introduced are: “leg of person”, “person wearing sock”, “person has head”, and “head of person”. After training on task 2 data, the Replay 10% model only remembers “man riding skateboard” from the previous task, forgetting “man above skateboard”. Moreover, Replay 10% model can only predict “leg of person” from the triplets introduced in task 2. The Naive model forgets all the triplets from task 1 ( “man riding skateboard”, “man above skateboard”) and only predicts “leg of person” from the triplets introduced in task 2. Similar to Naive model, EWC model forgets all the triplets from task 1 ( “man riding skateboard”, “man above skateboard”) and only predicts “leg of person” from the triplets introduced in task 2.
Figure S20: Visualization example for Learning Scenario 3 (S3). The leftmost column shows the standalone ground truth bounding boxes and scene graphs in the generalization test set regardless of which task it is in Learning Scenario 3 (S3). The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Naive, Replay 10%). Similar to Fig. S18, and S19, the red boxes in all scene graphs indicate objects , while green boxes represent relationships. As explained in Fig. S18 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet. The time arrow on the left side of the figure demonstrates that the model is exposed to new objects over time as described in Sec. 3.1. For easy referral of object instances in the predicted scene graphs, we numbered the unknown bounding boxes in this figure, where the numbers are not actually present in the model predictions. Concretely, unk_1 refers to "mountain"; unk_2 is "house"; and unk_3 is "sheep".
A.6.3 Learning Scenario 3 (S3)
Fig. S20 illustrates the performance of the Replay 10% and Naive models in locating unknown objects and recognizing the relationships between these objects and other nearby unknown objects. The ground truth in Fig. S20 consists of three unknown objects: "mountain", "sheep", and "house", along with three relationships: "near", "behind", and "infront of" (between "mountain" and "house"), and a "near" relationship (between "sheep" and "house").
After training on Task 1 data, the Naive model can accurately locate the objects "mountain" and "house". However, no new object, such as "sheep", is located by the Naive model after training on Task 2 data. After training on Task 3 data, the Naive model is also able to locate the object "sheep" in addition to "mountain" and "house". Even after training on Task 4 data, the Naive model continues to locate all three objects: "mountain", "house", and "sheep". In contrast, the Replay 10% model, after training on Task 1 data, can only locate "mountain" and "house". This remains the same even after training on Task 2 and Task 3 data, where the Replay 10% model can still only locate the objects "mountain" and "house". However, after training on Task 4 data, the Replay 10% model is able to locate all the objects: "mountain", "house", and "sheep".
Regarding relationship generalization on unknown objects, the Naive model, after training on Task 1, can only predict the "near" relationship between the located objects "mountain" and "house" out of the three possible relationships. This performance remains the same even after Task 2. However, after training on Task 3, the Naive model can predict the "behind" relationship in addition to the "near" relationship between the located objects "mountain" and "house". After Task 4, the Naive model can predict the "behind" and "near" relationships between the located objects "mountain" and "house", as well as the "near" relationship between the located objects "sheep" and "house". In contrast, the Replay 10% model, after training on Task 1, can predict the "near" and “behind” relationships between the located objects "mountain" and "house". After Task 2, it can also predict the "infront of" relationship between the located objects "mountain" and "house" along with “near” and “behind” relationships. Even after Task 3, the Replay 10% model is still able to predict the "near", "behind", and "infront of" relationships between the located objects "mountain" and "house". After Task 4, as it can now locate the object "sheep", the Replay 10% model can also predict the "near" relationship between the objects "house" and "sheep", in addition to the existing relationships between "mountain" and "house".

### v3

@@
-Adaptive Visual Scene Understanding: Incremental Scene Graph Generation
+
 Abstract
-Scene graph generation (SGG) involves analyzing images to extract meaningful information about objects and their relationships.
-Given the dynamic nature of the visual world, it becomes crucial for AI systems to detect new objects and establish their new relationships with existing objects.
-To address the lack of continual learning methodologies in SGG, we introduce the comprehensive Continual ScenE Graph Generation (CSEGG) dataset along with 3 learning scenarios and 8 evaluation metrics.
-Our research investigates the continual learning performances of existing SGG methods on the retention of previous object entities and relationships as they learn new ones.
-Moreover, we also explore how continual object detection enhances generalization in classifying known relationships on unknown objects.
-We conduct extensive experiments benchmarking and analyzing the classical two-stage SGG methods and the most recent transformer-based SGG methods in continual learning settings, and gain valuable insights into the CSEGG problem.
-We invite the research community to explore this emerging field of study.
+Scene graph generation (SGG) analyzes images to extract meaningful information about objects and their relationships.
+In the dynamic visual world, it is crucial for AI systems to continuously detect new objects and establish their relationships with existing ones.
+Recently, numerous studies have focused on continual learning within the domains of object detection and image recognition.
+However, a limited amount of research focuses on a more challenging continual learning problem in SGG.
+This increased difficulty arises from the intricate interactions and dynamic relationships among objects, and their associated contexts.
+Thus, in continual learning, SGG models are often required to expand, modify, retain, and reason scene graphs within the process of adaptive visual scene understanding.
+To systematically explore Continual Scene Graph Generation (CSEGG), we present a comprehensive benchmark comprising three learning regimes: relationship incremental, scene incremental, and relationship generalization.
+Moreover, we introduce a “Replays via Analysis by Synthesis" method named RAS.
+This approach leverages the scene graphs, decomposes and re-composes them to represent different scenes, and replays the synthesized scenes based on these compositional scene graphs.
+The replayed synthesized scenes act as a means to practice and refine proficiency in SGG in known and unknown environments.
+Our experimental results not only highlight the challenges of directly combining existing continual learning methods with SGG backbones but also demonstrate the effectiveness of our proposed approach, enhancing CSEGG efficiency while simultaneously preserving privacy and memory usage.
 All data and source code are publicly available at here.
+Keywords:
+Scene graph generation Continual learning In-context symbolic replay Long tail distribution Compositional scene graphs
 1 Introduction
 Scene graph generation (SGG) aims to extract object entities and their relationships in a scene.
-The resulting scene graph, carrying semantic scene structures, can be used for a variety of downstream tasks such as object detection(Szegedy et al., 2013), image captioning (Hassan et al., 2023; Aditya et al., 2015) , and visual question answering (Ghosh et al., 2019).
+The resulting scene graph, carrying semantic scene structures, can be used for a variety of downstream tasks such as object detection[55], image captioning [19, 1] , and visual question answering [16].
 Despite the notable advancements in SGG, current works have largely overlooked the critical aspect of continual learning.
-In the dynamic visual world, new objects and relationships are introduced incrementally, posing challenges for SGG models to adapt and generalize without forgetting previously acquired knowledge.
+In the dynamic visual world, new objects and relationships are introduced incrementally, posing challenges for SGG models to account for new changes without forgetting previously acquired knowledge.
 This problem of Continual ScenE Graph Generation (CSEGG) holds great potential for various applications, such as real-time robotic navigation in dynamic environments and adaptive augmented reality experiences.
-(a) Scene Graph Generation (SGG)
-(b) Continual ScenE Graph Generation (CSEGG)
 Figure 1: (a) A scene graph is a graph structure, where objects are represented as nodes (red boxes), and the relationships between objects are represented as edges connecting the corresponding nodes (green boxes).
@@
 (b) An example CSEGG application is presented, where a robot continuously encounters new objects (blue) and new relationships (yellow) over time across new scenes.
-Due to a scarcity of research specifically addressing the challenges of CSEGG, there is a pressing need for specialized investigations and methodologies to enable CSEGG.
-While the field of continual learning has witnessed significant growth in recent years, with a major focus on tasks such as image classification (Mai et al., 2021) , object detection (Wang et al., 2021a) , and visual question answering (Lei et al., 2022), these endeavors have largely neglected the distinctive complexities associated with CSEGG.
-Factors such as the long-tailed distribution of objects and relationships from each task in CSEGG and the intricate interplay in forgetting and generalization between incremental object detection and relationship classification remain unaddressed by previous works.
-In this study, we establish a methodology for investigating CSEGG.
-Building upon existing SGG datasets (Krishna et al., 2017; Kuznetsova et al., 2020), we contribute a comprehensive CSEGG dataset containing 11 tasks, 108,249 images, 150 object classes, and 50 relationships, along with 3 learning protocols, 8 metrics, and a benchmark of continual learning baselines.
-Our research focuses on analyzing the extent to which existing scene graph generation methods, when combined with common continual learning techniques, experience forgetting as they learn new object entities and relationships.
-Additionally, we assess the generalization ability of CSEGG models in classifying known relationships between unknown objects.
-We expand our discussions on the problem motivation in Sec.A.1.4 and highlight the existence of a significant performance gap in this field.
-We believe that our work provides a valuable testbed for the research community to explore and advance this emerging area of study.
-Main Contributions
-1. Building upon existing SGG datasets, we introduced a large CSEGG dataset containing 108,249 images, 150 object classes, and 50 relationships.
+The field of continual learning has witnessed significant growth in recent years, with a major focus on tasks such as image classification [37], object detection [58], and visual question answering [26].
+However, these endeavors have largely neglected the distinctive complexities associated with CSEGG.
+Here, we highlight several unique challenges of CSEGG: (1) In contrast to object detection, SGG involves understanding and capturing the relationships between objects, which can be intricate and diverse.
+Consequently, in CSEGG, conveying the spatial and semantic relationships between objects demands adaptive reasoning from the dynamic scene.
+(2) SGG introduces a higher level of combinatorial complexity than object detection and image classification because each detected object pair may have multiple potential spatial and functional relationships.
+Thus, as new objects are introduced to the scenes, the complexity of relationships among all the objects increases significantly in a non-linear fashion.
+(3) The long-tailed distribution in SGG can be attributed to the inherent characteristics of real-world scenes, where certain objects are more prevalent than others.
+Consequently, CSEGG requires the computational models to adapt continually to the evolving long-tailed distributions over different scenes.
+Due to a scarcity of research specifically addressing these challenges of CSEGG, there is a pressing need for specialized investigations and methodologies to enable computational models with the ability of CSEGG.
+In this study, we re-organize existing SGG datasets [23, 25] to establish a novel and comprehensive CSEGG benchmark with 3 learning protocols as shown in Fig.
 2.
-We designed a systematic methodology for conducting CSEGG experiments and evaluating CSEGG models over a sequence of SGG tasks across 3 learning protocols.
-This helps the community expand the studies of CSEGG and benchmark future AI models.
-3. In the experiments, we investigated the unique long-tailed challenge in CSEGG and observed that existing SGG models with continual learning methods do not perform well.
-4.
-We assessed the ability of CSEGG models to generalize and classify known relationships on unfamiliar objects.
-Our findings revealed that continual learning models exhibiting less forgetting demonstrated better generalization capabilities in this regard.
+(S1).
+Relationship-incremental setting: an SGG agent learns to recognize new relationships among familiar objects within the same scene.
+(S2).
+Scene-incremental setting: an SGG agent is deployed in new scenes where it has to jointly learn to detect new objects and classify new relationships.
+(S3).
+Relationship generalization setting: an SGG agent generalizes to recognize known relationships among unknown objects, as the agent learns to recognize new objects.
+We curate a set of competitive CSEGG baselines by directly combining three major categories of continual learning methods with two SGG backbones and benchmark them in our CSEGG dataset.
+Their inferior performances show the difficulties of our benchmark tasks, which require the ability to expand, modify, retain, and reason scene graphs within the process of adaptive visual scene understanding.
+Specifically, the weight-regularization methods fail to estimate the importance of learnable parameters given the complicated model design in SGG backbones.
+Although image-replay methods retain knowledge from prior tasks through replays, the extensive combinatorial complexity of relationships among objects surpasses the complexity accommodated by a restricted set of replay images with efficient storage.
+Additionally, none of these baseline methods consider the shifts inherent in long-tailed distributions in dynamic scenes.
+To address the CSEGG challenges, we present a method called "Replays via Analysis by Synthesis", abbreviated as RAS.
+RAS employs scene graphs from previous tasks, breaks them down and re-composes them to generate diverse scene structures.
+These compositional scene graphs are then used for synthesizing scene images for replays.
+Due to its nature of symbolic replays, RAS doesn’t require the storage of original images, which often carry excessive and redundant details.
+This also ensures data privacy preservation and data efficiency.
+Furthermore, by synthesizing scenes using composable scene graphs, RAS maintains the semantic context and structure of previous scenes and also enhances the diversity of scene generation.
+To prevent biased predictions stemming from long-tailed distributions, we moderate the distribution of replayed scene graphs by balancing tail and head classes.
+This ensures a uniform sampling of relationships and objects during replays.
+Extensive experiments underscore the effectiveness of our approach.
+Network analysis reveals our crucial design choices that can be beneficial for the future development of CSEGG models.
 2 Related Works
 Scene Graph Generation Datasets.
-Visual Phrase (Sadeghi & Farhadi, 2011) stands as one of the earliest datasets in the field of visual phrase recognition and detection.
-Over time, various large-scale datasets have emerged to tackle the challenges of Scene Graph Generation (SGG) (Johnson et al., 2015; Lu et al., 2016; Krishna et al., 2017; Kuznetsova et al., 2020; Liang et al., 2019; Zareian et al., 2020; Yang et al., 2019; Xu et al., 2017; Zhang et al., 2017; Dai et al., 2017; Li et al., 2017b; Zhang et al., 2019a).
-Among these, the Visual Genome dataset (Krishna et al., 2017) has played a pioneering role by providing comprehensive annotations of objects, attributes, and relationships in images.
-Another notable dataset is the Visual Relationship Detection dataset (Lu et al., 2016), which specifically focuses on detecting and classifying object relationships.
-More recently, the Open Image V6 dataset (Kuznetsova et al., 2020) has contributed to scene graph generation tasks by offering a vast collection of images accompanied by localized narratives.
-Despite the significant contributions of these datasets to SGG, none have been explicitly designed for the task of Continual ScenE Graph Generation (CSEGG).
-To address this gap, we leverage the existing datasets (Krishna et al., 2017) and curate a new one tailored specifically for CSEGG.
-Scene Graph Generation Models.
+Visual Phrase [50] stands as one of the earliest datasets in the field of visual phrase recognition and detection.
+Over time, various large-scale datasets have emerged to tackle the challenges of Scene Graph Generation (SGG) [21, 36, 23, 25, 32, 69, 65, 63, 71, 11, 31, 72].
+Among these, the Visual Genome dataset [23] has played a pioneering role by providing rich annotations of objects, attributes, and relationships in images.
+Despite the significant contributions of these datasets to SGG, none focuses on continual learning in SGG.
+As the preliminary efforts towards CSEGG, we re-structure the Visual Genome dataset [23] and establish a novel and comprehensive CSEGG benchmark, where AI models are deployed to dynamic scenes where new objects and new relationships are introduced.
+Scene Graph Generation (SGG) Models.
 SGG models are categorized into two main approaches: top-down and bottom-up.
-Top-down approaches(Liao et al., 2019; Yu et al., 2017) typically rely on object detection as a precursor to relationship prediction.
-They involve detecting objects and then explicitly modeling their relationships using techniques such as rule-based reasoning(Lu et al., 2016) or graph convolutional networks (Yang et al., 2018).
-On the other hand, bottom-up approaches focus on jointly predicting objects and their relationships in an end-to-end manner (Li et al., 2017a; b; Xu et al., 2017).
-These methods often employ graph neural networks (Li et al., 2021; Zhang et al., 2019b) or message-passing algorithms (Xu et al., 2017) to capture the contextual information and dependencies between objects.
-Furthermore, recent works have explored the integration of language priors (Plummer et al., 2017; Lu et al., 2016; Wang et al., 2019) and attention mechanisms in transformers(Andrews et al., 2019) to enhance the accuracy and interpretability of scene graph generation.
-However, none of these works evaluate SGG models in incremental relationship learning and continual adaptation to new object entities.
-We benchmark SGG models in CSEGG, identify research gaps, and provide valuable insights about the development of future CSEGG models.
+Top-down approaches[33, 68] typically rely on object detection as a precursor to relationship prediction.
+They involve detecting objects and then explicitly modeling their relationships using techniques such as rule-based reasoning[36] or graph convolutional networks [64].
+On the other hand, bottom-up approaches focus on jointly predicting objects and their relationships in an end-to-end manner [30, 31, 63].
+These methods often employ graph neural networks [29, 73] or message-passing algorithms [63] to capture the contextual information and dependencies between objects.
+Furthermore, recent works have explored the integration of language priors [42, 36, 60] and attention mechanisms in transformers[3] to enhance the accuracy and interpretability of scene graph generation.
+However, none of these works evaluate SGG models in the context of continual learning.
+In our work, we directly combine continual learning methods with SGG backbones and benchmark these competitive baselines in CSEGG.
+Our results reveal the limitations of these methods and highlight the challenges of our CSEGG learning protocols.
 Continual Learning Methods.
 Existing continual learning works can be categorized into several approaches.
-(1) Regularization-based methods (Chaudhry et al., 2018; Zenke et al., 2017; Aljundi et al., 2018; Benzing, 2022) aim to mitigate catastrophic forgetting by employing regularization techniques in the parameter space, such as Elastic Weight Consolidation (EWC)(Kirkpatrick et al., 2017) or Synaptic Intelligence (SI)(Zenke et al., 2017).
-(2) Replay-based methods (Rolnick et al., 2019; Chaudhry et al., 2019; Riemer et al., 2018; Vitter, 1985; Rebuffi et al., 2017; Castro et al., 2018) utilize a memory buffer to store and replay past data during training, enabling the model to revisit and learn from previously seen examples, thereby reducing forgetting.
-The variants of these methods include generative replays(Shin et al., 2017; Wu et al., 2018; Ye & Bors, 2020; Rao et al., 2019), where synthetic data is generated and replayed.
-(3) Dynamic architecture-based approaches(Wang et al., 2022a) adapt the model’s architecture dynamically to accommodate new tasks.
-Techniques like network expansion(Yoon et al., 2017; Hung et al., 2019; Ostapenko et al., 2019) grow network capacity to accommodate additional knowledge without interfering with the existing ones.
-Despite the extensive investigation of continual learning in domains like image classification (Mai et al., 2021; Wang et al., 2022b; Cha et al., 2021) and object detection(Wang et al., 2021b; Shieh et al., 2020; Menezes et al., 2023), there is a notable dearth of research focusing on CSEGG.
-This work aims to bridge this gap by addressing the distinctive challenges of SGG, including the issue of long-tailed distribution (Desai et al., 2021; Nan et al., 2021; Chiou et al., 2021), within continual learning.
+(1) Regularization-based methods [22, 9, 70, 2, 4] aim to mitigate catastrophic forgetting by employing regularization techniques in the parameter space.
+(2) Dynamic architecture-based approaches[57, 67, 20, 41] adapt the model’s architecture dynamically to accommodate new tasks without interfering with the existing ones.
+(3) Replay-based methods [48, 10, 47, 56, 44, 7] utilize a memory buffer to store and replay past data during training, enabling the model to revisit and learn from previously seen examples, thereby reducing forgetting.
+The special variants of these methods include generative replay methods, such as [52, 62, 66, 43], where synthetic data is generated and replayed.
+Although these generative replay methods, as well as other continual learning methods, have been extensively studied in image classification [8, 61, 37] and object detection[59, 51, 39], few works focus on the challenges in CSEGG, such as adaptive reasoning from the dynamic scenes, the evolving long-tailed distribution across scenes, and the combinatorial complexity involving objects and their multiple relationships.
+In this work, we introduce a continual learning method, abbreviated as RAS (Replays via Analysis by Synthesis).
+To address the distinct challenges of CSEGG, RAS involves creating in-context synthetic scene images based on re-composable scene graphs from previous tasks to reinforce continual learning.
+The components in RAS facilitate memory-efficient training and preserve privacy while maintaining the scene diversity and scene context for SGG in dynamic environments.
+Scenarios #Tasks #Objs #Rels
+Eval.
+metrics
+SGG
+backbones
+CL
+baselines
+S1 5 All tasks: 150 Each task: 10 F, R, mF, mR, FWT, BWT, R_bbox, Gen_R Transformer based (SGTR) CNN based (TCNN) Joint, Naive, Replays[48], EWC[22], PackNet[38]
+S2 2
+Task 1: 100
+Task 2: 25
+Task 1: 40
+Task 2: 5
+S3 4 Each task: 30 All tasks: 35
+Table 1: Overview of three CSEGG learning scenarios.
+This table summarizes the three learning scenarios in CSEGG, including the number of tasks, the number of object (#Objs) and relationship (#Rels) classes, the evaluation metrics, the SGG-Backbones used, and the continual learning (CL) baselines.
+See Sec.
+3 for details.
+3 Continual ScenE Graph Generation Benchmark
+In CSEGG, to cater to the three continual learning scenarios below, we re-organize the Visual Genome [23] dataset and follow its standard image splits for training, validation, and test sets specified in [63].
+In each learning scenario, we consider a sequence of $T$ tasks consisting of images and corresponding scene graphs with new objects, or new relationships, or both.
+Let $D_{t}=\{(I_{i},G_{i})\}_{i=1}^{N_{t}}$ represent the dataset at task $t$, where $I_{i}$ denotes the $i$-th image and $G_{i}$ represents the associated scene graph.
+The scene graph $G_{i}$ comprises a set of object nodes $O_{i}$ and their corresponding relationships $R_{i}$.
+Each object node $o_{j}$ is defined by its class label $c_{j}$ and its bounding box locations and sizes $b_{j}$.
+Each relationship $r_{k}$ is represented by a triplet $(o_{s},p_{k},o_{o})$, where $o_{s}$ and $o_{o}$ denote the subject and object nodes, and $p_{k}$ represents the relationship predicate.
+3.1 Learning Scenarios
 Figure 2: Three learning scenarios are introduced.
@@
 relationship (Rel.) incremental learning (Incre.); S2.
-relationship and object (Rel.
-$+$ Obj.) Incre.; and S3.
+scene incremental learning; and S3.
 relationship generalization (Rel.
 Gen.) in Object Incre..
-In S1 and S2, example triplets in the training (solid line) and test sets (dash line) from each task are presented.
+In S1 and S2, example triplet labels in the training (solid line) and test sets (dash line) from each task are presented.
 The training and test sets from the same task are color-coded.
@@
 learning models across all the tasks.
-3 Continual ScenE Graph Generation (CSEGG) Benchmark
-In CSEGG, we consider a sequence of tasks consisting of images and corresponding scene graphs with new objects or new relationships, or both.
-Let $D_{t}=\{(I_{i},G_{i})\}_{i=1}^{N_{t}}$ represent the dataset at task $t$, where $I_{i}$ denotes the $i$-th image and $G_{i}$ represents the associated scene graph.
-The scene graph $G_{i}$ comprises a set of object nodes $O_{i}$ and their corresponding relationships $R_{i}$.
-Each object node $o_{j}$ is defined by its class label $c_{j}$ and its spatial coordinates.
-Each relationship $r_{k}$ is represented by a triplet $(o_{s},p_{k},o_{o})$, where $o_{s}$ and $o_{o}$ denote the subject and object nodes, respectively, and $p_{k}$ represents the relationship predicate.
-The goal of CSEGG is to develop models that can incrementally learn new objects, new relationships, or both, without forgetting previously learned knowledge.
-Next, we introduce three continual learning scenarios.
-(a) S1: Rel. Incre.
-(b) S2: Rel. $+$ Obj. Incre.
-(c) S3 Rel. Gen. In Obj. Incre.
-Figure 3: Label distribution in each task in each learning scenario is presented.
+Scenario 1 (S1): Relationship Incremental Learning.
+To uncover contextual information and go beyond studies of object detection and recognition, we introduce this scenario consisting of 5 tasks where 10 new relationship classes are incrementally added in every task (Fig.
+2, left; Tab.
+1).
+All object classes and their locations are made known to all CSEGG models over all the tasks.
+This scenario resembles a human learning scenario where a parent gradually teaches a baby to recognize new relationships among all objects in the same room, focusing on one new relationship at a time during continual learning.
+This scenario also has implications in medical imaging where identical cell types may form new relationships with nearby cells depending on the context.
+See Sec.
+0.A.1.1 for details.
+Scenario 2 (S2): Scene Incremental Learning.
+To simulate the real-world cases when there are demands for detecting new objects and new relationships from old to new scenes, we introduce this scenario where new objects and new relationships are incrementally introduced over tasks (Fig.
+2, middle; Tab.
+1).
+There are 2 tasks in total with the first task containing 100 object classes and 40 relationship classes with 25 more object classes and 5 more relationship classes in the second task.
+This aligns with the real-world use cases where common objects and relationships are learned in the first scene, and incremental learning in the second scene only happens on less frequent relationships and objects.
+See Sec.
+0.A.1.2 for details.
+Scenario 3 (S3): Relationship Generalization.
+Humans have no problem at all recognizing the relationships of unknown objects with other nearby objects.
+This scenario is designed to investigate the relationship generalization ability of CSEGG models.
+This capability is essential for real-world implications, such as in robotic navigation where it often encounters unknown objects and requires classifying their relationships.
+In total, there are four tasks, each introducing an incremental addition of 30 new object classes.
+All relationship classes are made known to all CSEGG models over all the tasks (Fig.
+2, right; Tab.
+1).
+Different from scenarios S1 and S2, a standalone generalization test set is curated, where the objects are unknown but the relationship classes among these unknown objects are common to the training set of every task.
+The CSEGG models trained after every task are tested on this standalone generalization test set to predict relationships among the unknown objects.
+See Sec.
+0.A.1.3 for details.
+Data sampling and distributions.
+To allocate data for every task of each scenario, we perform the following sampling strategies.
+In S1 and S3 above, either object or relationship classes are randomly sampled from the Visual Genome dataset and incrementally added to every task.
+Due to the inherent characteristics of real-world scenes, the long-tailed class distribution is present in S1 and S3.
+However, in S2, only tail classes are sampled and added in subsequent tasks.
+3.2 Competitive CSEGG Baselines
+Due to the scarcity of CSEGG works, we contribute a diverse set of competitive CSEGG baselines and implement them on our own.
+Each CSEGG baseline requires three components: a backbone model for scene graph generation (SGG), a continual learning (CL) method to prevent the SGG model from forgetting, and an optional data sampling technique to deal with imbalanced data at every task for training SGG models.
+Next, we introduce the 2 SGG backbones, the 5 continual learning methods, and the 5 optional data sampling techniques.
+See Sec.
+0.A.2 for implementation and training details of CSEGG baselines.
+SGG Backbones.
+We use the two state-of-the-art backbones: (1) one-stage Scene graph Generation TRansformer (SGTR) [28] and (2) the traditional Two-stage SGG model (TCNN) [63].
+Briefly, SGTR (Fig.
+S2(a)) uses a transformer-based architecture for image feature extraction and fusion.
+During training, [28] formulates SGG as a bipartite graph construction and matching problem.
+In contrast, TCNN detects objects with Faster-RCNN[17] backbone and infers their relationships separately via Iterative message passing [63].
+We use public source codes from [28] and [59] with default hyperparameters.
+Continual Learning Methods.
+We include the following continual learning methods (Fig.
+S2(b)): (1) Naive (lower bound) is trained on each task in sequence without any measures to prevent catastrophic forgetting.
+(2) EWC[22] is a weight-regularization method, where the weights of the network are regularized in the parameter space, based on their “importance" to the previous tasks.
+(3) PackNet[38] is a parameter-isolation method, iteratively pruning the network parameters after every task, so that it can sequentially pack multiple tasks within one network.
+(4) Replay@M[48] includes a memory buffer with the capacity of storing $M$ percentages of images in the entire dataset as well as their corresponding ground truth object and predicate notations depending on the task at each learning scenario.
+We vary $M=$ 10%, 20%, and 100%.
+(5) Joint Training is an upper bound where the SGG model is trained on the entire CSEGG dataset.
+Sampling Methods to Handle Long-Tailed Distribution.
+We adopt the five data sampling techniques to alleviate the problem of imbalanced data distribution during training.
+(1) LVIS[18] is an image-level over-sampling strategy for the tailed classes.
+(2) Bi-level sampling (BLS) [29] balances the trade-off between image-level oversampling for the tailed classes and instance-level under-sampling for the head classes.
+(3) Equalized Focal Loss (EFL) [27] is an effective loss function, re-balancing the loss contribution of head and tail classes according to their imbalanced distribution.
+EFL is enabled all the time for all the CSEGG baselines.
+In addition to applying data sampling techniques to the training sets, we can also apply LVIS and BLS techniques to the data stored in the replay buffer.
+We name these data sampling techniques applied during both training and replays as (4) LVIS@Replay and (5) BLS@Replay.
+3.3 Evaluation Metrics
+Same as existing SGG works [63, 28], we adopt the evaluation metric recall@K (R@K) on the top $K$ predicted triplets in the scene graphs $G$.
+As CSEGG is long-tailed, we further report the results in mean recall (mR@K) over the head, body, and tail classes.
+Forgetfullness (F), Average (Avg.) performance, Forward Transfer (FWT) [34] and Backward Transfer (BWT) [35] are standard evaluation metrics used for continual learning in image recognition and object detection tasks.
+In Scenario 1 and 2, we adapt these metrics to recalls R@K and introduce F@K, Avg.R@K, FWT@K, and BWT@K respectively for CSEGG settings.
+Similarly, we also adapt these metrics to mR@K.
+We explored CSEGG with K=20, 50, and 100.
+Since our results are consistent among Ks, we omit “@K" and analyze all the results based on K=20 in the entire text.
+In scenario S3, we evaluate all CSEGG methods in the standalone generalization test set, shared over all the tasks.
+To benchmark generalization abilities in unknown object localization and relationship classification among these unknown objects, we introduce two evaluation metrics: Gen Rbbox@K and Gen R@K.
+As the CSEGG models have never been taught to classify unknown objects, we discard the class labels of the bounding boxes and only evaluate the predicted bounding box locations with Gen Rbbox@K.
+To evaluate whether the predicted bounding box location is correct, we apply a hard threshold of Intersection over Union (IoU) between the predicted bounding box locations and the ground truth.
+Any predicted bounding boxes with their IoU values above the hard threshold are deemed to be correct.
+We vary IoU thresholds from 0.3, 0.5, to 0.7.
+To assess whether the CSEGG model generalizes to detect known relationships over unknown objects, we evaluate the recall Gen R@K of the predicted relationships $r_{k}$ only on correctly predicted bounding boxes.
+See Sec.
+0.A.3 for detailed definitions of these metrics.
+All experimental results are averaged over three runs.
+4 Replays via Analysis by Synthesis (RAS)
+Figure 3: Schematic of our proposed Replays via Analysis by Synthesis (RAS) method.
+At task $t+1$, our RAS stores all the triplet labels $U_{t}$, such as <man, on, horse>, from the previous tasks.
+It then re-composes these triplet labels to create in-context prompts, utilizing them as inputs to generative image models to synthesize images for replays.
+For predicting scene graphs on these synthesized images, we employ the frozen model $M_{t}$ from the preceding task $t$, marked with “snowflakes".
+Subsequently, these predicted scene graph notations, along with their corresponding synthesized images, contribute to “pseudo" replays, preventing the current model $M_{t+1}$ from experiencing forgetting.
+See Sec.
+4 for more details.
+To address the complexities in CSEGG, we introduce our “Replays via Analysis by Synthesis" method, dubbed RAS.
+Our RAS belongs to the group of generative replay methods for continual learning.
+However, different from the existing generative replay methods [52, 14, 15], our RAS leverages symbolic replays with state-of-the-art diffusion models.
+Moreover, rather than generating any random images for replays, our RAS is capable of generating in-context images following semantic rules, such as object co-occurrences.
+The schematic of our RAS is presented in Fig.
+3.
+Next, we focus on how RAS generates images and predicts SGG annotations on these images for replays.
+Image Generation.
+At the current task $t+1$, our RAS requires storing the frozen old model snapshot $M_{t}$ at the end of the previous task $t$ and all the triplet labels $U_{t}$, which are parsed from all the scene graphs aggregated from all the previous tasks.
+These triplets contain object labels, subject labels, and relationships among them.
+For example, <man, on, horse> and <man, in front of, horse> are two unique triplet labels.
+Unlike the traditional replay methods in continual learning literature, our method refrains from storing original images $I_{i}$ or scene graphs $G_{i}$ in the training sets, thereby eliminating storage issues and privacy concerns.
+To generate images $I^{\prime}_{j}$ for replays, our RAS feeds text prompts, which are formed by a set of chosen triplet labels and describe the diverse in-context scenes, into the state-of-the-art Stable Diffusion model [49].
+The small object on the table is most likely to be a spoon rather than an elephant.
+Context plays important roles in visual perceptions [5].
+To generate text prompts describing context-congruent scenes, we employ a context checker.
+First, the context checker uses the pre-trained large language model BERT [13] to extract embeddings for each triplet label in $U_{t}$.
+As BERT has been pre-trained on a large corpus of text data, it learns to capture context-relevant representations of words.
+Next, hierarchical clustering is performed on these embeddings using the agglomerative clustering algorithm [40].
+This ensures that each cluster contains only embeddings that are semantically close.
+The threshold for the agglomerative clustering algorithm is set to 0.6.
+As real-world images often contain complex scenes involving multiple triplets, we select any cluster with more than 3 triplet labels to create a text prompt for the image generation model.
+In practical applications, conducting agglomerative clustering on all triplet labels in $U_{t}$ is computationally demanding, as it requires computing pairwise embedding similarities among all the triplet labels.
+To address this, RAS opts for a more efficient approach during replays by selecting a subset of triplet labels and clustering their embeddings.
+Recognizing that real-world scenes often exhibit a long-tailed distribution with certain objects or relationships being more prevalent, we introduce the Long-Tailed Distribution (LTD) module for balancing this distribution in $U_{t}$.
+Unlike image-level sampling methods like BLS and LVIS [29, 18] discussed in Sec.
+3.2, our LTD module in RAS operates at the triplet level.
+For each triplet label, its dropout rate is determined proportionally to its frequency in $U_{t}$.
+Specifically, we define the drop-out rate $d_{k}$ for the $k$-th triplet as: $d_{k}=f_{k}/(\sum_{i=1}^{i=N}f_{i})*\alpha$, where $N$ is the total number of triplets in $U_{t}$, $\alpha=0.7$ is a scaling factor, and $f_{i}$ is the frequency of the $i$-th triplet in $U_{t}$.
+This sampling formula enables RAS to select triplets from tail classes more frequently compared to those from head classes.
+To generate a text prompt from the chosen triplet labels, we employ a straightforward English language construct using the conjunction "and".
+This involves combining all the selected triplet labels into a sentence by starting with "Realistic Image of".
+For instance, if the triplet labels are <man, on, horse>, <house, behind, horse>, and <man, in front, house>, the generated prompt becomes “Realistic Image of man on horse and house behind horse and man in front of house." To increase exemplar diversity for replays, we use the Stable Diffusion model [49] to generate $\gamma$ number of images for the same text prompt.
+In practice, we set $\gamma=10$ over all the learning scenarios.
+Figure 4: Visualizations of example generated images given the text prompt from our RAS.
+Each row presents two examples of generated images.
+Given the input text prompt on the left, the generated image is displayed on the right.
+These visualizations provide qualitative validations into the capability of RAS to produce diverse and contextually relevant images based on our designed text prompts.
+We provide the visualization examples of some synthesized images along with the corresponding text prompts in Fig.
+4.
+From these examples, we found that the composed text prompts and the synthesized images are often of high quality and contextual coherence, thanks to the Stable Diffusion Model [49].
+Scene Graph Prediction on Synthesized Images.
+During replays, to train the model $M_{t+1}$ on $I^{\prime}_{j}$, we also need to predict their corresponding scene graph notations $G^{\prime}_{j}$ on $I^{\prime}_{j}$.
+As the frozen model snapshot $M_{t}$ at the end of task $t$ carries prior knowledge for SGG from the previous tasks, we use it to predict notations $G^{\prime}_{j}$ on $I^{\prime}_{j}$.
+These $G^{\prime}_{j}$ comprises object nodes $O^{\prime}_{j}$ with their respective classes $c^{\prime}_{j}$, along with object locations $b^{\prime}_{j}$.
+Additionally, it includes corresponding relationship nodes $R^{\prime}_{j}$ formed by triplets <$o^{\prime}_{s}$, $p^{\prime}_{k}$, $o^{\prime}_{j}$> representing subject, predicate, and object nodes, respectively.
+These generated notations $G^{\prime}_{j}$, along with $I^{\prime}_{j}$, serve to construct the exemplars, used for replays.
+5 Results
+SGTR[28]
+Methods Learning Scenario 1 (S1) Learning Scenario 2 (S2)
+[…]
+Joint 20.15 0 4.6 0 - - 12.64 0 9.84 0 - -
+Replay@100% 16.17 -12.24 3.32 -1.34 -1.77 -11.72 4.56 -4.13 4.56 -5.61 -1.045 -30.25
+Naive 1.33 -28.7 0.86 -1.74 -2.03 -60.67 0.51 -23.22 0.05 -11.31 -3.77 -62.34
+EWC[22] 1.89 -28.4 0.96 -1.72 -1.17 -52.45 0 -23.22 0 -11.31 -2.65 -50.12
+PackNet[38] 7.19 -25.67 1.35 -1.64 -1.03 -42.35 1.67 -22.77 0.9 -10.33 -1.4 -42.45
+Replay@10% 8.55 -22.21 4.33 -1.44 4.29 -38.35 1.81 -20.72 1.15 -9.64 -0.9 -40.67
+Replay@20% 9.25 -20.35 4.78 -1.42 3.21 -  31.98 2.57 -17.17 1.56 -8.07 -0.67 -38.27
+Ours* 10.78 -18.92 5.6 -1.39 2.3 -25.56 3.45 -10.23 2.75 -6.57 -0.54 -35.67
+TCNN[63]
+Methods Learning Scenario 1 (S1) Learning Scenario 2 (S2)
+[…]
+Joint 19.53 0 3.9 0 - - 4.3 0 3.7 0 - -
+Replay@100% 13.45 -8.83 3.6 -0.35 -1.5 -10.45 12.45 -4.13 3.2 -0.56 -2.1 -20.34
+Naive 0.98 -21.2 0.74 -1.35 -3.45 -43.87 0 -18.22 0.45 -2.67 -4.12 -53.12
+EWC[22] 2.36 -21.05 0.67 -1.34 -2.34 -39.89 0 -18.22 0.03 0 -3.77 -51.67
+PackNet[38] 3.2 -19.7 1.1 -1.13 -1.3 -32.45 1.1 -17.82 0.84 -1.97 -2.84 -40.34
+Replay@10% 5.67 -18.9 3.21 -1.05 1.45 -28.34 1.81 -16.72 1.03 -1.74 -1.4 -43.56
+Replay@20% 6.23 -17.45 3.5 -1.01 1.01 -24.32 2.37 -15.17 1.45 -1.53 -1.1 -38.56
+Ours* 7.8 -15.67 3.9 -0.95 0.5 -19.83 4.67 -11.31 2.2 -0.89 -0.97 -29.65
+Table 2: Results of CSEGG for various continual learning methods applied on the two SGG backbones (SGTR and TCNN) in Learning Scenarios 1 and 2.
+See Sec.
+3.2 for continual learning baselines.
+See Sec.
+3.3 for evaluation metrics.
+The higher the evaluation metrics, the better.
+The best are in bold.
+5.1 RAS outperforms all the CSEGG baselines in Scenarios 1 and 2
+[…]
+LVIS@Replay@10% 3.98 -1.54
+BLS@Replay@10% 4.34 -1.47
+RAS (ours) 5.6 -1.39
+Table 3: Results at Task 5 in Learning Scenario 1 when sampling techniques are applied to long-tailed distribution data.
+See Sec.
+3.2 for the introduction to techniques used for long-tailed distributions.
+We copy the results of our RAS from Tab.
+2 for easy comparisons.
+The best results are in bold.
+Figure 5: Results in Learning Scenario 3. See Sec. 3.3 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning methods. Line types denote the IoU thresholds.
+The results for Avg.
+R, F, mR, mF, FWT, and BWT in learning scenarios 1 (S1) and 2 (S2) are presented in Tab.
+2.
+Our observations align with established research in continual learning, especially in image classification and object detection: regardless of the SGG architectures, over both learning scenarios, Naive consistently performs the worst, showcasing significant catastrophic forgetting.
+Rehearsal-based methods, such as Replay@10% and Replay@20%, outperform techniques like EWC and PackNet.
+However, none of them surpass our RAS.
+Moreover, compared to Replay@20%, RAS achieves superior performance while requiring a smaller constant storage size, equivalent to storing 15% exemplary images from the training set in the Replay methods.
+This storage size is determined by the image dimensions of 800 by 600 pixels and the number of model parameters in RAS.
+It is also worth noting that RAS only requires storing the old model snapshot, the triplet labels in $U_{t}$, and the image generative model, eliminating privacy concerns associated with storing original images.
+In Learning Scenario 2 (S2), the task involves classifying both new objects and new relationships, significantly escalating the level of difficulty compared to S1.
+As evident from Tab.
+2, all CSEGG models, including Replay@100%, exhibit a decline in overall performance when compared to the upper bound Joint.
+Although our RAS achieves the leading performance among all the baselines, its performance is still far from Joint.
+This underscores the persistent challenge posed by S2 in the context of CSEGG.
+Future work should explore new approaches to address the complexities inherent in this scenario.
+To tackle the issue of imbalanced data distribution in real-world scenarios, we incorporate two established data sampling techniques (LVIS@Replay@10% and BLS@Replay@10%) into our experiment (Sec.
+3.2).
+The outcomes in learning scenario S1 are presented in Tab.
+3.
+We observed that their performance falls short of our RAS, underscoring the efficacy of RAS in addressing long-tailed distribution during generative replays.
+We also noticed that BLS@Replay@10% significantly outperforms LVIS@Replay@10%, contrary to findings in the classical SGG problem where BLS is considered more effective than LVIS [29].
+The performance difference may stem from variations in the number of replay instances between the two approaches after applying these data re-sampling methods to exemplar images in the memory buffer (Sec.
+3.2).
+This observation suggests that the original sampling methods designed for addressing long-tailed distributions in the classical SGG problem may not be as effective when applied to CSEGG.
+As the knowledge carried forward is important for the subsequent tasks, we also permuted the task sequences and explored their role in CSEGG performances.
+Aligning with the existing literature [54], we found a prominent effect of task sequences in CSEGG (Fig.
+S3 and Sec.
+0.A.4.1).
+To gain a qualitative understanding of CSEGG performances, we provide visualization results of the predicted scene graphs on example images over tasks for all the CSEGG baselines in Scenario 1 (Fig.
+S5 and Sec.
+0.A.5.1) and Scenario 2 (Fig.
+S6 and Sec.
+0.A.5.2).
+5.2 CSEGG Models Can Generalize in Unknown Scenes
+Fig.
+5 illustrates the generalization results for detecting unknown objects and classifying known relationships among these objects in Learning Scenario 3 (S3).
+In Fig.
+5 (a), an increasing trend in Gen Rbbox is observed with the increasing task number for all CSEGG methods, indicating improved generalization in detecting unknown objects.
+Notably, even with minimal training in Task 1, all CSEGG methods propose 23% reasonable object regions with IoU = 0.7, showcasing the SGTR model’s ability to generalize “objectness".
+As expected, with an increase in IoU threshold from 0.3 to 0.7, we found that Gen Rbbox decreases due to fewer bounding boxes being considered correct.
+Moreover, we also compared the generalization performance in object detection between Replay@10% and Naive.
+Contrary to previous observations in S1 and S2 (Tab.
+2), we found that Replay@10% show a decline in Gen Rbbox, possibly due to a fixed number of detected object bounding boxes output by CSEGG methods.
+Similarly, our RAS exhibits a reduced G Rbbox compared to Replay@10% and Naive, likely for the same underlying reason.
+In Fig.
+5 (b), Replay@10% outperforms Naive in Gen R when considering correctly detected unknown object locations, emphasizing that minimizing forgetting in continual learning enhances the SGTR model’s overall relationship generalization in unknown scene understanding.
+However, the performance of Replay@10% is still inferior to our RAS method; implying that our RAS is more proficient in generalizing to classify relationships among unknown objects.
+Interestingly, we also noted that even with minimal training in Task 1, all the CSEGG methods achieve 45% recall of known relationships among unknown objects, demonstrating the SGTR model’s ability to generalize “relationship".
+Visualization examples, when CSEGG models can generalize to recognize relationships, are presented in Fig.S7 and Sec.0.A.5.3.
+5.3 Ablation Studies on Our RAS Reveal Key Design Insights
+$\gamma$ Context LTD Triplet Avg.R $\uparrow$ F $\uparrow$ mR $\uparrow$ mF $\uparrow$
+A1 10 Multiple 8.23 -21.35 3.98 -1.98
+A2 10 Multiple 9.75 -20.65 4.12 -1.65
+A3 10 Single 7.75 -22.45 3.12 -2.48
+A4 2 Multiple 2.45 -27.43 0.45 -9.84
+A5 4 Multiple 5.67 -26.42 2.41 -3.26
+A6 8 Multiple 7.89 -22.42 3.89 -2.17
+Ours 10 Multiple 10.78 -18.92 5.6 -1.39
+Table 4: Ablation results of our RAS on learning scenario S1 reveals key design insights.
+This table presents the results of ablation studies conducted to identify key components of our method, as discussed in Sec.
+5.3.
+Results in Avg.R, F, mR, mF are reported after the last task in the scenario S1.
+We introduce our default method designs in Sec. 4. Here, we vary the components in our RAS to reveal key design insights.
+We propose a context checker in RAS.
+Here, we conduct an ablation by removing this module.
+Triplet labels are randomly selected and combined for text prompts.
+In A1 of Tab.
+4, we observe a performance decrease of approximately 2% across all evaluation metrics, compared with our RAS.
+This suggests that generating images adhering to real-world context rules is crucial for replays.
+The lower performance may be attributed to the challenge of generating good-quality out-of-context images for Stable Diffusion Models and the potential domain differences affecting the SGG model $M_{t}$ in predicting out-of-context SGG notations.
+The LTD sampling module in our RAS is designed to balance the distribution of head and tail triplet labels from $U_{t}$.
+Here, we remove the LTD sampling module and report the performance of the ablated method in A2 of Tab.
+4.
+Compared to our RAS, we observe an absolute decrease of 1-2% across all metrics.
+Notably, the relative decrease is more pronounced in mR and mF than Avg.R and F.
+As mR and mF indicate mean Recall and mean Forgetfulness over both tail and head classes, the larger drops in these metrics suggest that the absence of LTD sampling significantly hinders the SGG model’s ability to predict tail classes from previous tasks.
+In our RAS, we employ multiple triplet labels to construct text prompts for image generation.
+In contrast to single triplet labels, our approach yields rich text descriptions of complex scenes, allowing the SGG model to capture intricate relationships among multiple objects in the same scene.
+Additionally, using multiple triplets is more efficient in rehearsing, as it enables the model to practice predicting multiple triplets simultaneously within the same number of synthesized images.
+Indeed, when we replace multiple triplet labels with single triplet labels for text prompts, we note a decrease of approximately 3% across all metrics (compare A3 with ours in Tab.
+4).
+Lastly, we investigate the impact of generating $\gamma$ images using the same text prompt in RAS.
+We vary $\gamma$ from 2 to 8 and present their corresponding model performances in A4-6 of Tab.
+4.
+As expected, the SGG model’s performance improves with higher $\gamma$, indicating that generating more images increases sample diversity for rehearsals and hence, it enhances CSEGG performance.
+We set $\gamma=10$ as the default; however, in practical scenarios with ample compute resources, dynamically synthesizing more images could further improve CSEGG performances.
+This also highlights the advantage of our RAS, as it can dynamically generate numerous images for replays without expanding its storage usage.
+6 Discussion
+In the dynamic world, adapting scene graph generation (SGG) models to new objects and relationships poses challenges.
+Despite progress in SGG and continual learning, there is still a gap in understanding Continual Scene Graph Generation (CSEGG).
+We address this by operationalizing CSEGG, and introducing benchmarks, datasets, and evaluation protocols.
+Our study explores three learning scenarios, analyzing continual object detection and relationship classification in long-tailed class-incremental settings for CSEGG baselines.
+Our findings show that integrating sampling methods with CSEGG baselines to address long-tailed distributions moderately eliminates forgetting; however, a large performance gap between current CSEGG baselines and the joint training upper bound persists.
+To address CSEGG challenges, we propose RAS, a Replays via Analysis by Synthesis method.
+RAS parses previous task scene graphs into triplet labels for diverse in-context scene graph reconstruction.
+Based on these re-compositional context-congruent scene graphs, RAS synthesizes images with Stable Diffusion models for replays.
+Unlike other image replay methods, RAS stores only triplet labels and the model snapshot, maintaining constant memory usage and preserving privacy.
+Extensive experiments demonstrate our RAS’s superior performance over current CSEGG baselines in knowledge transfers and reducing forgetting.
+Interestingly, our RAS model is also capable of generalizing to classify known relationships among unfamiliar objects.
+We discuss ethical concerns in the Sec.0.A.6.
+7 Acknowledgement
+This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-025), its NRFF award NRF-NRFF15-2023-0001, Mengmi Zhang’s Startup Grant from Agency for Science, Technology, and Research (A*STAR), and Early Career Investigatorship from Center for Frontier AI Research (CFAR), A*STAR.
+The authors declare that they have no competing interests.
+The authors would like to thank Stan Weixian Lei, Difei Gao, and Mike Shou for their feedback and suggestions.
+Appendix 0.A Appendix
+0.A.1 Introduction to Three Learning Scenarios
+Within this section, we present more details of three learning scenarios and their practical applications.
+Figure S1: Label distribution in each task in each learning scenario is presented.
 In scenario S1 (a) and scenario S3 (c), we use different colors to denote different tasks.
@@
 See the legend for triplet sizes.
-See Fig.
-S1,S2,S3 in Sec.
-A.3 for the full statistics of S1-3.
-3.1 Learning Scenarios
-We re-organize and divide the data from Visual Genome (Krishna et al., 2017) to cater to three continual learning scenarios below in our CSEGG dataset.
-The dataset contains 108,249 images, 150 objects, and 50 relationship categories.
-We follow the standard image splits for training, validation, and test sets (Xu et al., 2017).
-Scenario 1 (S1): Relationship Incremental Learning.
-While existing continual object detection literature focuses on incrementally learning object attributes (Mai et al., 2021; Wang et al., 2022b; Cha et al., 2021; Wang et al., 2021b; Shieh et al., 2020; Menezes et al., 2023), incremental relationship classifications are equally important as it provides a deeper and more holistic understanding of the interactions and connections between objects within a scene.
-See Sec.
-A.1.1 for a concrete example application in medical imaging.
+0.A.1.1 Scenario 1 (S1): Relationship Incremental Learning
+While existing continual object detection literature focuses on incrementally learning object attributes [37, 61, 8, 59, 51, 39], incremental relationship classifications are equally important as it provides a deeper and more holistic understanding of the interactions and connections between objects within a scene.
 To uncover contextual information and go beyond studies of object attributes, we introduce this scenario where new relationship predicates $p_{k}$ are incrementally added in each task (Fig.
-2S1).
+2 S1).
 There are 5 tasks in S1.
-To simulate the naturalistic settings where the frequency of relationship distribution is often long-tailed, we randomly and uniformly sample relationship classes from head, body and tail categories in Visual Genome (Krishna et al., 2017), and form a set of 10 relationship classes for each task.
+To simulate the naturalistic settings where the frequency of relationship distribution is often long-tailed, we randomly and uniformly sample relationship classes from head, body and tail categories in Visual Genome [23], and form a set of 10 relationship classes for each task.
 Thus, the relationships within a task are long-tailed; and the number of relationships from the head categories of each task is of the same scale.
-Different from continual object recognition and detection tasks, it is difficult to find unique images, in which ground truth scene graphs only contain relationships belonging to one specific task.
 To tackle this issue, we allow CSEGG models to see the same images over tasks, but the relationship labels are only provided in their given task (see Sec.
-A.1.5 for the design motivation).
+0.A.1.4 for the design motivation).
 The same reasoning applies in S2 and S3.
 Example relationship classes from each task and their distributions are provided in Fig.
-3(a).
-Scenario 2 (S2): Scene Incremental Learning.
+S1.
+Here, we provide a concrete example application of Scenario 1 in medical imaging.
+Within medical imaging, an agent must acquire the ability to detect cancerous cells within primary tumors, like colon adenocarcinoma.
+Subsequently, it must extend this proficiency to identifying the same cell types within metastatic growths that manifest in different bodily regions, such as lymph nodes or the liver.
+In this instance, the identical cancer cell disseminates to fresh organs or tissues, progressively establishing new relationships with other cells over the course of time.
+0.A.1.2 Scenario 2 (S2): Scene Incremental Learning
 To simulate the real-world scenario when there are demands for detecting new objects and new relationships over time in old and new scenes, we introduce this learning scenario where new objects $O_{i}$ and new relationship predicates $p_{k}$ are incrementally introduced over tasks (Fig.
-2S2).
-See Sec.
-A.1.2 for two real-world example applications in robot collaborations on construction sites and video surveillance systems.
-To select the object and relationship classes from the original Visual Genome (Krishna et al., 2017) for S2, we have two design motivations in mind.
+2 S2).
+To select the object and relationship classes from the original Visual Genome [23] for S2, we have two design motivations in mind.
 First, in real-world applications, such as robotic navigation, robots might have already learned common relationships and objects in one environment.
@@
 Same as S1, both the object class and relationship class distributions are still long-tailed within a task (Fig.
-3(b)).
-Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning.
-We, as humans, have no problem at all recognizing the relationships of unknown objects with other nearby objects, even though we do not know the class labels of the unknown objects.
-This scenario is designed to investigate whether the CSEGG models can generalize as well as humans.
-See Sec.
-A.1.3 for two real-world applications in the deep sea and space explorations for autonomous navigation systems.
-Specifically, there are 4 tasks in total with each task containing 30 object classes and 35 relationship classes.
-In each subsequent task, the CSEGG models are trained to continuously learn to detect 30 more object classes and learn to classify the same set of 35 relationships among these objects.
-The class selection criteria for each task follow the same as S1, where the selections occur uniformly over head, body, and tail classes.
-Example object classes and their label distributions for each task are provided in Fig.
-3(c).
-Different from S1 and S2, a standalone generalization test set is curated, where the objects are unknown and their classes do not overlap with any object classes in the training set but the relationships among these unknown objects are common to the training set of every task.
-The CSEGG models trained after every task are tested on the same generalization test sets.
-3.2 Scene Graph Generation Backbone
-(a) Schematic of SGTR (Li et al., 2022b)
-(b) Overview of three continual learning baselines
-Figure 4: Introduction to backbone SGG models and continual learning baselines.
-We use Scene graph Generation TRansformer (SGTR) (Li et al., 2022b) as the backbone SGG model (Sec.
-3.2).
-SGTR consists of four modules indicated by each blue box.
-Arrows indicate the signal flows among modules.
-(b) Three continual learning baselines are listed: EWC (Kirkpatrick et al., 2017), Replay (Rolnick et al., 2019), Naive (Sec.
-3.3) and PackNet (Mallya & Lazebnik, 2018)(Sec.
-3.3).
-$\theta^{*}_{A}$ denotes the optimal network parameters after learning on task A.
-The arrows in colors indicate the shifts of network weights in the parameter space when learning Task B for different baselines.
-We use the state-of-the-art one-stage Scene graph Generation TRansformer (SGTR) (Li et al., 2022b) and the traditional two-stage SGG model (Xu et al., 2017), named as CNN-SGG.
-CNN-SGG detects objects with Faster-RCNN(Girshick, 2015) backbone and infers their relationships separately via Iterative message passing (IMP)(Xu et al., 2017).
-For simplicity and consistency, we focus discussions solely on SGTR in the main text.
-For CNN-SGG, see Sec.
-A.2.2 for the introduction, training and implementation details over all three learning scenarios.
-We observed consistent relative CSEGG performance among all continual learning methods across both backbones (Fig.
-S15 and Sec.
-A.5).
-Different from CNN-SGG, SGTR formulates the task as a bipartite graph construction problem.
-Briefly, we introduce how SGTR works (Fig.
-4a).
-Given a scene image $I_{i}$, SGTR utilizes a 2D-CNN followed by a transformer-based encoder to extract image features.
-These features are further incorporated into a transformer-based decoder to predict object and subject nodes $O_{i}$.
-After that, the object-aware predicate nodes $R_{i}$ are formed based on both image features and object node features.
-Finally, a bipartite graph $G_{i}$ is constructed to collectively represent the scene with object $o_{i}$ and predicate nodes $r_{k}$, where the correspondence between these nodes is established based on the Hungarian matching algorithm (Kuhn, 1955).
-All experimental results are based on the average over 3 runs.
-We adapt the public source codes from (Li et al., 2022b) and (Wang et al., 2021b) for implementations of the continual learning algorithms on SGTR.
-We use the hyper-parameters provided in the code as the default values.
-See Sec.
-A.2.1 for training and implementation details of SGTR.
-All source code and data are publicly available here.
-Techniques for Learning in Long-tailed Distribution.
-As shown in Fig.
-3 and Sec.
-3.1, the data in the training set of every task in all learning scenarios is long-tailed.
-We adopt the three existing techniques to alleviate the problem of imbalanced data distribution during training.
-LVIS(Gupta et al., 2019) is an image-level over-sampling strategy.
-The number of image repeats depends on the object classes with minimal frequency over the entire dataset.
-Bi-level sampling (BLS) (Li et al., 2021) combines image-level oversampling and instance-level undersampling.
-LVIS is used for image-level oversampling.
-At instance levels, a drop-out rate is applied.
-The more frequent an instance belonging to the common classes, the higher the dropout rate.
-The two-level data re-sampling achieves an effective trade-off between the head and tail classes.
-Equalized Focal Loss (EFL) (Li et al., 2022a) is an effective loss function, re-balancing the loss contribution of head and tail classes independently according to their imbalance degrees.
-During the training of SGTR, EFL is enabled all the time, regardless of the learning scenarios or continual learning algorithms.
-3.3 Continual Learning Algorithms
-We benchmark the continual learning baselines below (Fig.4b).
-Naive (lower bound) with the SGTR model as the backbone, is trained on each task in sequence without any measures to prevent catastrophic forgetting.
-EWC(Kirkpatrick et al., 2017) is a weight-regularization method, where the weights of the network are regularized in the parameter space, based on their “importance" to the previous tasks.
-PackNet(Mallya & Lazebnik, 2018) is a parameter-isolation method, where it iteratively prunes and pre-trains the network parameters so that it can sequentially pack multiple tasks within one network.
-Replay(Rolnick et al., 2019) includes a memory buffer with the capacity of storing $M$ percentages of images in the entire dataset as well as their corresponding ground truth object and predicate notations depending on the task at each learning scenario.
-We vary $M=$ 10%, 20%, and 100%.
-We randomly select images present in the current task and add them to the buffer.
-As the image may contain multiple object/relationship ground truths, the number of replays on ground truth notations in each task might vary.
-See Fig.
-S1(c)(d), Fig.
-S2(c)(d) and Fig.
-S3(c)(d) in Sec.
-A.3.4 where we report the number of ground truth notations stored in the memory buffer for each task in each learning scenario.
-However, these variations do not interfere with fair comparisons across replay methods with different memory buffer capacities.
-We also introduce Joint Training — an upper bound where the SGG model is trained on the entire CSEGG dataset.
-To alleviate the problem of long-tailed data distribution, we also introduce BLS and LVIS in replay methods (Replay + BLS and Replay + LVIS).
-In the data re-balancing replay, the data in the memory buffer of Replay gets first re-sampled and then mixed with the data from the current task for training.
-As LVIS is an image-level up-sampling technique, the images containing the tail classes in the memory buffer are over-sampled.
-In contrast, BLS adds an extra instance-level under-sampling technique, which under-samples instances of head classes on the replay images.
-With these two approaches, the exact number of ground truth notations per image in the memory buffer varies over tasks in each learning scenario.
-We report these numbers in the Fig.
-S4 in Sec.
-A.3.
-The total number of instance replays for Replay+LVIS is slightly larger than the ones for Replay+BLS.
-We show the results of this effect in Sec.
-4.2.
-3.4 Evaluation Metrics
-Same as existing SGG works (Xu et al., 2017; Li et al., 2022b), we adopt the evaluation metric recall@K (R@K) on the predicted scene graphs $G$.
-As CSEGG is long-tailed, we further report the results in mean recall (mR@K) over head, body, and tail relationship classes in Sec.
-A.4.2.
-For consistency, we provide the R@K results in the main text, while we provide and analyze results in mR@K in Fig.
-S8 and Fig.
-S9 in Sec.
-A.4.2.
-Figure 5: Results of average recall and forgetting over tasks in Learning Scenarios 1 (a and b) and 2 (c and d) based on the SGTR backbone as the SGG model.
-See Sec.
-3.3 for introduction to continual learning baselines.
-See Sec.
-3.4 for explanations about evaluation metrics.
-X-axis indicates the task numbers.
-The higher Avg.R@20 and F@20, the better.
-See Fig.
-S5,S6 in Sec.
-A.4.1 for results of Avg.R@K, F@K, and mR@K, when K$=50,100$.
-See Sec.
-A.4.3 for results on FWT@K and BWT@K.
-See Sec.
-A.5 for the results based on the CNN-SGG backbone (Sec.3.2).
-To assess the catastrophic forgetting of CSEGG models, we define Forgetfullness (F@K) as the difference in R@K on $D_{t=1}$ between the CSEGG models trained at task $t$ and task 1.
-An ideal CSEGG model could maintain the same $R@K$ on $D_{t=1}$ over tasks; thus, $F=0$ for all tasks.
-The more negative F is, the more severe in forgetting an model gets.
-To assess the overall recall of CSEGG models over tasks, we also report the continual average recall (Avg.
-R@K).
-Avg.
-R@K is computed as the average recall on all the data at the previous and current tasks $D_{i}$, where $i\in\{1,2,...,t\}$.
-To assess whether the knowledge at previous tasks facilitates learning the new task and whether the knowledge at new tasks enhances the performances at older tasks, we introduce Forward Transfer (FWT@K) (Lin et al., 2022) and Backward Transfer (BWT@K)(Lopez-Paz & Ranzato, 2017).
-See Sec.
-A.2.3 and Sec.
-A.4.3 for its definitions, results, and analysis.
-[…]
-4 Results
-4.1 Continual Scene Graph Generation Remains a Great Challenge.
-We present Avg.
-R@20, F@20, FWT@20, and BWT@20 results for learning scenario 1 (S1) in Fig.
-5 (a)(b) and Fig.S11(a)(b).
-Notably, all continual learning baselines start from a similar Avg.R@20 in Task 1 and their performance drops over subsequent tasks.
-This implies that catastrophic forgetting about learned relationships occurs when the CSEGG models learn new relationships.
-Among all the methods, the naive method takes no measures of preventing catastrophic forgetting, resulting in the largest drop in Avg.R@20 and F@20.
-In contrast, a replay method with all the old data to rehearse in the current task (Replay(100%)) yields the least forgetting and maintains a high Avg.R@20.
-Surprisingly, even though Replay(100%), as an upper bound, replays all the data in the current and previous tasks, there is still a drop in performance.
-This could possibly be due to the long-tailed data distribution in the memory buffer, which makes the rehearsal of tail classes even less frequent in new tasks, and thus, deteriorate the recall performances of tail classes.
-We also compared EWC versus the replay methods.
-Though EWC outperforms the naive baseline in earlier tasks, it fails in longer task sequences.
-Different from EWC, Replay with 10% still achieves a higher Avg.R@20 score of 8.55% and a higher F@20 of -22.21%.
-This aligns with the existing continual learning literature that replay methods are more effective than weight-regularization methods in eliminating catastrophic forgetting (Lesort et al., 2019).
-PackNet is a parameter isolation method.
-While PackNet outperforms EWC, its performance is inferior to that of Replay(10%).
-As expected, we also compare the replay methods with different memory buffer sizes.
-Replaying more old data helps CSEGG performances.
-Joint training demonstrates superior performance over all the tasks in Learning Scenario 1 as seen in Fig 5.
-This aligns with the existing continual learning literature that joint training is a superior upper bound than Replay(100%).
-As the knowledge carried forward is important for the subsequent tasks, we also permuted the task sequences and explored their role in CSEGG performances.
-Aligning with the existing literature (Singh et al., 2022), we found a prominent effect of task sequences in CSEGG (Fig.
-S13 and Sec.
-A.4.4).
-Learning scenario 2 (S2) approximates the real-world CSEGG setting where there are constantly new demands in detecting new objects and new relationships simultaneously.
-The results of S2 in Avg.R@20 and F@20 are provided in Fig.
-5(c)(d).
-Compared with S1, the overall Avg.R@20 and F@20 drop more significantly over tasks.
-For example, even with 20% memory buffer size, the replay method only achieves Avg.
-R@20 score of 6.57% and F@20 of -17.17% in Task 2.
-This suggests that the real-world CSEGG remains a challenging task and there still exists a large performance gap for state-of-the-art CSEGG methods.
-Moreover, we also made an interesting observation that Replay (100%) outperforms the upper bound of the joint training in the first task of Scenario 2.
-This performance difference could be attributed to the presence of long-tailed data distribution across tasks, with the first task containing more tailed classes than head classes.
-This is in contrast to the task splits in Scenario 1 where both head and tail classes are uniformly sampled for every task.
-Consequently, joint training struggles in the first task due to sub-optimal performance in tailed classes.
-To gain a qualitative understanding of CSEGG performances, we provide visualization results of the predicted scene graphs on example images over tasks for all the CSEGG baselines in Scenario 1 (Fig.
-S18 and Sec.
-A.6.1) and Scenario 2 (Fig.
-S19 and Sec.
-A.6.2).
-Model Avg.
-R@20 $\uparrow$ F@20 $\uparrow$ T1 T5 T1 T5 Replay(10%) 28.7 8.55 0 -22.21 Replay(10%)+LVIS 28.7 14.38 0 -15.39 Replay(10%)+BLS 28.7 9.56 0 -22.4 Replay(100%) 28.7 16.17 0 -12.24 Table 1: Results at Task 1 and 5 in Learning Scenario 1 when sampling techniques on long-tailed distribution are applied.
-See Sec.
-3.2 for the introduction to techniques used for long-tailed distributions.
-The best results are in bold.
-[Uncaptioned image] Figure 6: Generalization results in Learning Scenario 3.
-See Sec.
-3.4 for evaluation metrics.
-The higher the values, the better.
-Line colors indicate continual learning baselines.
-Line types denote the IoU thresholds for determining correctly predicted bounding box locations.
-4.2 Accounting for Long-tailed Distributions Enhances CSEGG
-Due to the imbalanced data distribution in the real world, long-tailed distribution remains a unique challenge for CSEGG.
-Here, we introduce two data sampling techniques (LVIS and BLS) to counter-balance the long-tailed data distributions in the memory buffers as well as the feed-forward training tasks (Sec.
-3.2 and Sec.
-3.3).
-We report the results of Replay(10%)+BLS and Replay(10%)+LVIS in learning scenario S1.
-In Tab.
-1, both long-tailed methods with replay(10%) outperform naive replays by an average margin of 3.42% in Avg.R@20 and 3.35% in F@20.
-This implies that data re-sampling techniques enhance general continual learning performances in long-tailed incremental learning settings.
-Indeed, we made the same observations after splitting the classes from each task into tail, body, and head classes and reporting their mR@K in the Fig.
-S10 in A.4.2.
-Interestingly, we see that Replay(10%)+BLS underperforms Replay(10%)+LVIS by 4.32% in Avg.R@20 and 6.42% in F@20.
-This contradicts the findings that BLS is more effective than LVIS in the classical SGG problem (Li et al., 2021).
-The performance discrepancy could be due to the difference in the number of replay instances in both approaches after these two data re-sampling methods are applied to the memory buffer (see Sec.
-3.3).
-This emphasizes that the long-tailed learning methods explored in the SGG problem may not be effective in CSEGG.
-We need to explore new long-tailed learning methods specifically for CSEGG.
-4.3 CSEGG Improves Generalization in Unknown Scene Understanding
-[…]
-Notably, we also found that the SGTR model is very good at generalizing to classify relationships Fig.
-6 (b).
-Even in Task 1, both the naive method and the Replay(10%) achieve 45% recall of known relationships among unknown objects in the generalization test set.
-As the CSEGG models continuously learn to detect more new objects and classify their relationships in subsequent tasks, their relationship generalization ability among unknown objects saturates around Task 3.
-See Fig.
-S20 in Sec.
-A.6.3 for visualization examples.
-5 Discussion
-In the dynamic world, the incremental introduction of new objects and new relationships in scenes presents significant challenges for scene graph generation (SGG) models to effectively adapt and generalize without forgetting previously acquired semantic knowledge.
-However, despite the progress made in SGG and continual learning research, there remains a lack of comprehensive investigations specifically targeting the unique challenges of Continual Scene Graph Generation (CSEGG).
-To close this research gap, we take the initial steps of operationalizing CSEGG and introducing benchmarks, datasets, and evaluation protocols.
-Our study delves into three distinct learning scenarios, thoroughly examining the interplay between continual object detection and continual relationship classification for existing CSEGG methods under long-tailed class-incremental settings.
-Our experimental results reveal intriguing insights.
-First, applying standard continual learning approaches combined with long-tailed techniques to SGG models yields moderate improvements.
-However, a notable performance gap persists between current CSEGG methods and the joint training upper bound.
-Second, we investigated the model’s generalization ability and found that the models are capable of generalizing to classify known relationships involving unfamiliar objects.
-Third, we compared the CSEGG performance of the traditional CNN-based and the transformer-based SGG models as backbones.
-We observed consistent relative CSEGG performance across all continual learning methods using both backbones, with CNN-SGG models underperforming SGTR-based ones.
-Moving forward, there are several key avenues for future research.
-Our current endeavors focus on learning CSEGG problems from static images in Independent and Identically Distributed (i.d.d) manner, diverging from human learning from video streams.
-Future research can look into CSEGG problems on video streams.
-Our plans also involve expanding continual learning baselines and integrating more long-tailed distribution sampling techniques.
-Furthermore, we aim to construct a synthetic SGG dataset to systematically quantify the aspects of SGG that influence continual learning performance under controlled conditions.
-Although the CSEGG method holds promise for many downstream applications like monitoring systems, medical imaging, and autonomous navigation, we should also be aware of its misuse in privacy, data biases, fairness, security concerns, and misinterpretation.
-We invite the research community to join us in maintaining and updating the safe use of CSEGG benchmarks, thereby fostering its advancements in this field.
-Ethics Statement
-The development and deployment of Scene Graph Generation (SGG) technology present potential negative societal impacts that warrant careful consideration (Li et al., 2022b).
-Firstly, privacy concerns arise as SGG may inadvertently capture sensitive information from images, potentially violating privacy rights and raising surveillance issues.
-Secondly, bias and fairness challenges persist, as SGG algorithms can perpetuate biases present in training data, leading to discriminatory outcomes that reinforce societal inequalities.
-Misinterpretation and misclassification by SGG algorithms could result in misinformation and incorrect actions, impacting decision-making.
-The risk of manipulation and misuse of SGG-generated scene representations for malicious purposes is also a concern.
-For example, attackers might manipulate scene graphs to deceive systems or disrupt applications that rely on scene understanding.
-Reproducibility Statement
-We are committed to ensuring the reproducibility and transparency of our research.
-In accordance with the guidelines set forth by ICLR 2024, we provide detailed information to facilitate the replication of our experiments and results.
-  1. 1.
-    Code Availability: All code used for our experiments is available at here.
-  2. 2.
-    Data Availability: Any publicly accessible datasets used in our research are specified in the paper, along with their sources and access information.
-  3. 3.
-    Experimental Details: We have documented the specific details of our experiments, including hyper-parameters, model architectures, and pre-processing steps, to enable others to replicate our results.
-We are dedicated to supporting the scientific community in replicating and building upon our work.
-We welcome feedback and collaboration to ensure the robustness and reliability of our research findings.
-Acknowledgement
-This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-025), its NRFF award NRF-NRFF15-2023-0001, Mengmi Zhang’s Startup Grant from Agency for Science, Technology, and Research (A*STAR), and Early Career Investigatorship from Center for Frontier AI Research (CFAR), A*STAR.
-The authors declare that they have no competing interests.
-The authors would like to thank Stan Weixian Lei, Difei Gao, and Mike Shou for their feedback and suggestions.
-Appendix A Appendix
-A.1 Motivation for Learning Scenarios
-Within this section, we present practical applications that demonstrate the real-world relevance of investigating CSEGG across all three learning scenarios.
-A.1.1 Scenario 1 (S1): Relationship Incremental Learning
-Within medical imaging, an agent must acquire the ability to detect cancerous cells within primary tumors, like colon adenocarcinoma.
-Subsequently, it must extend this proficiency to identifying the same cell types within metastatic growths that manifest in different bodily regions, such as lymph nodes or the liver.
-In this instance, the identical cancer cell disseminates to fresh organs or tissues, progressively establishing new relationships with other cells over the course of time.
-A.1.2 Scenario 2 (S2): Scene Incremental Learning
-The CSEGG model’s capacity to incorporate new objects and new relationships while retaining existing knowledge finds pivotal application in security contexts.
+S1).
+Next, we provide two real-world example applications in robot collaborations on construction sites and video surveillance systems.
+The CSEGG model’s capacity to incorporate new objects and new relationships while retaining existing knowledge finds pivotal application in video surveillance contexts.
 Consider a company developing video-based security systems for indoor environments, capturing prevalent indoor objects and relationships.
@@
 This ensures that the robots can effectively communicate and collaborate while comprehending the evolving scene and tasks, optimizing their construction efficiency and accuracy.
-A.1.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
+0.A.1.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
+We, as humans, have no problem at all recognizing the relationships of unknown objects with other nearby objects, even though we do not know the class labels of the unknown objects.
+This scenario is designed to investigate whether the CSEGG models can generalize as well as humans.
+Specifically, there are 4 tasks in total with each task containing 30 object classes and 35 relationship classes.
+In each subsequent task, the CSEGG models are trained to continuously learn to detect 30 more object classes and learn to classify the same set of 35 relationships among these objects.
+The class selection criteria for each task follow the same as S1, where the selections occur uniformly over head, body, and tail classes.
+Example object classes and their label distributions for each task are provided in Fig.
+S1.
+Different from S1 and S2, a standalone generalization test set is curated, where the objects are unknown and their classes do not overlap with any object classes in the training set but the relationships among these unknown objects are common to the training set of every task.
+The CSEGG models trained after every task are tested on the same generalization test sets.
+Here, we provide two real-world applications of Scenario 3 in the deep sea and space explorations for autonomous navigation systems.
 A prime example is the ongoing research on deep sea exploration for autonomous navigation systems, where undiscovered flora and fauna reside beneath the ocean’s surface.
 Encountering new and unidentified species becomes manageable through SGG’s ability to understand spatial relations.
-The robot discerns the object’s proximity or orientation even without precise identification of the spieces, enhancing its autonomous navigation ability.
+The robot discerns the object’s proximity or orientation even without precise identification of the species, enhancing its autonomous navigation ability.
 Likewise, in deep space exploration, SGG aids in recognizing spatial relationships with previously unseen space debris, aiding in path-planning.
 In essence, SGG’s relationship generalization empowers robots to navigate and plan routes in unfamiliar terrains, such as deep sea and deep space, where novel encounters demand adaptable responses.
-A.1.4 Profound Impacts Beyond the Problem of CSEGG
-We wish to emphasize that the challenge of continuous Scene Graph Generation (SGG) applied to real-world images exhibits two distinctive attributes.
-Techniques developed to address these attributes possess the potential to be widely applicable across various domains and subsequent tasks.
-Firstly, SGG in a continuous learning context frequently deals with data stemming from distributions that exhibit long-tailed characteristics.
-The methodologies formulated within the realm of Continual SGG can be adapted to address broader long-tailed continuous learning issues, such as classifying bird sounds in ecological studies.
-Secondly, the process of continual SGG demands the ability to engage in reasoning and integrate knowledge over time.
-As an illustration, in the context of retail inventory management, merely learning to identify durians (a tropical fruit uncommon in US supermarkets) is insufficient.
-The model must also cultivate the capability to integrate this new information into its existing knowledge database, ensuring that this new product is positioned alongside other fruits within the grocery store.
-A.1.5 Design Motivations for Overlapping Images and Non-overlapping Labels across Tasks
+0.A.1.4 Design Motivations for Overlapping Images and Non-overlapping Labels across Tasks
 The design of such images and label splits over tasks aligns with human learning scenarios where a parent teaches the baby to recognize different toys and objects in the bedroom.
@@
 In the future, we will expand our studies to cases where the SGG models learn from non-overlapping sets of training images for each task.
-A.2 Implementation Details
-A.2.1 SGTR
+0.A.2 Implementation and Training Details
+Figure S2: Introduction to backbone SGG models and continual learning baselines.
+We use Scene graph Generation TRansformer (SGTR) [28] as the backbone SGG model (Sec.
+3.2).
+SGTR consists of four modules indicated by each blue box.
+Arrows indicate the signal flows among modules.
+(b) Four continual learning baselines are listed: EWC [22], Replay [48], Naive (Sec.
+3.2) and PackNet [38](Sec.
+3.2).
+$\theta^{*}_{A}$ denotes the optimal network parameters after learning on task A.
+The arrows in colors indicate the shifts of network weights in the parameter space when learning Task B for different baselines.
+0.A.2.1 SGTR Backbone
+For SGTR as depicted in (Fig.
+S2 (a)), the approach uniquely formulates the task as a bipartite graph construction problem.
+Starting with a scene image ($I_{i}$), SGTR utilizes a 2D-CNN and transformer-based encoder to extract image features.
+These features are then incorporated into a transformer-based decoder, predicting object and subject nodes ($O_{i}$).
+Predicate nodes ($R_{i}$) are formed based on both image features and object node features, and a bipartite graph ($G_{i}$) is constructed to represent the scene collectively.
+The correspondence between object nodes ($o_{i}$) and predicate nodes ($r_{k}$) is established using the Hungarian matching algorithm [24].
+Experimental results are based on the average over three runs, and the implementation leverages public source codes from [28] and [59] with default hyperparameters.
 The SGTR is trained in two stages in a supervised manner.
-In stage 1, only object detection losses in DETR is (Carion et al., 2020) applied on $O_{i}$.
-In stage 2, only predicate entity prediction loss is applied on $R_{i}$, which can be further decomposed into L1 and GIOU losses for object/subject/predicate localization (Rezatofighi et al., 2019) and cross-entropy loss for object/subject/predicate classification.
-In learning scenario S1, we skip Stage 1, and directly load pre-trained weights of DETR for object detection on the entire training set of Visual Genome (Li et al., 2022b).
+In stage 1, only object detection losses in DETR is [6] applied on $O_{i}$.
+In stage 2, only predicate entity prediction loss is applied on $R_{i}$, which can be further decomposed into L1 and GIOU losses for object/subject/predicate localization [46] and cross-entropy loss for object/subject/predicate classification.
+In learning scenario S1, we skip Stage 1, and directly load pre-trained weights of DETR for object detection on the entire training set of Visual Genome [28].
 In stage 2 of S1, we freeze the feature extractor, and fine-tune the rest parts of SGTR for predicate entity predictions.
 As only relationship classes are incrementally introduced in S1, we freeze the entire weights of DETR for detecting all the objects in the scene over tasks.
-However, empirical results suggest that fine-tuning transformer-based encoders in DETR helps downstream predicate predictions (Li et al., 2022b).
+However, empirical results suggest that fine-tuning transformer-based encoders in DETR helps downstream predicate predictions [28].
 Even with fine-tuning DETR in S1, we verify that there is minimal forgetting of detecting all objects in the scene over tasks (see Sec.
-S14 and Sec.
-A.4.5).
+S4 and Sec.
+0.A.4.2).
 Thus, the forgetting observed in S1 could only be attributed to incremental relationship learning.
-In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the feature extractor, pre-trained on ImageNet (Deng et al., 2009), and fine-tune the entire DETR (Carion et al., 2020) over all the tasks.
+In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the feature extractor, pre-trained on ImageNet [12], and fine-tune the entire DETR [6] over all the tasks.
 Stage 2 of S2 and S3 is the same as S1.
@@
 All models are trained on 4 A5000 GPUs.
-A.2.2 CNN-SGG Backbone
-Given a scene image $I_{i}$, CNN-SGG utilizes Faster-RCNN(Girshick, 2015) to generate a set of object proposals.
+0.A.2.2 TCNN Backbone
+As for TCNN, it employs Faster-RCNN [17] to generate object proposals from a scene image ($I_{k}$).
+The model extracts visual features for nodes and edges from these proposals.
+Through message passing, both edge and node GRUs output a structured scene graph.
+Experimental results are based on the average over three runs.
+Given a scene image $I_{i}$, TCNN utilizes Faster-RCNN[17] to generate a set of object proposals.
 The model subsequently extracts visual features of nodes and edges from the set of object proposals.
 Finally, both edge and node GRUs output a structured scene graph via message passing.
-The CNN-SGG is trained in two stages in a supervised manner.
-In stage 1, only object detection losses in Faster-RCNN(Ren et al., 2015) are applied on $O_{i}$.
+The TCNN is trained in two stages in a supervised manner.
+In stage 1, only object detection losses in Faster-RCNN[45] are applied on $O_{i}$.
 We use the cross entropy loss for the object class and $L1$ loss for the bounding box offsets.
-In stage 2, the visual feature extractor (VGG-16(Simonyan & Zisserman, 2014) pre-trained on ImageNet (Deng et al., 2009)) and GRUs layers are trained to predict the final object classes, bounding boxes, and relationship predicates using cross-entropy loss and $L1$ loss.
-In Learning Scenario 1 (S1), similar to the implementation details of SGTR in Sec.3.2, we skip Stage 1, and directly load pre-trained weights of Faster-RCNN for object detection on the entire training set of Visual Genome (Li et al., 2022b).
+In stage 2, the visual feature extractor (VGG-16[53] pre-trained on ImageNet [12]) and GRUs layers are trained to predict the final object classes, bounding boxes, and relationship predicates using cross-entropy loss and $L1$ loss.
+In Learning Scenario 1 (S1), similar to the implementation details of SGTR in Sec.
+0.A.2.1, we skip Stage 1, and directly load pre-trained weights of Faster-RCNN for object detection on the entire training set of Visual Genome [28].
 In stage 2 of S1, we load the pre-trained weights of the visual feature extractor (pre-trained on ImageNet) and fine-tune the rest parts of the model.
 In stage 2 of S1, we load the pre-trained weights of visual feature extractor (pre-trained on ImageNet) and fine-tune the rest parts of the model.
-In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the Faster-RCNN, pre-trained on ImageNet Deng et al.
-(2009), and fine-tune it over all the tasks.
+In Stage 1 of S2 and S3 where object classes are also incrementally introduced over tasks, we load weights of the Faster-RCNN, pre-trained on ImageNet [12], and fine-tune it over all the tasks.
 Stage 2 of S2 and S3 follows the same training regimes as Stage 2 of S1.
@@
 All models are trained on 4 A5000 GPUs.
-A.2.3 Definitions of BWT and FWT Evaluation Metrics
-To assess the influence that learning a task $t$ has on the performance of any previous tasks in CSEGG models, we also report Backward Transfer (BWT@K)(Lopez-Paz & Ranzato, 2017).
+0.A.3 Evaluation Metrics
+To assess the catastrophic forgetting of CSEGG models, we define Forgetfullness (F@K) as the difference in R@K on $D_{t=1}$ between the CSEGG models trained at task $t$ and task 1.
+An ideal CSEGG model could maintain the same $R@K$ on $D_{t=1}$ over tasks; thus, $F=0$ for all tasks.
+The more negative F is, the more severe in forgetting an model gets.
+To assess the overall recall of CSEGG models over tasks, we also report the continual average recall (Avg.
+R@K).
+Avg.
+R@K is computed as the average recall on all the data at the previous and current tasks $D_{i}$, where $i\in\{1,2,...,t\}$.
+To assess whether the knowledge at previous tasks facilitates learning the new task and whether the knowledge at new tasks enhances the performances at older tasks, we introduce Forward Transfer (FWT@K) [34] and Backward Transfer (BWT@K)[35].
 BWT@K is defined as $BWT@K=\frac{1}{T-1}\sum_{i=1}^{T-1}R@K_{T,i}\>-\>R@K_{i,i}$, where $T$ denotes the total number of tasks in a learning scenario and $R@K_{i,j}$ denotes the continual learning model trained after task $i$ and tested in task $j$.
-To assess the influence of previous tasks on the current task $t$ in CSEGG models, we also report Forward Transfer (FWT@K) (Lin et al., 2022).
 FWT is defined as $FWT@K=\frac{1}{T-1}\sum_{i=2}^{T}R@K_{i,i}\>-\>\overline{b@K}_{i,i}$, where $\overline{b@K}_{i,i}$ is the R@K for an independent model with random initialization trained in task $i$ and tested in task $i$.
-A.3 Data Statistics
-In this section, we provide various types of data statistics for all three learning scenarios.
-Specifically, we present statistics regarding the number of images, objects, and relationships involved in each task of each learning scenario.
-Additionally, we include data statistics pertaining to replay buffers of sizes 10%, 20%, and 100%.
-A.3.1 Scenario 1 (S1): Relationship Incremental Learning
-In this scenario, new relationship predicates $p_{k}$ are incrementally added in each task.
-The Learning Scenario 1 (S1) comprises five tasks, with each task consisting of 10 mutually exclusive relationships.
-Across all tasks, there is a common set of 150 objects present.
-Fig.
-S1 shows comprehensive data statistics for stages 1 and 2 over all the tasks.
-As mentioned in Sec.
-3.2, we skip the Stage 1 training in S1 and directly load pre-trained weights of DETR (Zhu et al.
-(2020)) for object detection on the entire training set of Visual Genome.
-Fig.
-S1(a) shows the distribution of objects in the entire training set of Visual Genome.
-Fig.
-S1(b) displays the distribution of relationships during Training Stage 2 for each task in the Learning Scenario 1 (S1).
-Each task consists of 10 relationships that are mutually exclusive.
-Notably, a long tail pattern is observed in the distribution of relationships for each task.
-The legend of Fig.
-S1(b) indicates that there is a relatively uniform number of images present for each task in Stage 2 of the Learning Scenario 1 (S1).
-The test sets for each task in S1 exhibit the distributions of the number of images and relationships that closely align with the distributions of the training sets depicted in Fig. S1(b).
-A.3.2 Scenario 2 (S2): Scene Incremental Learning
-In this learning scenario, new objects $O_{i}$ and new relationship predicates $p_{k}$ are incrementally introduced over tasks.
-The Learning Scenario 2 (S2) comprises of 2 tasks with the first task containing 100 objects and 40 relationships and the second task containing 25 objects and 5 relationships.
-Fig.
-S2 shows comprehensive data statistics pertaining to Learning Scenario 2 (S2).
-As mentioned in Sec.
-3.2, both the Stage 1 and Stage 2 training are present in S2.
-Fig.
-S2(a) presents the distribution of objects during Stage 1 for each task in the S2.
-The first task consists of 100 objects and the second task consists of 25 objects which are mutually exclusive.
-Notably, a long tail pattern is observed in the distribution of objects for each task.
-The legend of Fig.
-S2(a) indicates that there is a uniform number of images present for each task in Stage 1 of the S2.
-Fig.
-S2(b) displays the distribution of relationships during Stage 2 for each task in the S2.
-The first task consists of 40 relationships and the second task consists of 5 relationships that are mutually exclusive.
-Notably, a long tail pattern is observed in the distribution of relationships for each task.
-The legend of Fig.
-S2(b) highlights that in Stage 2 of the S2, Task 1 has a noticeably higher number of images compared to Task 2.
-However, the number of relationship notations per class from Task 2 is higher than Task 1, since relationships from head classes are assigned to Task 2 (see Sec.
-3.1).
-The test sets for each task in S2 exhibit similar distributions in the number of images, objects, and relationships to the distributions of the training sets depicted in Fig. S2(a)(b).
-A.3.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
-In this learning scenario, there are a total of four tasks, with each task encompassing 30 distinct objects and 35 common relationships.
-Fig.
-S3 shows comprehensive data statistics pertaining to Learning Scenario 3 (S3).
-As mentioned in Sec.
-3.2, both Stage 1 and Stage 2 training are present in S3.
-Fig.
-S3(a) presents the distribution of objects during Stage 1 for each task in the S3.
-Each task consists of 30 objects which are mutually exclusive.
-Notably, a long tail pattern is observed in the distribution of objects for each task.
-The legend of Fig.
-S3(a) indicates that there is a relatively uniform number of images present for each task in Stage 1 of the S3.
-Fig.
-S3(b) displays the distribution of relationships during Stage 2 for each task in the S3.
-Each task consists of 35 relationships which are common over all the tasks.
-Notably, a long tail pattern is observed in the distribution of relationships for each task.
-As mentioned in Sec.
-3.1 of Scenario 3 (S3), a standalone generalization test set is curated for S3.
-This test set consists of 1942 images and features a completely different set of objects compared to those present in the training data.
-However, the test set maintains the same set of relationships as observed in the training data.
-A.3.4 Replay Buffers
-In our benchmarking process, we evaluate multiple replay baselines using a memory buffer with a fixed capacity to store a certain percentage (M) of images from the entire training dataset, along with their corresponding ground truth object and relationship notations specific to each learning scenario.
-We vary the value of M to be 10%, 20%, and 100%.
-As there could be multiple ground truth objects and relationships per image, we provide statistics on the number of ground truth notations stored in the memory buffer for each task in all the learning scenarios in this section.
-Specifically, we report the histograms for ground truth notations in the replay buffer for scenario 1 (Fig.
-S1), scenario 2 (Fig.
-S2), and scenario 3 (Fig.
-S3).
-In general, over all three learning scenarios, we observe a long-tail distribution of notations per object or relationship class within each task.
-At stage 1, as we fix the memory capacity over tasks, the number of images stored in the memory buffer remains constant over tasks in most cases, except for memory capacity $M=100\%$.
-This is to maximize the diversity of information for replays.
-Note that although the number of images remain constant in the memory buffer, the number of images allocated to each task decreases given a fixed memory buffer capacity; and hence, we saw a decrease in the number of notations from previous tasks.
-At stage 2, the images stored in stage 1 are carried over.
-However, as some images might not contain task-relevant relationship classes, the number of images used for training at stage 2 might vary over tasks.
-The notations stored in the memory buffer also follow the long-tail distributions.
-In learning scenario 1, we reported the histograms of ground truth notations after two sampling techniques, LVIS and BLS, are applied in the memory buffers (Fig.
-S4).
-Both LVIS and BLS resampling methods aim to address the long-tail distribution issues by either reducing the prominence of head classes or over-sampling the object instances from the tail classes.
-Indeed, by comparing the histograms in Fig.
-S4(a)(b), we noticed such changes in histograms for tail and head classes.
-A similar trend in the class distribution per task emerges when applying long-tail distribution techniques to the exemplar dataset, as indicated in Fig.S4(c)(d).
-Figure S1: Data Statistics for Learning Scenario 1 (S1): (a) Distribution of objects in the entire training set of Visual Genome during Stage 1.
-(b) Distribution of relationships during Stage 2 for each task in S1.
-(c) and (d) show distributions of relationships in the memory buffer for Replay 10% and Replay 100 % during Stage 2 for each task in S1.
-See the legend for the color codes of each task.
-The numbers in brackets in the legend in (b-d) denote the number of training images in the particular task.
-Zoom in to the figure to get the exact labels and the frequency associated with them.
-Figure S2: Data Statistics for Learning Scenario 2 (S2): (a) Distribution of objects during Stage 1 for each task in S2.
-(b) Distribution of relationships during Stage 2 for each task in S2.(c) and (d) show distributions of objects in the memory buffer for Replay 10% and Replay 20% during Stage 1 for each task in S2.
-(e) and (f) show distributions of relationships in the memory buffer for Replay 10% and Replay 20% during Stage 2 for each task in S2.
-See the legend for the color codes of each task.
-The numbers in brackets in the legend in (a-f) denote the number of training images in the particular task.
-Zoom in to the figure to get the exact labels and the frequency associated with them.
-Figure S3: Data Statistics for Learning Scenario 3 (S3): (a) Distribution of objects during Stage 1 for each task in S3.
-(b) Distribution of relationships during Stage 2 for each task in S3.
-(c) Distribution of objects in the memory buffer for Replay 10% during Stage 1 for each task in S3.
-(d) Distribution of relationships in the memory buffer for Replay 10% during Stage 2 for each task in S3.
-See the legend for the color codes of each task.
-The numbers in brackets in the legend in (a-d) denote the number of training images in the particular task.
-Zoom in to the figure to get the exact labels and the frequency associated with them.
-Figure S4: Data Statistics for Learning Scenario 1 (S1) when sampling techniques on long-tailed distribution are applied: (a) Distribution of relationships during Stage 2 for each task in S1 when LVIS is applied.
-(b) Distribution of relationships in the memory buffer for Replay 10% during Stage 2 for each task in S1 when LVIS is applied.(c) Distribution of relationships during Stage 2 for each task in S1 when BLS is applied.
-(d) Distribution of relationships in the memory buffer for Replay 10% during Stage 2 for each task in S1 when BLS is applied.
-The numbers in brackets in the legend in (a-d) denote the number of training images in the particular task.
-Zoom in to the figure to get the exact labels and the frequency associated with them.
-A.4 More Quantitative Result Analysis on Continual SGTR Methods
-A.4.1 Results for K = 50, 100
-Fig.
-S5 and S6 present the results for Avg.R@K and F@K with K values of 50 and 100, focusing on the Learning Scenarios 1 and 2.
-These results align with the findings discussed in Sec.
-4.1, which considered Avg.R@20 and F@20.
-Notably, when K is increased to 50 and 100, both Avg.R and F exhibit higher absolute values compared to K=20 but the trends remain consistent.
-Fig.
-S7 presents the results of Gen.R@K for K=50,100 for Learning Scenario 3.
-Similar findings in Sec.
-4.3 can be made here as well.
-Figure S5: Results in AvgR@K and F@K over tasks in Learning Scenario 1 (S1) when K=50,100. See Fig. 5 for design conventions.
-Figure S6: Results in AvgR@K and F@K over tasks in Learning Scenario 2 (S2) when K=50,100. See Fig. 5 for design conventions.
-Figure S7: Results in Gen R@K over tasks in Learning Scenario 3 (S3) when K=50,100. See Fig. 6 for design conventions.
-A.4.2 Mean Recall Results
-Mean Recall@K (mR@K) calculates Recall@K for each relationship category independently and then reports their mean.
-This is a more fair evaluation metric when dealing with long tail distributions as it averages out the recall over all the relationship classes.
-In this section, we define mF@K and Avg.mR@K based on mR as described in Sec.
-3.4.
-We report the results of mF@K and Avg.mR@K for learning scenarios 1 and 2 in Fig.
-S8 and Fig.
-S9.
-Note that the evaluations with mF@K and Avg.mR@K are not applicable in learning scenario 3.
-Though the absolute values of mF@K and Avg.mR@K are generally smaller than the values of F@K and Avg.R@K in Fig.
-5, the same observations made in Sec.
-4.1 can be applied here as well.
-Figure S8: Results in Avg. mR@K and mF@K over tasks in Learning Scenario 1 (S1) when K=20,50,100. See Fig. 5 for design conventions.
-Figure S9: Results in Avg. mR@K and mF@K over tasks in Learning Scenario 2 (S2) when K=20,50,100. See Fig. 5 for design conventions.
-Figure S10: Results in Avg.
-mR@K and mF@K over tasks in Learning Scenario 1 (S1) when sampling techniques on long-tailed distribution are applied..
-See Sec.
-3.2 for the introduction to techniques used for long-tailed distributions.
-See Tab.
-1 for design conventions.
-A.4.3 Result Analysis on BWT@K and FWT@K
-We provide the results of BWT@K and FWT@K in Fig.
-S11 for Learning Scenario 1.
-Consistent with the results of Avg.R@20 and F@20, we observe that among all methods, the naive method has the lowest FWT@20 and BWT@20 implying catastrophic forgetting.
-Once again, EWC outperforms the naive baseline by having higher FWT@20 and BWT@20 though the difference is small as it fails in longer task sequences.
-Although all the baselines have negative BWT@20 indicating forgetting, Replay(10%) yields a higher BWT@20 than PackNet, EWC and naive baseline pointing to a reduced level of forgetting.
-Replay(100%) exhibits the greatest BWT@20, suggesting the significance of revisiting previous samples to prevent forgetting and facilitate backward knowledge transfer.
-Similar to BWT@20, most of the baselines have negative FWT@20 implying that knowledge acquired from prior tasks slightly interferes with the learning process of the current task.
-In contrast, Replay(10%) shows positive FWT@20 indicating that knowledge acquired from prior tasks enhances the learning process of the current task.
-Surprisingly, while Replay(100%) excels in Avg.R@20, F@20, and BWT@20, Replay(10%) achieves a superior FWT@20.
-This suggests that Replay(100%) might not grasp the current task as effectively as Replay(10%), as it preserves the knowledge from previous tasks by replaying more old samples and forgetting less at the sacrifice of slow learning about the current task.
-Similar trends are observed and the same analysis can be applied in Learning Scenario 2 (Fig.
-S12).
-Figure S11: Results of FWT (a) and BWT (b) over baselines in Learning Scenario 1 based on the SGTR backbone.
-See Sec.
-3.3 for introduction to continual learning baselines.
-See Sec.
-3.4 for explanations about evaluation metrics.
-The x-axis indicates the various baselines.
-The higher FWT@20 and BWT@20, the better.
-Figure S12: Results of FWT (a) and BWT (b) over baselines in Learning Scenario 2 based on the SGTR backbone.
-See Sec.
-3.3 for introduction to continual learning baselines.
-See Sec.
-3.4 for explanations about evaluation metrics.
-The x-axis indicates the various baselines.
-The higher FWT@20 and BWT@20, the better.
-A.4.4 Results for Different Task Sequences
-Recent work (Singh et al., 2022) has highlighted the consistent and substantial curriculum effects in class-incremental learning and continual visual question-answering tasks.
+In learning scenario S3, we evaluate CSEGG models on their abilities to generalize to detect unknown objects and classify known relationships on these objects, in the standalone generalization test set over all tasks.
+To benchmark these, we introduce two evaluation metrics: the recall of the predicted bounding boxes on unknown objects (Gen Rbbox@K) and the recall of the predicted graph $G_{i}$ (Gen R@K).
+As the CSEGG models have never been taught to classify unknown objects, we discard the class labels of the bounding boxes and only evaluate the predicted box locations with Gen Rbbox@K.
+To evaluate whether the predicted box location is correct, we apply a hard threshold of Intersection over Union (IoU) between the predicted bounding box locations and the ground truth.
+Any predicted bounding boxes with their IoU values above the hard threshold are deemed to be correct.
+We vary IoU thresholds from 0.3, 0.5, and 0.7.
+To assess whether the CSEGG model generalizes to detect known relationships over unknown objects, we evaluate the recall Gen R@K of the predicted relationships $r_{k}$ only on correctly predicted bounding boxes.
+0.A.4 More Quantitative Result Analysis on Continual SGTR Methods
+0.A.4.1 Results for Different Task Sequences
+Recent work [54] has highlighted the consistent and substantial curriculum effects in class-incremental learning and continual visual question-answering tasks.
 Inspired by these findings, we conducted experiments to assess the impact of the curricula in the context of CSEGG.
@@
 Notably, in Fig.
-S13(d), a large difference in Avg.R@20 between Order 1 and Order 3 emerges for the Replay (100%) baseline.
+S3(d), a large difference in Avg.R@20 between Order 1 and Order 3 emerges for the Replay (100%) baseline.
 Similarly, Fig.
-S13(a) reveals a substantial Avg.R@20 disparity between Order 2 and Order 3 for the Naive baseline.
+S3(a) reveals a substantial Avg.R@20 disparity between Order 2 and Order 3 for the Naive baseline.
 This trend extends to F@20, as depicted in Fig.
-S13(e)(f)(g)(h).
+S3(e)(f)(g)(h).
 These insights collectively affirm the significance of the curricula within CSEGG.
@@
 (b),(f) denotes Avg.R@K and F@K for EWC baseline.
-(c),(g) denotes Avg.R@K and F@K for Replay(10%) baseline.
-(d),(h) denotes Avg.R@K and F@K for Replay(100%) baseline.
+(c),(g) denotes Avg.R@K and F@K for Replay@10% baseline.
+(d),(h) denotes Avg.R@K and F@K for Replay@100% baseline.
 See Sec.
@@
 The higher F, Avg.R@20, the better.
-A.4.5 Minimal Forgetting in DETR
-To validate the impact of fine-tuning the DETR model in training Stage 2 of learning scenario S1 on relationship predicate predictions and to ensure minimal forgetting occurs in object detection (Sec 3.2), we compare the mean Average Precision (mAP) for object detection on the entire test set of Visual Genome between the pre-trained DETR checkpoint from the paper (Li et al., 2022b) and the DETR models after fine-tuning on each task of S1.
+0.A.4.2 Minimal Forgetting in DETR
+To validate the impact of fine-tuning the DETR model in training Stage 2 of learning scenario S1 on relationship predicate predictions and to ensure minimal forgetting occurs in object detection (Sec 0.A.2), we compare the mean Average Precision (mAP) for object detection on the entire test set of Visual Genome between the pre-trained DETR checkpoint from the paper [28] and the DETR models after fine-tuning on each task of S1.
 As shown in Fig.
-S14, the results indicate a slight decrease of 0.4 in mAP from the pre-trained checkpoint to the DETR models over 5 tasks.
+S4, the results indicate a slight decrease of 0.4 in mAP from the pre-trained checkpoint to the DETR models over 5 tasks.
 This study provides evidence that fine-tuning DETR in S1 has negligible effects on forgetting.
@@
 Figure S4: mAP performance on the entire Visual Genome Test Set.
-We used a pre-trained DETR checkpoint (Li et al., 2022b) as well as a naive baseline of SGTR continually trained on each task from S1, evaluated them on the entire Visual Genome’s test set, and reported their mAP performances.
+We used a pre-trained DETR checkpoint [28] as well as a naive baseline of SGTR continually trained on each task from S1, evaluated them on the entire Visual Genome’s test set, and reported their mAP performances.
 We observed a minimal decrease in mAP across all tasks.
-A.5 Comparision between CNN-SGG Backbone and the SGTR Model
-[…]
-Figure S15: Results of average recall and forgetting over tasks in Learning Scenario 1 (a and b) for CNN-SGG, when K-20.
-See Sec.
-3.3 for introduction to continual learning baselines.
-See Sec.
-3.4 for explanations about evaluation metrics.
-X-axis indicates the task numbers.
-The higher Avg.R@20 and F@20, the better.
-Figure S16: Results of average recall and forgetting over tasks in Learning Scenario 2 (a and b) for CNN-SGG (K=20).
-See Sec.
-3.3 for introduction to continual learning baselines.
-See Sec.
-3.4 for explanations about evaluation metrics.
-X-axis indicates the task numbers.
-The higher Avg.R@20 and F@50, the better.
-Fig.
-S17 provides the generalization results in detecting unknown objects and classifying known relationships among these objects in Learning Scenario 3 (S3) for CNN-SGG backbone.
-As elaborated in Sec.
-4.3, an intriguing trend emerges: the CSEGG models demonstrate an increasing ability to identify unknown objects as the number of tasks rises.
-Furthermore, similar to our observations in Sec.
-4.3, the CNN-SGG model exhibits proficiency in identifying known relationships between unfamiliar objects.
-However, this capability reaches a plateau as the task count escalates.
-Figure S17: Generalization results in Learning Scenario 3 for CNN-SGG (K=20).
-See Sec.
-3.4 for evaluation metrics.
-The higher the values, the better.
-Line colors indicate continual learning baselines.
-Line types denote the IoU thresholds for determining correctly predicted bounding box locations.
-A.6 Visualization Examples for All Learning Scenarios for Continual SGTR-based Models
-In this section, we present visualization examples from each learning scenario to showcase the performance of the three continual SGTR-based models, namely Replay10%, EWC, and Naive, in three learning scenarios.
-A.6.1 Learning Scenario 1 (S1)
+0.A.5 Visualization Examples for All Learning Scenarios for Continual SGTR-based Models
+In this section, we present visualization examples from each learning scenario to showcase the performance of the three continual SGTR-based models, namely Replay@10%, EWC, and Naive, in three learning scenarios.
+0.A.5.1 Learning Scenario 1 (S1)
 Figure S5: Visualization example for Learning Scenario 1 (S1).
 The leftmost column in the figure displays the ground truth bounding boxes and scene graphs for each task in Learning Scenario 1 (S1).
-The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay 10%, Naive, and EWC).
+The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay@10%, Naive, and EWC).
 In all the scene graphs, red boxes indicate objects, while green boxes represent relationships.
@@
 The leftmost column shows the ground truth bounding boxes and scene graphs in each task of Learning Scenario 2 (S2).
-The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay 10%, Naive, and EWC).
+The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Replay@10%, Naive, and EWC).
 In all the scene graphs, red boxes indicate objects, while green boxes represent relationships.
 As explained in Fig.
-S18 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet.
+S5 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet.
 The time arrow on the left side of the figure demonstrates that the model is exposed to new data over time, with new objects and relationships incrementally added, as described in Sec.
@@
 From Fig.S5 we observe that, in Task 1, the ground truth scene graph contains triplets of "on" relationship: "plate on table" and "hair on women".
-After training on task 1, all three models (Replay 10%, EWC, Naive) can accurately predict these triplets of "on" relationship.
+After training on task 1, all three models (Replay@10%, EWC, Naive) can accurately predict these triplets of "on" relationship.
 In Task 2, triplets of "has" relationship are introduced: "plate has food" and "women has hair".
-After training on task 2 data, the Replay 10% model successfully remembers the triplets of "on" relationship ("plate on table", "hair on women") from Task 1 and predicts "women has hair".
+After training on task 2 data, the Replay@10% model successfully remembers the triplets of "on" relationship ("plate on table", "hair on women") from Task 1 and predicts "women has hair".
 The Naive model forgets the triplets of "on" relationship and only predicts "women has hair".
@@
 In Task 3, triplet of "at" relationship is introduced: "women at table".
-After training on task 3 data, the Replay 10% model remembers the previous triplets ("plate on table", "women has hair", "hair on women") and predicts "women at table".
+After training on task 3 data, the Replay@10% model remembers the previous triplets ("plate on table", "women has hair", "hair on women") and predicts "women at table".
 The Naive model forgets the previous triplets and only predicts "women at table".
@@
 In Task 4, triplets related to "of" relationship are introduced: "hand of women", "neck of women", "arm of women", and "head of women".
-After training on task 4 data, the Replay 10% model remembers the triplets related to "on" and "has" relationships ("plate on table", "women has hair", "hair on women") from previous tasks but forgets the "at" relationship triplet (“women at table”).
+After training on task 4 data, the Replay@10% model remembers the triplets related to "on" and "has" relationships ("plate on table", "women has hair", "hair on women") from previous tasks but forgets the "at" relationship triplet (“women at table”).
 It only predicts "head of women" from the triplets introduced in Task 4.
@@
 In Task 5, triplet belonging to the "behind" relationship is introduced: "trunk behind women".
-After training on task 5 data, the Replay 10% model forgets the triplets related to "on" relationship ("plate on table", "hair on women") and only remembers the triplets related to "has" and "of" relationships ("women has hair", "head of women") learned from the previous task.
+After training on task 5 data, the Replay@10% model forgets the triplets related to "on" relationship ("plate on table", "hair on women") and only remembers the triplets related to "has" and "of" relationships ("women has hair", "head of women") learned from the previous task.
 It is not able to predict "trunk behind women".
@@
 The EWC model also fails to remember any triplets from the previous task and only predicts "trunk behind women".
-A.6.2 Learning Scenario 2 (S2)
-From Fig.S19 we observe that, in Task 1, the ground truth scene graph contains triplets: “man riding skateboard”, “man above skateboard”, and “shoe of skateboard”.
-After training on task 1 data, all three models (Replay 10%, Naive, EWC) predict “man riding skateboard” and “man above skateboard”.
+0.A.5.2 Learning Scenario 2 (S2)
+From Fig.S6, we observe that, in Task 1, the ground truth scene graph contains triplets: “man riding skateboard”, “man above skateboard”, and “shoe of skateboard”.
+After training on task 1 data, all three models (Replay@10%, Naive, EWC) predict “man riding skateboard” and “man above skateboard”.
 None of the models predict “shoe of skateboard”.
 In Task 2, new triplets introduced are: “leg of person”, “person wearing sock”, “person has head”, and “head of person”.
-After training on task 2 data, the Replay 10% model only remembers “man riding skateboard” from the previous task, forgetting “man above skateboard”.
-Moreover, Replay 10% model can only predict “leg of person” from the triplets introduced in task 2.
+After training on task 2 data, the Replay@10% model only remembers “man riding skateboard” from the previous task, forgetting “man above skateboard”.
+Moreover, Replay@10% model can only predict “leg of person” from the triplets introduced in task 2.
 The Naive model forgets all the triplets from task 1 ( “man riding skateboard”, “man above skateboard”) and only predicts “leg of person” from the triplets introduced in task 2.
@@
 The leftmost column shows the standalone ground truth bounding boxes and scene graphs in the generalization test set regardless of which task it is in Learning Scenario 3 (S3).
-The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Naive, Replay 10%).
+The remaining columns, from left to right, represent the bounding boxes and scene graphs generated by each baseline model (Naive, Replay@10%).
 Similar to Fig.
-S18, and S19, the red boxes in all scene graphs indicate objects , while green boxes represent relationships.
+S5, and S6, the red boxes in all scene graphs indicate objects , while green boxes represent relationships.
 As explained in Fig.
-S18 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet.
+S5 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet.
 The time arrow on the left side of the figure demonstrates that the model is exposed to new objects over time as described in Sec.
@@
 Concretely, unk_1 refers to "mountain"; unk_2 is "house"; and unk_3 is "sheep".
-A.6.3 Learning Scenario 3 (S3)
+0.A.5.3 Learning Scenario 3 (S3)
 Fig.
-S20 illustrates the performance of the Replay 10% and Naive models in locating unknown objects and recognizing the relationships between these objects and other nearby unknown objects.
+S7 illustrates the performance of the Replay@10% and Naive models in locating unknown objects and recognizing the relationships between these objects and other nearby unknown objects.
 The ground truth in Fig.
-S20 consists of three unknown objects: "mountain", "sheep", and "house", along with three relationships: "near", "behind", and "infront of" (between "mountain" and "house"), and a "near" relationship (between "sheep" and "house").
+S7 consists of three unknown objects: "mountain", "sheep", and "house", along with three relationships: "near", "behind", and "infront of" (between "mountain" and "house"), and a "near" relationship (between "sheep" and "house").
 After training on Task 1 data, the Naive model can accurately locate the objects "mountain" and "house".
@@
 Even after training on Task 4 data, the Naive model continues to locate all three objects: "mountain", "house", and "sheep".
-In contrast, the Replay 10% model, after training on Task 1 data, can only locate "mountain" and "house".
-This remains the same even after training on Task 2 and Task 3 data, where the Replay 10% model can still only locate the objects "mountain" and "house".
-However, after training on Task 4 data, the Replay 10% model is able to locate all the objects: "mountain", "house", and "sheep".
+In contrast, the Replay@10% model, after training on Task 1 data, can only locate "mountain" and "house".
+This remains the same even after training on Task 2 and Task 3 data, where the Replay@10% model can still only locate the objects "mountain" and "house".
+However, after training on Task 4 data, the Replay@10% model is able to locate all the objects: "mountain", "house", and "sheep".
 Regarding relationship generalization on unknown objects, the Naive model, after training on Task 1, can only predict the "near" relationship between the located objects "mountain" and "house" out of the three possible relationships.
@@
 After Task 4, the Naive model can predict the "behind" and "near" relationships between the located objects "mountain" and "house", as well as the "near" relationship between the located objects "sheep" and "house".
-In contrast, the Replay 10% model, after training on Task 1, can predict the "near" and “behind” relationships between the located objects "mountain" and "house".
+In contrast, the Replay@10% model, after training on Task 1, can predict the "near" and “behind” relationships between the located objects "mountain" and "house".
 After Task 2, it can also predict the "infront of" relationship between the located objects "mountain" and "house" along with “near” and “behind” relationships.
-Even after Task 3, the Replay 10% model is still able to predict the "near", "behind", and "infront of" relationships between the located objects "mountain" and "house".
-After Task 4, as it can now locate the object "sheep", the Replay 10% model can also predict the "near" relationship between the objects "house" and "sheep", in addition to the existing relationships between "mountain" and "house".
+Even after Task 3, the Replay@10% model is still able to predict the "near", "behind", and "infront of" relationships between the located objects "mountain" and "house".
+After Task 4, as it can now locate the object "sheep", the Replay@10% model can also predict the "near" relationship between the objects "house" and "sheep", in addition to the existing relationships between "mountain" and "house".
+0.A.6 Ethical Concerns
+The development and deployment of Scene Graph Generation (SGG) technology present potential negative societal impacts that warrant careful consideration [28].
+Firstly, privacy concerns arise as SGG may inadvertently capture sensitive information from images, potentially violating privacy rights and raising surveillance issues.
+Secondly, bias and fairness challenges persist, as SGG algorithms can perpetuate biases present in training data, leading to discriminatory outcomes that reinforce societal inequalities.
+Misinterpretation and misclassification by SGG algorithms could result in misinformation and incorrect actions, impacting decision-making.
+The risk of manipulation and misuse of SGG-generated scene representations for malicious purposes is also a concern.
+For example, attackers might manipulate scene graphs to deceive systems or disrupt applications that rely on scene understanding.

### v4

@@
-
+Adaptive Visual Scene Understanding: Incremental Scene Graph Generation
 Abstract
@@
 Our experimental results not only highlight the challenges of directly combining existing continual learning methods with SGG backbones but also demonstrate the effectiveness of our proposed approach, enhancing CSEGG efficiency while simultaneously preserving privacy and memory usage.
-All data and source code are publicly available at here.
-Keywords:
-Scene graph generation Continual learning In-context symbolic replay Long tail distribution Compositional scene graphs
+All data and source code are publicly available here.
 1 Introduction
@@
 Thus, as new objects are introduced to the scenes, the complexity of relationships among all the objects increases significantly in a non-linear fashion.
-(3) The long-tailed distribution in SGG can be attributed to the inherent characteristics of real-world scenes, where certain objects are more prevalent than others.
+(3) The long-tailed distribution in both objects and relationships in SGG can be attributed to the inherent characteristics of real-world scenes, where certain objects are more prevalent than others.
 Consequently, CSEGG requires the computational models to adapt continually to the evolving long-tailed distributions over different scenes.
@@
 These compositional scene graphs are then used for synthesizing scene images for replays.
-Due to its nature of symbolic replays, RAS doesn’t require the storage of original images, which often carry excessive and redundant details.
+Due to its nature of symbolic replays, RAS does not require the storage of original images, which often carry excessive and redundant details.
 This also ensures data privacy preservation and data efficiency.
@@
 Visual Phrase [59] stands as one of the earliest datasets in the field of visual phrase recognition and detection.
-Over time, various large-scale datasets have emerged to tackle the challenges of Scene Graph Generation (SGG) [21, 36, 23, 25, 32, 69, 65, 63, 71, 11, 31, 72].
-Among these, the Visual Genome dataset [23] has played a pioneering role by providing rich annotations of objects, attributes, and relationships in images.
+Over time, various large-scale datasets have emerged to tackle the challenges of Scene Graph Generation (SGG) on static images [23, 42, 25, 27, 37, 78, 74, 72, 80, 12, 35, 81].
+Subsequent works further extend the SGG to dynamic videos [22, 50, 56].
 Despite the significant contributions of these datasets to SGG, none focuses on continual learning in SGG.
-As the preliminary efforts towards CSEGG, we re-structure the Visual Genome dataset [23] and establish a novel and comprehensive CSEGG benchmark, where AI models are deployed to dynamic scenes where new objects and new relationships are introduced.
+As the preliminary efforts towards CSEGG, we start with fundamental and straightforward settings of SGG on static images.
+Among all the SGG datasets on static images, the Visual Genome dataset [25] has played a pioneering role by providing rich annotations of objects, attributes, and relationships in images.
+Thus, we re-structure the Visual Genome dataset [25] and establish a novel and comprehensive CSEGG benchmark, where AI models are deployed to dynamic scenes where new objects and new relationships are introduced.
 Scene Graph Generation (SGG) Models.
@@
 Our results reveal the limitations of these methods and highlight the challenges of our CSEGG learning protocols.
+S. #Tasks #Objs #Rels
+Eval.
+metrics
+SGG Backbone CL base. Kn. Unk.
+S1 5 150 (All) 10 per task F, R, mF, mR, FWT, BWT, Gen Rbbox, Gen R Transformer based (SGTR) CNN based (IMP) Joint, Naive, Replay M%, EWC, PackNet, RAS_GT
+Objs
+(bbox, labels)
+Rels
+S2 2
+Task 1: 100
+Task 2: 25
+Task 1: 40
+Task 2: 5
+None
+Rels and Objs
+(bbox, labels)
+S3 4 30 per task 35 per task
+Rels
+Objs
+(bbox, labels)
+Table 1: Overview of three CSEGG learning scenarios.
+This table summarizes the three learning scenarios (Column 1) in CSEGG, including the number of tasks, the number of object (#Objs) and relationship (#Rels) classes, the evaluation metrics, the SGG-Backbones used, and the continual learning (CL) baselines.
+The Kn.
+and Unk.
+columns provide information regarding what is known to the CSEGG models during training in that scenario and what is being incrementally learned by the models.
+Unknown information is being incrementally learned by the models.
+See Sec.
+3 for details.
 Continual Learning Methods.
@@
 The components in RAS facilitate memory-efficient training and preserve privacy while maintaining the scene diversity and scene context for SGG in dynamic environments.
-Scenarios #Tasks #Objs #Rels
-Eval.
-metrics
-SGG
-backbones
-CL
-baselines
-S1 5 All tasks: 150 Each task: 10 F, R, mF, mR, FWT, BWT, R_bbox, Gen_R Transformer based (SGTR) CNN based (TCNN) Joint, Naive, Replays[48], EWC[22], PackNet[38]
-S2 2
-Task 1: 100
-Task 2: 25
-Task 1: 40
-Task 2: 5
-S3 4 Each task: 30 All tasks: 35
-Table 1: Overview of three CSEGG learning scenarios.
-This table summarizes the three learning scenarios in CSEGG, including the number of tasks, the number of object (#Objs) and relationship (#Rels) classes, the evaluation metrics, the SGG-Backbones used, and the continual learning (CL) baselines.
-See Sec.
-3 for details.
+With the rise of pretrained vision-language models (VLMs) [49, 31, 83, 40], various SGG methods [11, 30] have been proposed to tackle open-vocabulary and zero-shot SGG challenges.
+However, these settings differ fundamentally from CSEGG, where we aim to simulate scenarios where the model encounters novel predicates or objects, unseen by any models, including LLMs or multi-modal models.
+Using pre-learned information from LLMs or frozen encoders conflicts with the continual learning setting we address in CSEGG.
 3 Continual ScenE Graph Generation Benchmark
@@
 Gen.) in Object Incre..
-In S1 and S2, example triplet labels in the training (solid line) and test sets (dash line) from each task are presented.
+In S1 and S2, example triplet labels in the training (solid line) and test sets (dotted line) from each task are presented.
 The training and test sets from the same task are color-coded.
+Blue color indicates task 1 and orange color indicates task 2.
 The new objects or relationships in each task are bold and underlined.
-In S3, one single test set (dashed gray box) is used for benchmarking the relationship generalization ability of object incre.
+In S3, one single test set (dotted gray box) is used for benchmarking the relationship generalization of object incre.
 learning models across all the tasks.
@@
 To uncover contextual information and go beyond studies of object detection and recognition, we introduce this scenario consisting of 5 tasks where 10 new relationship classes are incrementally added in every task (Fig.
-2, left; Tab.
+2, left; Fig.
+S1; Tab.
 1).
@@
 This scenario resembles a human learning scenario where a parent gradually teaches a baby to recognize new relationships among all objects in the same room, focusing on one new relationship at a time during continual learning.
-This scenario also has implications in medical imaging where identical cell types may form new relationships with nearby cells depending on the context.
-See Sec.
-0.A.1.1 for details.
+This scenario also has implications in medical imaging where identical cell types may form new relationships with nearby cells depending on the context (Sec.
+A.1.1).
 Scenario 2 (S2): Scene Incremental Learning.
 To simulate the real-world cases when there are demands for detecting new objects and new relationships from old to new scenes, we introduce this scenario where new objects and new relationships are incrementally introduced over tasks (Fig.
-2, middle; Tab.
+2, middle; Fig.
+S1; Tab.
 1).
@@
 See Sec.
-0.A.1.2 for details.
+A.1.2 for details.
 Scenario 3 (S3): Relationship Generalization.
@@
 All relationship classes are made known to all CSEGG models over all the tasks (Fig.
-2, right; Tab.
+2, right; Fig.
+S1; Tab.
 1).
@@
 See Sec.
-0.A.1.3 for details.
+A.1.3 for details.
 Data sampling and distributions.
@@
 However, in S2, only tail classes are sampled and added in subsequent tasks.
+The number of tasks in each scenario is experimentally determined to optimize the training data configuration, ensuring sufficient training samples in each task while maximizing the number of tasks.
+For detailed statistics, see Fig.
+S2 and Sec.
+A.2.
 3.2 Competitive CSEGG Baselines
@@
 See Sec.
-0.A.2 for implementation and training details of CSEGG baselines.
+A.3 for implementation and training details of CSEGG baselines.
 SGG Backbones.
@@
 Briefly, SGTR (Fig.
-S2(a)) uses a transformer-based architecture for image feature extraction and fusion.
+S3 left) uses a transformer-based architecture for image feature extraction and fusion.
 During training, [32] formulates SGG as a bipartite graph construction and matching problem.
 In contrast, TCNN detects objects with Faster-RCNN[18] backbone and infers their relationships separately via Iterative message passing [72].
-We use public source codes from [28] and [59] with default hyperparameters.
-Continual Learning Methods.
+We use implementations from [32] and [68] with default hyperparameters.
+Baselines.
 We include the following continual learning methods (Fig.
-S2(b)): (1) Naive (lower bound) is trained on each task in sequence without any measures to prevent catastrophic forgetting.
+S3 right): (1) Naive (lower bound) is trained on each task in sequence without any measures to prevent catastrophic forgetting.
 (2) EWC[24] is a weight-regularization method, where the weights of the network are regularized in the parameter space, based on their “importance" to the previous tasks.
@@
 (5) Joint Training is an upper bound where the SGG model is trained on the entire CSEGG dataset.
+(6) RAS_GT is a baseline in which we use the ground truth scene graph labels from each task to create replay buffers using an image generation model explained in detail in Sec.
+4.
+See Fig.
+S4 for schematics of CSEGG baselines.
+We provide mathematical formulations of these baselines in Sec.
+A.4.
 Sampling Methods to Handle Long-Tailed Distribution.
@@
 See Sec.
-0.A.3 for detailed definitions of these metrics.
-All experimental results are averaged over three runs.
+A.5 for details.
+All results are averaged over 3 runs.
 4 Replays via Analysis by Synthesis (RAS)
@@
 Our RAS belongs to the group of generative replay methods for continual learning.
+We create an exemplar set $E_{t}$ for replays.
+At task $t$, we jointly train the scene graph generation model $M_{t}$ on $E_{t}$ and the current training dataset $D_{t}$.
 However, different from the existing generative replay methods [61, 15, 16], our RAS leverages symbolic replays with state-of-the-art diffusion models.
@@
 3.
-Next, we focus on how RAS generates images and predicts SGG annotations on these images for replays.
+Next, we focus on how RAS creates $E_{t}$ containing the generated images and their SGG annotations on these images for replays.
 Image Generation.
@@
 To generate images $I^{\prime}_{j}$ for replays, our RAS feeds text prompts, which are formed by a set of chosen triplet labels and describe the diverse in-context scenes, into the state-of-the-art Stable Diffusion model [58].
-The small object on the table is most likely to be a spoon rather than an elephant.
-Context plays important roles in visual perceptions [5].
+As previous works suggests context plays important roles in visual perceptions [5].
 To generate text prompts describing context-congruent scenes, we employ a context checker.
@@
 The threshold for the agglomerative clustering algorithm is set to 0.6.
-As real-world images often contain complex scenes involving multiple triplets, we select any cluster with more than 3 triplet labels to create a text prompt for the image generation model.
+As real-world images often contain complex scenes involving multiple triplets, we select any cluster with more than 3 triplet labels to create a text prompt for image generation.
 In practical applications, conducting agglomerative clustering on all triplet labels in $U_{t}$ is computationally demanding, as it requires computing pairwise embedding similarities among all the triplet labels.
@@
 In practice, we set $\gamma=10$ over all the learning scenarios.
-Figure 4: Visualizations of example generated images given the text prompt from our RAS.
-Each row presents two examples of generated images.
-Given the input text prompt on the left, the generated image is displayed on the right.
-These visualizations provide qualitative validations into the capability of RAS to produce diverse and contextually relevant images based on our designed text prompts.
 We provide the visualization examples of some synthesized images along with the corresponding text prompts in Fig.
-4.
-From these examples, we found that the composed text prompts and the synthesized images are often of high quality and contextual coherence, thanks to the Stable Diffusion Model [49].
+S5.
+From these examples, we found that the composed text prompts and the synthesized images are often of high quality and contextual coherence.
 Scene Graph Prediction on Synthesized Images.
@@
 Additionally, it includes corresponding relationship nodes $R^{\prime}_{j}$ formed by triplets <$o^{\prime}_{s}$, $p^{\prime}_{k}$, $o^{\prime}_{j}$> representing subject, predicate, and object nodes, respectively.
-These generated notations $G^{\prime}_{j}$, along with $I^{\prime}_{j}$, serve to construct the exemplars, used for replays.
+These generated notations $G^{\prime}_{j}$, along with $I^{\prime}_{j}$, serve to construct the exemplars $E_{t}$, used for replays.
 5 Results
@@
 EWC[24] 1.89 -28.4 0.96 -1.72 -1.17 -52.45 0 -23.22 0 -11.31 -2.65 -50.12
+RAS_GT 5.78 -26.51 1.43 -1.54 -1.2 -44.27 0.98 -23.11 0.76 -10.86 -1.6 -43.25
 PackNet[44] 7.19 -25.67 1.35 -1.64 -1.03 -42.35 1.67 -22.77 0.9 -10.33 -1.4 -42.45
@@
 The best are in bold.
+* means the experiment is still running, we will report the results in the final version.
 5.1 RAS outperforms all the CSEGG baselines in Scenarios 1 and 2
@@
 See Sec.
-3.2 for the introduction to techniques used for long-tailed distributions.
+3.2 for the sampling techniques on long-tailed distributions.
 We copy the results of our RAS from Tab.
@@
 The best results are in bold.
-Figure 5: Results in Learning Scenario 3. See Sec. 3.3 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning methods. Line types denote the IoU thresholds.
+Figure 4: Results in Scenario 3. See Sec. 3.3 for evaluation metrics. The higher the values, the better. Line colors indicate continual learning methods. Line types denote the IoU thresholds.
 The results for Avg.
@@
 Our observations align with established research in continual learning, especially in image classification and object detection: regardless of the SGG architectures, over both learning scenarios, Naive consistently performs the worst, showcasing significant catastrophic forgetting.
-Rehearsal-based methods, such as Replay@10% and Replay@20%, outperform techniques like EWC and PackNet.
+Replay-based methods, such as Replay@10% and Replay@20%, outperform techniques like EWC and PackNet.
 However, none of them surpass our RAS.
-Moreover, compared to Replay@20%, RAS achieves superior performance while requiring a smaller constant storage size, equivalent to storing 15% exemplary images from the training set in the Replay methods.
-This storage size is determined by the image dimensions of 800 by 600 pixels and the number of model parameters in RAS.
-It is also worth noting that RAS only requires storing the old model snapshot, the triplet labels in $U_{t}$, and the image generative model, eliminating privacy concerns associated with storing original images.
+RAS achieves superior performance compared to Replay@20% ($\sim$2 Gb) while requiring less storage ($\sim$1.2 Gb), equivalent to storing 15% exemplary images.
+This storage efficiency is due to only needing to store the old model snapshot, triplet labels in $U_{t}$, and the image generation model, thus avoiding privacy concerns.
+Also, RAS outperforms RAS_GT (Tab.
+2), indicating that decomposing scene graphs into smaller, more diverse ones, along with more comprehensible prompts, is more effective than storing the ground truth scene graphs and directly using them for image generation.
 In Learning Scenario 2 (S2), the task involves classifying both new objects and new relationships, significantly escalating the level of difficulty compared to S1.
@@
 This underscores the persistent challenge posed by S2 in the context of CSEGG.
-Future work should explore new approaches to address the complexities inherent in this scenario.
+Future work should explore new approaches to address this gap.
 To tackle the issue of imbalanced data distribution in real-world scenarios, we incorporate two established data sampling techniques (LVIS@Replay@10% and BLS@Replay@10%) into our experiment (Sec.
@@
 This observation suggests that the original sampling methods designed for addressing long-tailed distributions in the classical SGG problem may not be as effective when applied to CSEGG.
-As the knowledge carried forward is important for the subsequent tasks, we also permuted the task sequences and explored their role in CSEGG performances.
-Aligning with the existing literature [54], we found a prominent effect of task sequences in CSEGG (Fig.
-S3 and Sec.
-0.A.4.1).
-To gain a qualitative understanding of CSEGG performances, we provide visualization results of the predicted scene graphs on example images over tasks for all the CSEGG baselines in Scenario 1 (Fig.
-S5 and Sec.
-0.A.5.1) and Scenario 2 (Fig.
-S6 and Sec.
-0.A.5.2).
+We explored the impact of task sequence permutations on CSEGG performance, finding an effect consistent with existing literature [63] (Fig.
+S6; Sec.
+A.6.1).
+We also observed that fine-tuning DETR in S1 has minimal impact on forgetting, indicating that any forgetfulness in S1 is solely due to relationship incremental learning (Fig.
+S7; Sec.
+A.6.2).
+Moreover, to gain qualitative insights, we provide visualizations of predicted scene graphs for all CSEGG baselines in Scenario 1 (Fig.
+S8 and Sec.
+A.7.1) and Scenario 2 (Fig.
+S9 and Sec.
+A.7.2).
 5.2 CSEGG Models Can Generalize in Unknown Scenes
@@
 4 (a), an increasing trend in Gen Rbbox is observed with the increasing task number for all CSEGG methods, indicating improved generalization in detecting unknown objects.
-Notably, even with minimal training in Task 1, all CSEGG methods propose 23% reasonable object regions with IoU = 0.7, showcasing the SGTR model’s ability to generalize “objectness".
+Notably, even with minimal training in Task 1, all CSEGG methods propose 23% reasonable object regions with threshold IoU = 0.7, showcasing the SGTR model’s ability to generalize to locate “objectness".
 As expected, with an increase in IoU threshold from 0.3 to 0.7, we found that Gen Rbbox decreases due to fewer bounding boxes being considered correct.
@@
 However, the performance of Replay@10% is still inferior to our RAS method; implying that our RAS is more proficient in generalizing to classify relationships among unknown objects.
-Interestingly, we also noted that even with minimal training in Task 1, all the CSEGG methods achieve 45% recall of known relationships among unknown objects, demonstrating the SGTR model’s ability to generalize “relationship".
-Visualization examples, when CSEGG models can generalize to recognize relationships, are presented in Fig.S7 and Sec.0.A.5.3.
+Interestingly, we also noted that even with minimal training in Task 1, all the CSEGG methods achieve 45% recall of known relationships among unknown objects, demonstrating the SGTR model’s ability to generalize to classify “relationship".
+Visualization examples, when CSEGG models can generalize to recognize relationships, are presented in Fig.
+S10 and Sec.
+A.7.3.
 5.3 Ablation Studies on Our RAS Reveal Key Design Insights
@@
 5.3.
-Results in Avg.R, F, mR, mF are reported after the last task in the scenario S1.
-We introduce our default method designs in Sec. 4. Here, we vary the components in our RAS to reveal key design insights.
+Results in Avg.R, F, mR, mF are reported after the last task in Scenario S1.
+We introduce our default method designs in Sec.
+4.
+Here, we vary the components in our RAS to reveal key design insights.
 We propose a context checker in RAS.
@@
 4).
-Lastly, we investigate the impact of generating $\gamma$ images using the same text prompt in RAS.
-We vary $\gamma$ from 2 to 8 and present their corresponding model performances in A4-6 of Tab.
-4.
-As expected, the SGG model’s performance improves with higher $\gamma$, indicating that generating more images increases sample diversity for rehearsals and hence, it enhances CSEGG performance.
-We set $\gamma=10$ as the default; however, in practical scenarios with ample compute resources, dynamically synthesizing more images could further improve CSEGG performances.
-This also highlights the advantage of our RAS, as it can dynamically generate numerous images for replays without expanding its storage usage.
+Lastly, we investigate the impact of generating $\gamma$ images using the same text prompt in RAS, varying $\gamma$ from 2 to 8 (Tab.
+4, A4-6).
+As expected, performance improves with higher $\gamma$, showing that increased sample diversity enhances CSEGG performance.
+With ample computing resources, dynamically synthesizing more images could further improve performance.
+This highlights RAS’s advantage in generating numerous images for replays without expanding storage usage.
 6 Discussion
@@
 Extensive experiments demonstrate our RAS’s superior performance over current CSEGG baselines in knowledge transfers and reducing forgetting.
-Interestingly, our RAS model is also capable of generalizing to classify known relationships among unfamiliar objects.
-We discuss ethical concerns in the Sec.0.A.6.
-7 Acknowledgement
-This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-025), its NRFF award NRF-NRFF15-2023-0001, Mengmi Zhang’s Startup Grant from Agency for Science, Technology, and Research (A*STAR), and Early Career Investigatorship from Center for Frontier AI Research (CFAR), A*STAR.
-The authors declare that they have no competing interests.
-The authors would like to thank Stan Weixian Lei, Difei Gao, and Mike Shou for their feedback and suggestions.
-Appendix 0.A Appendix
-0.A.1 Introduction to Three Learning Scenarios
+Interestingly, our RAS model is also capable of generalizing to classify known relationships among unseen objects.
+Moving forward, there are several key avenues for future research.
+First, our current endeavors focus on tackling CSEGG problems from static images in an Independent and Identically Distributed (i.d.d) manner, diverging from how humans learn from video streams.
+Future research can look into CSEGG problems on video SGG datasets.
+Second, our plans also involve expanding the set of continual learning baselines and integrating more long-tailed distribution sampling techniques.
+Third, we aim to construct a synthetic SGG dataset to systematically quantify the aspects of SGG that influence continual learning performance under controlled conditions.
+In RAS, SGG annotations for synthesized images in the replay buffer are predicted by the preceding SGG model, which can lead to error propagation across training iterations.
+In the future work, integrating a generative model with fine-grained control signals (such as bounding boxes and captions) [36] could provide more precise supervision, potentially mitigating these accumulated errors and further enhancing the performance of our approach.
+Although the CSEGG method holds promise for many downstream applications like monitoring systems, medical imaging, and autonomous navigation, we should also be aware of its misuse in privacy, data biases, fairness, security concerns, and misinterpretation (see Sec.
+A.8 for an expanded discussion).
+We invite the research community to join us in expanding and updating the safe use of CSEGG benchmarks, thereby fostering its advancements in research and technology.
+Acknowledgement
+This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-025), and its NRFF award NRF-NRFF15-2023-0001.
+We also acknowledge Mengmi Zhang’s Startup Grant from Agency for Science, Technology, and Research (A*STAR), Startup Grant from Nanyang Technological University, and Early Career Investigatorship from Center for Frontier AI Research (CFAR), A*STAR.
+Appendix A Appendix
+A.1 Introduction to Three Learning Scenarios
 Within this section, we present more details of three learning scenarios and their practical applications.
-Figure S1: Label distribution in each task in each learning scenario is presented.
-In scenario S1 (a) and scenario S3 (c), we use different colors to denote different tasks.
-The color gradient indicates the frequency of data within a task, with the lighter color denoting the smaller frequency of data in that category.
-Only the most frequent labels (relationship labels in (a) and object labels in (c)) are provided.
-See the legend for the total data size per task.
-In (b) scenario S2 on both objects and relationships, data distributions are presented in the form of small-world networks, where nodes denote object categories and the edges linking object pairs indicate relationships.
-Thickness in edges implies the diversity of relationships between object pairs.
-Same color conventions as (a) and (c) are applied.
-See the legend for triplet sizes.
-0.A.1.1 Scenario 1 (S1): Relationship Incremental Learning
+A.1.1 Scenario 1 (S1): Relationship Incremental Learning
 While existing continual object detection literature focuses on incrementally learning object attributes [43, 70, 8, 68, 60, 45], incremental relationship classifications are equally important as it provides a deeper and more holistic understanding of the interactions and connections between objects within a scene.
@@
 Thus, the relationships within a task are long-tailed; and the number of relationships from the head categories of each task is of the same scale.
-To tackle this issue, we allow CSEGG models to see the same images over tasks, but the relationship labels are only provided in their given task (see Sec.
-0.A.1.4 for the design motivation).
+To tackle this issue, we allow CSEGG models to see the same images over tasks, but the relationship labels are only provided in their given task
+The design of such images and label splits over tasks aligns with human learning scenarios where a parent teaches the baby to recognize different toys and objects in the bedroom.
+Though the baby is exposed to the same bedroom scenes multiple times, the parent only teaches the baby to detect and recognize one object at a time in a continual learning setting.
+In the future, we will expand our studies to cases where the SGG models learn from non-overlapping sets of training images for each task.
 The same reasoning applies in S2 and S3.
@@
 In this instance, the identical cancer cell disseminates to fresh organs or tissues, progressively establishing new relationships with other cells over the course of time.
-0.A.1.2 Scenario 2 (S2): Scene Incremental Learning
+A.1.2 Scenario 2 (S2): Scene Incremental Learning
 To simulate the real-world scenario when there are demands for detecting new objects and new relationships over time in old and new scenes, we introduce this learning scenario where new objects $O_{i}$ and new relationship predicates $p_{k}$ are incrementally introduced over tasks (Fig.
@@
 This ensures that the robots can effectively communicate and collaborate while comprehending the evolving scene and tasks, optimizing their construction efficiency and accuracy.
-0.A.1.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
+Figure S1: Label distribution in each task in each learning scenario is presented.
+In scenario S1 (a) and scenario S3 (c), we use different colors to denote different tasks.
+The color gradient indicates the frequency of data within a task, with the lighter color denoting the smaller frequency of data in that category.
+Only the most frequent labels (relationship labels in (a) and object labels in (c)) are provided.
+See the legend for the total data size per task.
+In (b) scenario S2 on both objects and relationships, data distributions are presented in the form of small-world networks, where nodes denote object categories and the edges linking object pairs indicate relationships.
+Thickness in edges implies the diversity of relationships between object pairs.
+Same color conventions as (a) and (c) are applied.
+See the legend for triplet sizes.
+A.1.3 Scenario 3 (S3): Scene Graph Generalization In Object Incremental Learning
 We, as humans, have no problem at all recognizing the relationships of unknown objects with other nearby objects, even though we do not know the class labels of the unknown objects.
@@
 In essence, SGG’s relationship generalization empowers robots to navigate and plan routes in unfamiliar terrains, such as deep sea and deep space, where novel encounters demand adaptable responses.
-0.A.1.4 Design Motivations for Overlapping Images and Non-overlapping Labels across Tasks
-The design of such images and label splits over tasks aligns with human learning scenarios where a parent teaches the baby to recognize different toys and objects in the bedroom.
-Though the baby is exposed to the same bedroom scenes multiple times, the parent only teaches the baby to detect and recognize one object at a time in a continual learning setting.
-In the future, we will expand our studies to cases where the SGG models learn from non-overlapping sets of training images for each task.
-0.A.2 Implementation and Training Details
+Figure S2: Data Statistics for all Learning Scenarios: (a) Distribution of objects in the entire training set of Visual Genome during Stage 1 of S1.
+(b) Distribution of relationships during Stage 2 for each task in S1.
+(c) Distribution of objects during Stage 1 for each task in S2.
+(d) Distribution of relationships during Stage 2 for each task in S2.
+(e) Distribution of objects during Stage 1 for each task in S3.
+(f) Distribution of relationships during Stage 2 for each task in S3.
+The numbers in brackets in the legend in (b-f) denote the number of training images in the particular task.
+Zoom in to the figure to get the exact labels and the frequency associated with them.
+A.2 Data Statistics
+In this section, we provide various types of data statistics for all three learning scenarios.
+Specifically, we present statistics regarding the number of images, objects, and relationships involved in each task of each learning scenario.
+This information is provided in Fig.
+S2.
+A.3 Implementation and Training Details
 Figure S3: Introduction to backbone SGG models and continual learning baselines.
-We use Scene graph Generation TRansformer (SGTR) [28] as the backbone SGG model (Sec.
+We use Scene graph Generation TRansformer (SGTR) [32] as the SGG backbone (Sec.
 3.2).
@@
 The arrows in colors indicate the shifts of network weights in the parameter space when learning Task B for different baselines.
-0.A.2.1 SGTR Backbone
-For SGTR as depicted in (Fig.
-S2 (a)), the approach uniquely formulates the task as a bipartite graph construction problem.
+Figure S4: Detailed Schematic of three CSEGG Baselines.
+In this schematic, we show the operations of each baseline at tasks $t-1$, $t$, $t+1$.
+(a) shows the schematic for exemplar-based methods like replay and RAS.
+(b) shows the schematic for EWC.
+(c) shows the schematic for PackNet.
+Here, $D_{i}$ refers to the dataset of task $i$.
+$E_{i}$ denotes the exemplar set for replays at task $i$.
+$W_{i}$ is the weights of $M_{i}$ at task $i$.
+Figure S5: Visualizations of example generated images given the text prompt from our RAS.
+Each row presents two examples of generated images.
+Given the input text prompt on the left, the generated image is displayed on the right.
+These visualizations provide qualitative validations into the capability of RAS to produce diverse and contextually relevant images based on our designed text prompts.
+A.3.1 SGTR Backbone
+For SGTR in Fig.
+S3 (a), the approach uniquely formulates the task as a bipartite graph construction problem.
 Starting with a scene image ($I_{i}$), SGTR utilizes a 2D-CNN and transformer-based encoder to extract image features.
@@
 However, empirical results suggest that fine-tuning transformer-based encoders in DETR helps downstream predicate predictions [32].
-Even with fine-tuning DETR in S1, we verify that there is minimal forgetting of detecting all objects in the scene over tasks (see Sec.
-S4 and Sec.
-0.A.4.2).
+Even with fine-tuning DETR in S1, we verify that there is minimal forgetting of detecting all objects in the scene over tasks (see Fig.
+S7 and Sec.
+A.6.2).
 Thus, the forgetting observed in S1 could only be attributed to incremental relationship learning.
@@
 All models are trained on 4 A5000 GPUs.
-0.A.2.2 TCNN Backbone
+A.3.2 TCNN Backbone
 As for TCNN, it employs Faster-RCNN [18] to generate object proposals from a scene image ($I_{k}$).
@@
 In Learning Scenario 1 (S1), similar to the implementation details of SGTR in Sec.
-0.A.2.1, we skip Stage 1, and directly load pre-trained weights of Faster-RCNN for object detection on the entire training set of Visual Genome [28].
+A.3.1, we skip Stage 1, and directly load pre-trained weights of Faster-RCNN for object detection on the entire training set of Visual Genome [32].
 In stage 2 of S1, we load the pre-trained weights of the visual feature extractor (pre-trained on ImageNet) and fine-tune the rest parts of the model.
@@
 All models are trained on 4 A5000 GPUs.
-0.A.3 Evaluation Metrics
+A.4 Mathematical formulations of Continual learning baselines
+Here, we introduce mathematical formulations of these baselines: (1) To train Naive baseline on CSEGG, at task $t$, we take the previously trained model $M_{t-1}$ with weights $W_{t-1}$ and train it on the dataset $D_{t}$ at task $t$, to obtain weights $W_{t}$ of the model $M_{t}$.
+(2) For EWC, we use $W_{t}$ and $M_{t}$ to calculate $F_{t+1}$, where $F$ is the Fisher information matrix using the equation, $F_{t+1}=-E[\frac{\partial^{2}}{\partial W_{t}^{2}}log(M_{t}(x)|W_{t})]$.
+During the training of task $t+1$, we add $L_{EWC}$ to the training loss, where $L_{EWC}=\sum F_{t+1}(W_{t+1}-W_{t})^{2}$.
+(3) For PackNet, after training of task $t$, we take $W_{t}$ and apply a pruning algorithm to obtain the pruned weights $W_{t}^{\prime}$.
+At task $t+1$, we obtain $M_{t+1}$ by training $M_{t}$ with $W_{t}^{\prime}$ on $D_{t+1}$.
+(4) For Replay@M, we create exemplar set $E_{t}$ at the replay buffer by storing data points from $D_{t-1}$ after training $M_{t-1}$ on task $t-1$ At task $t$, we obtain weights $W_{t}$ of $M_{t}$ by training $M_{t}$ on $E_{t}$ and $D_{t}$.
+(5) For the joint training, we train one model $M_{T}$ jointly on all the datasets $\{D_{t}\}$, where $t\in\{1,2,...,T\}$.
+(6) For RAS_GT, we adopt the similar math formulations as Replay@M.
+See more details in Sec.
+4.
+A.5 Evaluation Metrics
 To assess the catastrophic forgetting of CSEGG models, we define Forgetfullness (F@K) as the difference in R@K on $D_{t=1}$ between the CSEGG models trained at task $t$ and task 1.
@@
 FWT is defined as $FWT@K=\frac{1}{T-1}\sum_{i=2}^{T}R@K_{i,i}\>-\>\overline{b@K}_{i,i}$, where $\overline{b@K}_{i,i}$ is the R@K for an independent model with random initialization trained in task $i$ and tested in task $i$.
-In learning scenario S3, we evaluate CSEGG models on their abilities to generalize to detect unknown objects and classify known relationships on these objects, in the standalone generalization test set over all tasks.
-To benchmark these, we introduce two evaluation metrics: the recall of the predicted bounding boxes on unknown objects (Gen Rbbox@K) and the recall of the predicted graph $G_{i}$ (Gen R@K).
-As the CSEGG models have never been taught to classify unknown objects, we discard the class labels of the bounding boxes and only evaluate the predicted box locations with Gen Rbbox@K.
-To evaluate whether the predicted box location is correct, we apply a hard threshold of Intersection over Union (IoU) between the predicted bounding box locations and the ground truth.
-Any predicted bounding boxes with their IoU values above the hard threshold are deemed to be correct.
-We vary IoU thresholds from 0.3, 0.5, and 0.7.
-To assess whether the CSEGG model generalizes to detect known relationships over unknown objects, we evaluate the recall Gen R@K of the predicted relationships $r_{k}$ only on correctly predicted bounding boxes.
-0.A.4 More Quantitative Result Analysis on Continual SGTR Methods
-0.A.4.1 Results for Different Task Sequences
-Recent work [54] has highlighted the consistent and substantial curriculum effects in class-incremental learning and continual visual question-answering tasks.
-Inspired by these findings, we conducted experiments to assess the impact of the curricula in the context of CSEGG.
-To delve into its potential influence, we trained baselines with three distinct task sequences for Learning Scenario 1.
-Our results demonstrated that curriculum learning indeed shapes CSEGG performance within class-incremental settings.
-Notably, in Fig.
-S3(d), a large difference in Avg.R@20 between Order 1 and Order 3 emerges for the Replay (100%) baseline.
-Similarly, Fig.
-S3(a) reveals a substantial Avg.R@20 disparity between Order 2 and Order 3 for the Naive baseline.
-This trend extends to F@20, as depicted in Fig.
-S3(e)(f)(g)(h).
-These insights collectively affirm the significance of the curricula within CSEGG.
 Figure S6: Results of F@K=20, Avg.
@@
 The higher F, Avg.R@20, the better.
-0.A.4.2 Minimal Forgetting in DETR
-To validate the impact of fine-tuning the DETR model in training Stage 2 of learning scenario S1 on relationship predicate predictions and to ensure minimal forgetting occurs in object detection (Sec 0.A.2), we compare the mean Average Precision (mAP) for object detection on the entire test set of Visual Genome between the pre-trained DETR checkpoint from the paper [28] and the DETR models after fine-tuning on each task of S1.
+In learning scenario S3, we evaluate CSEGG models on their abilities to generalize to detect unknown objects and classify known relationships on these objects, in the standalone generalization test set over all tasks.
+To benchmark these, we introduce two evaluation metrics: the recall of the predicted bounding boxes on unknown objects (Gen Rbbox@K) and the recall of the predicted graph $G_{i}$ (Gen R@K).
+As the CSEGG models have never been taught to classify unknown objects, we discard the class labels of the bounding boxes and only evaluate the predicted box locations with Gen Rbbox@K.
+To evaluate whether the predicted box location is correct, we apply a hard threshold of Intersection over Union (IoU) between the predicted bounding box locations and the ground truth boxes.
+If the IoU exceeds the predetermined threshold, it is considered a true positive ($TP$).
+We used IoU thresholds of 0.3, 0.5, and 0.7 for our experiments.
+If the IoU exceeds the threshold multiple times for the same predicted bounding box, we consider it a single positive prediction.
+This is because the metric aims to define the model’s performance in locating bounding boxes of objects that have not been learned during training.
+Counting multiple times for the same box would be misleading, as it would inflate the number of $TP$s and recall, while the actual number of unknown bounding boxes the model can generate might be low.
+The total possible positives ($TP+FN$) are determined by the total number of ground truth boxes in the image.
+In general, Gen Rbbox@K helps evaluate how well the CSEGG model locates unknown objects within an image.
+True positives ($TP$) represent successful identification of the location of an unknown object, while $TP+FN$ represents all possible unknown objects the CSEGG model could locate.
+Thus, object classification labels are not required to calculate Gen Rbbox@K.
+Instead, we need the total number of ground truth bounding boxes in an image and the number of predicted boxes that meet the IoU thresholds.
+To assess whether the CSEGG model generalizes to detect known relationships over unknown objects, we evaluate the recall Gen R@K of the predicted relationships $r_{k}$ only on correctly predicted bounding boxes.
+A.6 More Result Analysis on Continual SGTR Methods
+A.6.1 Results for Different Task Sequences
+Recent work [63] has highlighted the consistent and substantial curriculum effects in class-incremental learning and continual visual question-answering tasks.
+Inspired by these findings, we conducted experiments to assess the impact of the curricula in the context of CSEGG.
+To delve into its potential influence, we trained baselines with three distinct task sequences for Learning Scenario 1.
+Our results demonstrated that curriculum learning indeed shapes CSEGG performance within class-incremental settings.
+Notably, in Fig.
+S6(d), a large difference in Avg.R@20 between Order 1 and Order 3 emerges for the Replay (100%) baseline.
+Similarly, Fig.
+S6(a) reveals a substantial Avg.R@20 disparity between Order 2 and Order 3 for the Naive baseline.
+This trend extends to F@20, as depicted in Fig.
+S6(e)(f)(g)(h).
+These insights collectively affirm the significance of the curricula within CSEGG.
+A.6.2 Minimal Forgetting in DETR
+To validate the impact of fine-tuning the DETR model in training Stage 2 of learning scenario S1 on relationship predicate predictions and to ensure minimal forgetting occurs in object detection (Sec A.3), we compare the mean Average Precision (mAP) for object detection on the entire test set of Visual Genome between the pre-trained DETR checkpoint from the paper [32] and the DETR models after fine-tuning on each task of S1.
 As shown in Fig.
-S4, the results indicate a slight decrease of 0.4 in mAP from the pre-trained checkpoint to the DETR models over 5 tasks.
+S7, the results indicate a slight decrease of 0.4 in mAP from the pre-trained checkpoint to the DETR models over 5 tasks.
 This study provides evidence that fine-tuning DETR in S1 has negligible effects on forgetting.
@@
 We observed a minimal decrease in mAP across all tasks.
-0.A.5 Visualization Examples for All Learning Scenarios for Continual SGTR-based Models
+A.7 Visualization Examples for All Learning Scenarios for Continual SGTR-based Models
 In this section, we present visualization examples from each learning scenario to showcase the performance of the three continual SGTR-based models, namely Replay@10%, EWC, and Naive, in three learning scenarios.
-0.A.5.1 Learning Scenario 1 (S1)
+A.7.1 Learning Scenario 1 (S1)
 Figure S8: Visualization example for Learning Scenario 1 (S1).
@@
 As explained in Fig.
-S5 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet.
+S8 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet.
 The time arrow on the left side of the figure demonstrates that the model is exposed to new data over time, with new objects and relationships incrementally added, as described in Sec.
@@
 The EWC model also fails to remember any triplets from the previous task and only predicts "trunk behind women".
-0.A.5.2 Learning Scenario 2 (S2)
+A.7.2 Learning Scenario 2 (S2)
 From Fig.S9, we observe that, in Task 1, the ground truth scene graph contains triplets: “man riding skateboard”, “man above skateboard”, and “shoe of skateboard”.
@@
 Similar to Fig.
-S5, and S6, the red boxes in all scene graphs indicate objects , while green boxes represent relationships.
+S8, and S9, the red boxes in all scene graphs indicate objects , while green boxes represent relationships.
 As explained in Fig.
-S5 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet.
+S8 caption, the direction of the arrows between the red (object) and green (relationship) boxes indicates the subject and object ordering in the triplet.
 The time arrow on the left side of the figure demonstrates that the model is exposed to new objects over time as described in Sec.
@@
 Concretely, unk_1 refers to "mountain"; unk_2 is "house"; and unk_3 is "sheep".
-0.A.5.3 Learning Scenario 3 (S3)
+A.7.3 Learning Scenario 3 (S3)
 Fig.
-S7 illustrates the performance of the Replay@10% and Naive models in locating unknown objects and recognizing the relationships between these objects and other nearby unknown objects.
+S10 illustrates the performance of the Replay@10% and Naive models in locating unknown objects and recognizing the relationships between these objects and other nearby unknown objects.
 The ground truth in Fig.
-S7 consists of three unknown objects: "mountain", "sheep", and "house", along with three relationships: "near", "behind", and "infront of" (between "mountain" and "house"), and a "near" relationship (between "sheep" and "house").
+S10 consists of three unknown objects: "mountain", "sheep", and "house", along with three relationships: "near", "behind", and "infront of" (between "mountain" and "house"), and a "near" relationship (between "sheep" and "house").
 After training on Task 1 data, the Naive model can accurately locate the objects "mountain" and "house".
@@
 After Task 4, as it can now locate the object "sheep", the Replay@10% model can also predict the "near" relationship between the objects "house" and "sheep", in addition to the existing relationships between "mountain" and "house".
-0.A.6 Ethical Concerns
+A.8 Ethical Concerns
 The development and deployment of Scene Graph Generation (SGG) technology present potential negative societal impacts that warrant careful consideration [32].
@@
 For example, attackers might manipulate scene graphs to deceive systems or disrupt applications that rely on scene understanding.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Sec. 5
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Sec. 6
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: We don’t have any theoritical proofs.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Sec. A.3
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We have provided the code as the Supplementary material.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Sec. A.3
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: In Sec. A.3, we have provided information about the statistical significance of the experiments.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Sec. A.3
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We have followed the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Sec. A.8
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [Yes]
+  54. Justification: We have used Stable Diffusion models for the image generation. So we have ensured to put their safegaurds during the image generation process.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We have credited the authors of the models and methods we are using for our work.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: Sec. 4
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: We have not done any crowdsourcing or research involving human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: We have not done any crowdsourcing or research involving human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.